### v1

# Rotation sets and axes in the fine curve graph for torus homeomorphisms

Abstract.
We expand the dictionary between the action of a torus homeomorphism on the fine curve graph and its rotation set. More precisely, we show that the fixed points at infinity of a loxodromic element determine the rotation set up to scale. A key ingredient is a metric version of the classical WPD property from geometric group theory.
As a consequence we find new stable criteria for positive scl, and for two homeomorphisms to generate a free group.
1. Introduction
1.1. Rotation sets and algebra
The aim of the present paper concerns the group $\mathrm{Homeo}_{0}(T^{2})$ of homeomorphisms of the two-dimensional torus that are isotopic to the identity. Our results relate the rotation set, a classical invariant from dynamical systems, to the algebra of this group, via its action on the fine curve graph. We will recall the definition of the rotation set in section 2 below; for the moment, let us just say that it is a compact and convex subset of the plane, associated to a given element of our group $\mathrm{Homeo}_{0}(T^{2})$, which is defined up to integer translations.
Theorem 1.
Suppose $f,g\in\mathrm{Homeo}_{0}(T^{2})$ so that the rotation sets $\mathrm{Rot}(f),\mathrm{Rot}(g)$ have interior, and are not homothetic. Then there exists $N>0$ such that $f^{N}$ and $g^{N}$ generate a free group, all elements of which have a rotation set with non-empty interior.
Here two sets $C,C^{\prime}$ are said to be homothetic if there exist $\lambda\in\mathbb{R}\setminus\{0\}$ and $v\in\mathbb{R}^{2}$ such that $C^{\prime}=\lambda C+v$. In the next theorem the set $C$ is said to have a point symmetry if there is some vector $v$ such that C=C+vC=-C+v.
Theorem 2.
Suppose $f\in\mathrm{Homeo}_{0}(T^{2})$ so that the rotation set $\mathrm{Rot}(f)$ has interior, and $\mathrm{Rot}(f)$ has no point symmetry. Then $\mathrm{scl}(f)>0$ in $\mathrm{Homeo}_{0}(T^{2})$.
Here is the definition of the stable commutator length ($\mathrm{scl}$). Given a group $G$ generated by a subset $\mathcal{S}$, the word length $\ell_{\mathcal{S}}(g)$ of $g\in G$ is the smallest number of elements of $\mathcal{S}$ whose product is $g$. The stable word length is
$s\ell_{\mathcal{S}}(g)=\lim\frac{1}{n}\ell_{\mathcal{S}}(g^{n})$
where the limit always exists by sub-additivity. If the group $G$ is perfect, i.e. generated by the set $\mathcal{C}$ of commutators, then $\ell_{\mathcal{C}}(g)$ and $s\ell_{\mathcal{C}}(g)$ are denoted $\mathrm{cl}(g),\mathrm{scl}(g)$ and called the (stable) commutator length. By Bavard duality, positivity of $\mathrm{scl}(g)$ is equivalent to the existence of a homogeneous quasimorphism $\Phi$ such that $\Phi(g)\neq 0$ ([Bav91]). The identity component of the group of homeomorphisms of any compact manifold is known to be perfect ([And58]), in particular the definition applies to our group $\mathrm{Homeo}_{0}(T^{2})$.
By Kwapisz realization theorem [Kwa92], every polygon with all vertices having rational coordinates (for short, rational polygon) can be realized as the rotation set of some element of $\mathrm{Homeo}_{0}(T^{2})$, and the proof gives an explicit construction as an axiom A diffeomorphism. This provides plenty of examples to illustrate both theorems. Note that a symmetric polygon has an even number of sides; in particular, triangles are never symmetric. The easiest example satisfying Theorem 2 is obtained for a homeomorphism whose rotation set is the triangle $(0,0),(1,0),(0,1)$; we will give a construction close to that of Kwapisz in section 11.2. On the other hand every symmetric quadrilateral $Q$ (parallelogram!) which is rational is the rotation set of some $f$ which has a power conjugate to its inverse, which easily implies $\mathrm{scl}(f)=0$, see section 11.1. This shows some kind of optimality for Theorem 2. By the continuity property of the rotation set, the hypotheses of both theorems are stable: the second theorem describes an open set in $\mathrm{Homeo}_{0}(T^{2})$, and the first one an open set of its cartesian square.
The existence of elements with positive scl is not new: in [BHW22], Bowden, Hensel and Webb proved this in $\mathrm{Homeo}_{0}(\Sigma)$ for every surface $\Sigma$ of genus at least one. This was the original motivation for the introduction of the fine curve graph (see below), the elephant in the room in Theorem 1 and 2. The proof in [BHW22] was rather constructive but did not provide any dynamical feature of the exhibited elements. In [BHW24], the same authors provide more specific examples, namely (relative) pseudo-Anosov maps. We refer the reader to these previous papers for more context and motivation. The following corollary is new.
Corollary 1.1.
Positive scl happens arbitrarily close to the identity in $\mathrm{Homeo}_{0}(T)$.
Proof of Corollary 1.1.
Take any $f$ satisfying the hypotheses of Theorem 2, e.g. whose rotation set is a rational triangle. For $m>0$, let $B_{m}=m\mathrm{Id}$, and consider the rescaling $f_{m}=B_{m}^{-1}fB_{m}$ (see Lemma 2.2 below). The sequence $(f_{m})$ converges uniformly to the identity. On the other hand the rotation set of $f_{m}$ is homothetic to that of $f$, thus it also satisfies the hypotheses, and the conclusion, of Theorem 2, namely $\mathrm{scl}(f_{m})>0$. ∎
Finally, note that Theorem 2, together with Kwapisz’ construction, also provides $C^{\infty}$-diffeomorphisms whose stable commutator length is positive within the group of $C^{\infty}$-diffeormophisms, since the commutator length in a subgroup is obviously not less than the commutator length in the group.
1.2. Rotation sets and the fine curve graph
The fine curve graph $\mathcal{C}^{\dagger}(\mathbb{T}^{2})$, introduced by Bowden, Hensel and Webb in [BHW22], is the graph whose vertices are essential simple closed curves, and edges join curves that are either disjoint or have only one intersection point, at which they cross transversely. This graph is Gromov-hyperbolic ([BHW22, Theorem 3.10]). The group $\mathrm{Homeo}_{0}(T^{2})$ acts naturally on this graph, and identifies with its automorphism group ([LMP+21, LRW24]). The translation length $\mathrm{tl}(f)$ of $f\in\mathrm{Homeo}_{0}(T^{2})$ is defined as
$\mathrm{tl}(f)=\lim\frac{1}{n}d_{\mathcal{C}^{\dagger}(T^{2})}(\alpha,f^{n}\alpha),$
where $d_{\mathcal{C}^{\dagger}(T^{2})}$ denotes the distance in the graph, the limit exists by sub-additivity and does not depend on the choice of $\alpha$. An element $f$ is said to act loxodromically if $\mathrm{tl}(f)>0$. The starting point for our work is the following result by Bowden, Hensel, Mann, Militon and Webb, who discovered that the action on the fine curve graph is intimately related to the rotation set.
Theorem ([BHM+22, Theorem 1.3]).
Let $f:T\to T$ be a torus homeomorphism which is isotopic to the identity. Then the following are equivalent:
  1. i)
    $f$ acts loxodromically on the fine curve graph $\mathcal{C}^{\dagger}(T)$, and
  2. ii)
    The rotation set $\mathrm{Rot}(f)$ has nonempty interior.
We remark that the theorem is still true in higher genus if one replaces the rotation set by the ergodic rotation set ([GM23]), and that both conditions are equivalent to $f$ defining a pseudo-Anosov mapping class relative to a finite set of periodic points. Neither of these are relevant for the current work.
A Gromov hyperbolic space $X$ has a natural boundary $\partial_{\infty}X$, which roughly speaking represents the different ways to go to infinity (see section 3). A loxodromic element $f$ acting on the hyperbolic space $X$ has exactly two fixed points at infinity, an attracting fixed point $\xi^{+}(f)$ that attracts all points of $X$ and a repulsive fixed point $\xi^{-}(f)$ that attract all points in the past. We will prove that the attracting fixed point determines the rotation set up to homothety.
Theorem 3.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be two homeomorphisms acting loxodromically on the fine curve graph. Assume $\xi^{+}(f)=\xi^{+}(g)$. Then the scaled rotation sets
$\frac{1}{\mathrm{tl}(f)}\mathrm{Rot}(f)\text{ and }\frac{1}{\mathrm{tl}(g)}\mathrm{Rot}(g)$
are translates of each other.
The existence of free groups claimed by Theorem 1 follows easily from theorem 3 by the classical ping-pong argument (see [dlH00]). In section 11 we show how to determine the scaled rotation sets on some interesting examples. In particular, we also get a free group generated by two elements having the same rotation set but distinct scaled rotation sets. By using a different invariant (the twist number, see section 1.4 below), we also get free groups generated by elements sharing the same scaled rotation sets.
To get positive stable commutator length, we need a stronger property. In [BF02], Bestvina and Fujiwara define an equivalence relation on loxodromic elements of the isometry group of a Gromov-hyperbolic space, namely $f\sim g$ whenever $g$ has conjugate elements whose repulsive and attracting fixed points in $\partial_{\infty}X$ are arbitrarily close respectively to the repulsive and attracting fixed points of $f$ (see section 3.4 for more details). The following statement generalizes Theorem 3.
Theorem 4.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be two homeomorphisms acting loxodromically on the fine curve graph. Assume $f\sim g$. Then the scaled rotation sets of $f$ and $g$ are translates of each other.
We will actually get the stronger result that, roughly speaking, the rotation set depends semi-continuously on the attracting fixed point. For a precise statement, see Theorem 10.2 below. We can now get positive scl.
Theorem 2 from Theorem 4.
Let $f\in\mathrm{Homeo}_{0}(T^{2})$ whose rotation set has non-empty interior and no point symmetry. Since $\mathrm{Rot}(f^{-1})=-\mathrm{Rot}(f)$, and $\mathrm{tl}(f^{-1})=\mathrm{tl}(f)$, the absence of point symmetry yields that the scaled rotation sets of $f$ and $f^{-1}$ are not translates of each other. From Theorem 4 we get that $f\not\sim f^{-1}$. The result now follows from the work of Bestvina-Fujiwara [BF02], see section 3.4 below. ∎
1.3. The metric WPD property
The heart of our argument is the discovery that loxodromic elements of the fine curve graph share some kind of weak proper discontinuity property that we describe now. This property is a natural extension of Bestvina-Fujiwara WPD property ([BF02], see also [Osi16]) to the setting of Rosendal’s geometric group theory [Ros21], which enlarges geometric group theory to enclose some non locally compact groups.
Let $G$ be a topological group, which for simplicity we assume to be connected. Call a subset $A$ of $G$ bounded if for every neighborhood $V$ of the identity, there exists $n$ such that $A\subseteq V^{n}$, where $V^{n}$ denotes the set of products of at most $n$ elements of VV. [^1] These sets are called relatively (OB) by Rosendal. Assume $G$ acts by isometries on a metric space $X$. For every $x,y\in X$ and $r>0$, we define the set $Z_{r}(x,y)$ of elements $h$ of $G$ such that $d(x,hx)\leq r$ and $d(y,hy)\leq r$.
Let $f\in G$. Assume $f$ acts loxodromically, that is,
$\lim\frac{1}{n}d(x,f^{n}x)>0$
for some (every) $x\in X$.
Definition 1.2.
We say that $f$ satisfies the metric WPD property if for every $x\in X$ and every $r>0$, there exists $n_{0}$ such that the set $Z_{r}(x,f^{n_{0}}x)$ is bounded in $G$.
This definition is particularly relevant when the group $G$ is ”locally (OB)”, which means that there exists a neighborhood $O$ of the identity which is bounded. Then the word distance for the generating set $O$ is maximal in the sense of Rosendal: a subset of $G$ is bounded (according to the above definition) if and only if it is bounded for this distance, if and only if it is bounded for every right-invariant distance which is compatible with the topology. In this context, Rosendal’s result is that these bounded neighborhoods of the identity can compensate for the absence of compact neighborhoods of the identity, and plays their role as ”small” generating sets of $G$. Note that when $G$ is a finitely generated group endowed with the discrete topology, the maximal metric is given by the usual word length, and bounded sets are just finite sets. Thus in this case our metric WPD property, with the appropriate definition of bounded sets, amounts to the classical WPD property.
Mann-Rosendal’s theorem states that for every compact manifold, the identity component of the group of homeomorphisms is locally (OB), and the fragmentation norm with respect to any given cover by open balls is quasi-isometric to the maximal metric [MR18]. Back to our context, we propose to use the following concrete incarnation of the maximal metric. For $f,g\in\mathrm{Homeo}_{0}(T^{2})$, define
[…]
Our theorems will all be consequences of the following proposition.
Proposition 1.3.
Every element of $\mathrm{Homeo}_{0}(S)$ acting loxodromically satisfies the metric WPD property.
Another consequence of this metric WPD property is the following characterization of attracting fixed points of loxodromic elements.
Corollary 1.4 (Metric characterization of fixed points of loxodromic).
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph. Then $\xi^{+}(f)=\xi^{+}(g)$ if and only if there exists a sequence of positive numbers $(m_{n})_{n\geq 0}$ such that the sequence $(f^{-n}g^{m_{n}})_{n\geq 0}$ is bounded in the group $\mathrm{Homeo}_{0}(T^{2})$.
Theorem 3 above is actually an easy consequence of this corollary. Let us describe another dynamical consequence. Given $f\in\mathrm{Homeo}_{0}(T^{2})$, consider the equivalence relation on the torus defined by
$x\sim_{f^{-}}y\Leftrightarrow\text{ the sequence }(\|\widetilde{f}^{n}(\widetilde{x})-\widetilde{f}^{n}(\widetilde{y})\|)_{n\leq 0}\text{ is bounded}$
where the norm is the sup norm in $\mathbb{R}^{2}$, $\widetilde{f}$ denotes any lift of $f$, and $\widetilde{x},\widetilde{y}$ any lifts of $x,y$ (the property does not depend on the choices of the lifts). We call the partition by equivalence classes the backward shadowing partition. Note that all the points of a given class share the same backward rotation property, e.g. if any point admits a rotation vector then all the other points admit the same rotation vector (again, see section 2 for the definitions). The following corollary is immediate.
Corollary 1.5.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ be acting loxodromically on the fine curve graph, and assume $\xi^{+}(f)=\xi^{+}(g)$. Then the backward shadowing partitions of $f$ and $g$ coincide. Any $h\in\mathrm{Homeo}_{0}(T^{2})$ that fixes $\xi^{+}(f)$ preserves the relation $\sim_{f^{-}}$, that is, every image of an equivalence class is an equivalence class.
In the previous corollary, it may seems paradoxical that the attracting fixed point on the fine curve graph is related to the backward shadowing relation on the torus. This comes from the fact that the action on the fine curve graph is a left action, whereas the natural metric on the group is right invariant. Namely, if we have a bound, say, on the $C^{0}$ distance between $f^{-n}$ and $g^{-m}$ for large positive numbers $m$ and $n$, then we have the same bound on $f^{-n}g^{m}$, this will translate into a bound in the fine curve graph between a given curve $\alpha$ and $f^{-n}g^{m}\alpha$, and thus also between $f^{n}\alpha$ and $g^{m}\alpha$.
Inhyeok Choi has independently proved that relative pseudo-Anosov homeomorphisms have metric WPD in every genus, see [Cho25]. His proof uses very different methods than our approach, using an explicit description of the fixed points at infinity as singular foliations on the surface from [BHW24].
1.4. Ingredients for metric WPD
The strategy for proving the metric WPD property goes through the fine marking graph, whose vertices are fundamental domains of the torus homeomorphic to the unit square, and whose edges correspond to moving one side of the square in the fine curve graph. This graph is a combinatorial model for our group $\mathrm{Homeo}(T^{2})$: indeed, the orbit map of the natural group action is a quasi-isometry. Thus, bounding the distance in the group, as required by the WPD property, amounts to bounding the distance in the fine marking graph.
Inspired by the marking graph hierarchy formula of Masur-Minsky [MM00] in the classical curve graph setting, we get an upper bound for the distance in the fine marking graph, in terms of (1) the distance in the fine curve graph between the curves corresponding to the sides of the fine markings, and (2) the sum of the twist numbers in suitable annuli (Lemma 8.1). Here the twist number in an annulus $A$ between two curves that cross $A$ is the distance in the fine arc graph of $A$ between the pieces of the curves that cross $A$. In other words, this is the fine version, in genus 1, of subsurface projections. Another key ingredient is a fine version of the Bounded Geodesic Image Theorem (Lemmas 6.4 and 6.5), from which we deduce that twist numbers between two curves along a given loxodromic axis are bounded independently of the annulus where we project (Lemma 6.7). The WPD property is then an easy consequence of these ingredients.
The twists numbers seem to be of independent interest: indeed, taking the supremum on all annuli provides a new conjugacy invariant on $\mathrm{Homeo}_{0}(T^{2})$ that allows to distinguish elements sharing both the same rotation sets and translation length. We will derive the first properties of this invariant, including some semi-continuity, and show that it can take essentially all possible values by computing it on a classical family of examples.
A first proof of Theorem 3 involved train tracks via Masur-Minsky nesting lemma and the fact that iterates of a given curve determines the rotation set (see [Le 24]). We were not able to prove positive stable commutator length along the same lines.
1.5. Organization of the paper
Sections 23 and 4 are preliminaries dealing respectively with the rotation set, hyperbolic geometry and the fine curve graph. The fine marking graph and its relation with the group $\mathrm{Homeo}(T^{2})$ is described in section 5. In section 6 we introduce the fine twist numbers, prove the bounded geodesic image theorems, and deduce that twist numbers are bounded along loxodromic axes. In Section 7 we discuss a variant of the fine curve graph and prove technical distance bounds which are the main ingredient for a distance formula in the fine marking graph that is proved in Section 8. We prove the metric WPD property in section 9. The main theorems are proved in section 10. The last section 11 is dedicated to describing several examples, and also contains the definition of the twist number of a homeomorphism.
2. The rotation set
2.1. Definition
We recall the definition and main properties of the rotation set, following [MZ90]. Here the torus is $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. Let $f\in\mathrm{Homeo}_{0}(T^{2})$, and fix a lift $\widetilde{f}$ of $f$: this is an element of the group $\mathrm{Homeo}(\mathbb{R}^{2},\mathbb{Z}^{2})$ of homeomorphisms of the plane that commutes with integer translations. Consider the function D(f~)D($\widetilde${f}}):
[…]
This function descends to a function $T^{2}\to\mathbb{R}^{2}$ that we still denote D(f~)D($\widetilde${f}).
Definition 2.1.
[…]
$v=\lim\frac{1}{n_{k}}D(\widetilde{f}^{n_{k}})(x_{k}).$
In this definition, whenever the sequence $(x_{k})$ is constantly equal to some point $x$, and the sequence $(n_{k})$ may be taken to be the sequence of all integers, then $v$ is also said to be the rotation vector of $x$, and we write $\mathrm{Rot}_{\widetilde{f}}(x)=v$. In particular, a fixed point $x$ of $f$ has a rotation vector $\mathrm{Rot}_{\widetilde{f}}(x)=D(\widetilde{f})(x)$ which is an integer. Choosing another lift $\widetilde{f}+(p,q)$ of $f$ has the effect of translating the rotation set by $(p,q)$. Thus the rotation set $\mathrm{Rot}(f)$ is defined as a subset of the plane, defined modulo integer translations.
The rotation set has the following properties.
Theorem (Misiurewicz-Ziemian, [MZ90]).
[…]
$\frac{1}{n}\widetilde{f}^{n}(\Delta)$
[…]
$).\mathrm{Rot}(A\widetilde{f}A^{-1})=A(\mathrm{Rot}(\widetilde{f}$)).
Let $B$ be a matrix with integer coefficients which is invertible as a real matrix, but with determinant maybe not equal to $1$. The homeomorphism $B^{-1}\widetilde{f}B$ still belongs to the group $\mathrm{Homeo}(\mathbb{R}^{2},\mathbb{Z}^{2})$, and thus it induces an element of $\mathrm{Homeo}_{0}(T^{2})$; beware that this element depends on the choice of the lift $\widetilde{f}$. When applied with $B=B_{m}=m\mathrm{Id}$ for some $m>1$, we call this element a rescaling of $f$ of order $m$. The previous formula still holds for $A=B^{-1}$.
Lemma 2.2 ([Kwa92]).
For every matrix $B$ with integer coefficients and non-zero determinant,
$).\mathrm{Rot}(B^{-1}\widetilde{f}B)=B^{-1}(\mathrm{Rot}(\widetilde{f}$)).
The following lemma states that the set of realizable rotation sets is invariant under the rational affine group. It is probably well known, we will use it to construct examples in section 11.
Lemma 2.3.
The set
$\{\mathrm{Rot}(\widetilde{f}),\widetilde{f}\text{ lift of some }f\in\mathrm{Homeo}_{0}(T^{2})\}$
is invariant under the action of the affine group $\mathrm{GA}_{2}(\mathbb{Q})$.
More precisely, given any lift $\widetilde{f}$ of some $f\in\mathrm{Homeo}_{0}(T^{2})$, and $\Phi\in\mathrm{GA}_{2}(\mathbb{Q})$, we can find an invertible matrix $A$ with integer coefficients, a vector $v\in\mathbb{Q}^{2}$ and a positive integer $p$ such that the map $\widetilde{g}=A^{-1}\widetilde{f}^{p}A$ commutes with the translation by vector $v$, and $)\mathrm{Rot}(\widetilde{g}+v)=\Phi(\mathrm{Rot}(\widetilde{f}$)).
Proof.
Fix a lift $\widetilde{f}$ of some element $f\in\mathrm{Homeo}_{0}(T^{2})$. Given a positive integer $m$, let $B=m\mathrm{Id}$. By Lemma 2.2 and the homogeneity, the map $\widetilde{g}=B^{-1}\widetilde{f}^{m}B$ has the same rotation set as $\widetilde{f}$, and if we pick some vector $v=\frac{1}{m}(p,q)$, then $\widetilde{g}$ commutes with the translation by $v$. We deduce that $\mathrm{Rot}(\widetilde{g}+v)=\mathrm{Rot}(\widetilde{f})+v$. Thus the set of realizable rotation sets is invariant under the action of rational translations.
It remains to check invariance under the linear group $\mathrm{GL}_{2}(\mathbb{Q})$. Let $A$ be a diagonal matrix with diagonal entries $\frac{p}{q},\frac{p^{\prime}}{q}$, and $B$ the diagonal matrix with diagonal entries $p^{\prime}q,pq$. Then
$).\mathrm{Rot}(B^{-1}\widetilde{f}^{pp^{\prime}}B)=A(\mathrm{Rot}(\widetilde{f}$)).
Since $\mathrm{GL}_{2}(\mathbb{Q})$ is generated by $\mathrm{SL}_{2}(\mathbb{Z})$ together with such matrices $A$, the result follows. ∎
2.2. Rotation set bounds
We start with a well known lemma giving an easy upper bound for the rotation set. Here, and throughout, we denote by $\mathrm{Conv}(X)$ the convex hull of a set $X$ in the plane.
Lemma 2.4 (Upper bound lemma).
For every lift $\widetilde{f}$ of $f\in\mathrm{Homeo}_{0}(T^{2})$, for every $n>0$, we have
$\mathrm{Rot}(\widetilde{f})\subset\mathrm{Conv}\left(\frac{1}{n}(\widetilde{f}^{n}-\mathrm{Id})([0,1]^{2})\right).$
Proof.
[…]
$\frac{1}{n}D(\widetilde{f}^{n})(x)=\frac{1}{n}\sum_{k=0}^{n}D(\widetilde{f})(\widetilde{f}^{k}(x)).$
[…]
The following corollary will be useful for constructing examples.
Corollary 2.5.
For every respective lifts $\widetilde{f},\widetilde{g}$ of $f,g\in\mathrm{Homeo}_{0}(T^{2})$, we have
[…]
Proof.
Let $\varepsilon>0$. Note that $\widetilde{f}$ is at bounded distance from the identity, and thus $\widetilde{f}\widetilde{g}^{n}$ is at bounded distance from $\widetilde{g}^{n}$. By the Misiurewicz-Ziemian theorem above, for every $n$ large enough, the set
$E_{n}=\frac{1}{n}(\widetilde{f}\widetilde{g}^{n}-\mathrm{Id})([0,1]^{2})$
is included in the $\varepsilon$-neighborhood $N_{\varepsilon}$ of $\mathrm{Rot}(g)$. Consider such an integer $n$. Since the rotation set is convex, $N_{\varepsilon}$ is also convex, and thus it also contains $\mathrm{Conv}(E_{n})$. We apply the lemma to $fg^{n}$ to get
$\frac{1}{n}\mathrm{Rot}(\widetilde{f}\widetilde{g}^{n})\subseteq\mathrm{Conv}(E_{n})\subseteq N_{\varepsilon}.$
Unfortunately we do not know if the equality holds in the corollary.
Question 2.6.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ and assume their rotation sets have non empty interior. Do we always have that
$\lim\frac{1}{n}\mathrm{Rot}(fg^{n})=\mathrm{Rot}(g)?$
The difficulty to answer this question positively comes from the lack of reasonable lower bounds. This also explains why we could only prove semi-continuity, and not continuity, in our last theorem (see section 10.3 and the question inside).
The next lemma will be useful to control the rotation sets in the main proofs. Remember from the introduction that a sequence in $\mathrm{Homeo}_{0}(T^{2})$ is said to be bounded if it is bounded in the sense of Mann-Rosendal, or equivalently, bounded for the distance $\widetilde{d}$ defined as the minimal $C^{0}$ distance between lifts to the plane $\mathbb{R}^{2}$. The next lemma shows how such boundedness may provide information on the rotation set up to translation. We emphasize that the translations in this statement are by any vector in $\mathbb{R}^{2}$, and not necessarily integer vectors.
Lemma 2.7 (Boundedness and the rotation set).
  1. i)
    Let $f\in\mathrm{Homeo}_{0}(T^{2})$. Let $\varepsilon>0$, $D>0$. Then there is $n_{0}>0$ such that for every $n\geq n_{0}$, for every $g\in\mathrm{Homeo}_{0}(T^{2})$ such that $\widetilde{d}(f^{n},g)\leq D$, the set $\frac{1}{n}\mathrm{Rot}(g)$ has a translate included in the $\varepsilon$-neighborhood of $\mathrm{Rot}(f)$.
  2. ii)
    Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$. Assume that there is a sequence $(m_{n})_{n\geq 0}$ in $\mathbb{N}$ such that the sequence $(f^{n}g^{-m_{n}})_{n\geq 0}$ is bounded. Assume furthermore that $\lim\frac{m_{n}}{n}$ exists and is a positive number, and denote it $\lambda$. Then $\lambda\mathrm{Rot}(g)$ and $\mathrm{Rot}(f)$ are translates of each other.
Proof of Lemma 2.7.
Let us prove i). Choose some lifts $\widetilde{f}$ of $f$, and let $n_{0}$ be large enough so that $\frac{D}{n_{0}}<\varepsilon$ and for every $n\geq n_{0}$ the set
$\frac{1}{n}({\widetilde{f}}^{n}-\mathrm{Id})([0,1]^{2})$
[…]
$\frac{1}{n}\mathrm{Rot}(\widetilde{g})\subset\mathrm{Conv}\left(\frac{1}{n}(\widetilde{g}-\mathrm{Id})([0,1]^{2})\right).$
By the distance estimate, this last set is included in the $\frac{D}{n}$-neighborhood of
$\mathrm{Conv}\left(\frac{1}{n}(\widetilde{f}^{n}-\mathrm{Id})([0,1]^{2})\right)$
[…]
Now let us prove ii). Denote $D$ the bound for the $\widetilde{d}-$distance from $(f^{n}g^{-m_{n}})_{n\geq 0}$ to the identity in $\mathrm{Homeo}_{0}(T^{2})$. Consider some $\varepsilon>0$. Let $n_{0}$ be given by i): for every $n\geq n_{0}$, since $\widetilde{d}(f^{n},g^{m_{n}})\leq D$, we get that $\frac{1}{n}\mathrm{Rot}(g^{m_{n}})=\frac{m_{n}}{n}\mathrm{Rot}(g)$ has a translate included in the $\varepsilon$-neighborhood of $\mathrm{Rot}(f)$. Letting $n$ tend to $\infty$, and then $\varepsilon$ to 0, we get that $\lambda\mathrm{Rot}(g)$ has a translate included in $\mathrm{Rot}(f)$. For the reverse inclusion, we apply i) with the roles of $f$ and $g$ exchanged. We get some $n_{0}$ so that, for every $n\geq n_{0}$, $\frac{1}{m_{n}}\mathrm{Rot}(f^{n})=\frac{n}{m_{n}}\mathrm{Rot}(f)$ has a translate included in the $\varepsilon$-neighborhood of $\mathrm{Rot}(g)$. We conclude as above that $\frac{1}{\lambda}\mathrm{Rot}(f)$ has a translate included in $\mathrm{Rot}(g)$. ∎
3. Hyperbolic preliminaries
In this section, we collect some preliminaries on Gromov hyperbolic spaces. We refer the reader to e.g. [BH99] for details on hyperbolic spaces; we will only collect the required results. Most of them are well-known, but especially in Subsections 3.2 or 3.3 we discuss results specifically adapted to our situation.
3.1. Basic definitions and the boundary
Recall that a geodesic metric space $X$ is called $\delta$–hyperbolic if every geodesic triangle is $\delta$-thin, i.e. if every side is contained in the $\delta$-neighbourhood of the other two sides.
For a basepoint $b\in X$, the Gromov product is defined as
$(x\cdot y)_{b}=\frac{1}{2}\left(d(b,x)+d(b,y)-d(x,y)\right).$
If $X$ is $\delta$–hyperbolic, and $\alpha$ is a geodesic joining $x$ to $y$, then
$(x\cdot y)_{b}\leq d(b,\alpha)\leq(x\cdot y)_{b}+3\delta.$
One can also think of the Gromov product as the distance where geodesics from $b$ to $x,y$ start to diverge.
Using the Gromov product, we can define the boundary at infinity of a hyperbolic space. Namely, call a sequence $(x_{i})$ admissible if
$(x_{i}\cdot x_{j})_{b}\to\infty,$
and two sequences $(x_{i}),(y_{i})$ equivalent if
$(x_{i}\cdot y_{j})_{b}\to\infty.$
Then the Gromov boundary $\partial_{\infty}X$ is the set of equivalence classes of admissible sequences. An admissible sequence $(x_{i})$ defining the boundary point $\xi$ is said to converge to $\xi$.
If $X$ is proper, then $\partial_{\infty}X$ can be identified with the set of geodesic rays up to bounded distance – but since the spaces we deal with are very much not proper, it is not clear if every boundary point is reached by an actual geodesic ray in our setting. See e.g. [LT25, Section 2.A] for discussion.
3.2. Geodesics and quasigeodesics
A central property of geodesics in hyperbolic spaces is the following stability (see e.g. [BH99, Chapter III.H, Theorem 1.7])
Lemma (Morse lemma).
Suppose $X$ is a $\delta$–hyperbolic space. Then for any $K>0$ there is a $B=B(K,\delta)$ with the following property. Suppose that $\gamma$ a (finite) geodesic, and $q$ is a KK–quasigeodesic with the same endpoints. Then the Hausdorff distance between $\gamma,q$ is at most $B$.
Note that the Morse lemma implies that two quasigeodesics with the same endpoints are bounded Hausdorff distance apart (with a bound depending only on the hyperbolicity and quasigeodesic constants).
Another result we will use several times is the following local-to-global property for quasigeodesics. Variants of this result are well-known; see e.g. [Min05, Lemma 4.2] for a version immediately implying our required result. We adopt the usual convention that $[xy]$ denotes any geodesic joining two given points $x$ and $y$.
Lemma 3.1 (Local-to-global).
Let $X$ be a $\delta$-hyperbolic space. For every $B$ there are constants $L,K$ so that the following is true: if
$g=[x_{1}x_{2}]\ast\cdots\ast[x_{k}x_{k+1}]$
is a concatenation of geodesics, so that for every $1\leq i\leq k$ we have
  1. (1)
    $d(x_{i},x_{i+1})\geq L$, and
  2. (2)
    $(x_{i},x_{i+2})_{x_{i+1}}\leq B$,
then $g$ is a KK–quasigeodesic.
Lemma 3.2 (Fellow traveling geodesics).
Let $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ be four points in a $\delta$-hyperbolic space, and let $b$ be such that
$d(\alpha,\beta)\leq b,\quad d(\alpha^{\prime},\beta^{\prime})\leq b.$
Let $\ell>0$, and $\alpha_{\ell},\beta_{\ell}$ respectively denote points of the geodesics $\alpha,\alpha^{\prime}]$, $\beta,\beta^{\prime}]$ that are distance $\ell$ from $\alpha,\beta$. Then $d(\alpha_{\ell},\beta_{\ell})\leq 5b+4\delta$.
Proof.
$\alpha$
$\beta$
$\alpha^{\prime}$
$\beta^{\prime}$
$\alpha_{\ell}$
$\beta_{\ell}$
$\beta_{\overline{\ell}}$
Figure 1. Lemma 3.2
Note that the geodesics $\alpha,\alpha^{\prime}]$, $\beta,\beta^{\prime}]$ have $b$-close end-points, so their length differ by no more than $2b$. If $\ell\leq 2\delta+b$, then the lemma follows at once from the triangular inequality. Similarly, we also get the result if $d(\alpha_{\ell},\alpha^{\prime})\leq 2\delta+b$, since in this case the above remark on the length of the geodesics gives $d(\beta_{\ell},\beta^{\prime})\leq 2\delta+3b$.
From now on we assume that the distance from $\alpha_{\ell}$ to both endpoints $\alpha,\alpha^{\prime}$ is more than $2\delta+b$. Consider the geodesic quadrilateral with vertices $\alpha,\beta,\beta^{\prime},\alpha^{\prime}$. Quadrilaterals in a $\delta$-hyperbolic space are $2\delta$ thin. Thus the point $\alpha_{\ell}\in[\alpha,\alpha^{\prime}]$ is $2\delta$ close to some point in the union of the three other sides. Since $\alpha_{\ell}$ is not too close to $\alpha$ neither to $\alpha^{\prime}$, this point has to be on the geodesic $\beta,\beta^{\prime}]$. Let us denote it $\beta_{\bar{\ell}}$, with $\bar{\ell}=d(\beta,\beta_{\bar{\ell}})$; and we remember that $d(\alpha_{\ell},\beta_{\bar{\ell}})\leq 2\delta$. From the bound on end-points of the geodesics $\alpha,\alpha_{\ell}]$ and $\beta,\beta_{\bar{\ell}}]$, we get that $\ell$ and $\bar{\ell}$ differ by no more than $b+2\delta$. Thus $d(\beta_{\bar{\ell}},\beta_{\ell})\leq b+2\delta$, and now the triangular inequality gives $d(\alpha_{\ell},\beta_{\ell})\leq b+4\delta$. ∎
Lemma 3.3 (Fellow traveling quasigeodesics).
Let $b>0$. Let $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ be four points in a $\delta$-hyperbolic space, and assume that
$d(\alpha,\beta)\leq b,\quad d(\alpha^{\prime},\beta^{\prime})\leq b.$
[…]
Proof.
$\alpha$
$\beta$
$\alpha^{\prime}$
$\beta^{\prime}$
$\alpha_{\overline{k}}$
$\beta_{\overline{\ell}}$
$\hat{\alpha}$
$\hat{\beta}$
Figure 2. Lemma 3.3
Consider a geodesic $\alpha,\alpha^{\prime}]$. By the Morse lemma, the point $\hat{\alpha}$ is distance at most $M$ from some point $\alpha_{\bar{k}}$ on this geodesic, where $\bar{k}=d(\alpha,\alpha_{\bar{k}})$. Let $\beta_{\bar{k}}$ denote the point on a geodesic $\beta,\beta^{\prime}]$ which is distance $\bar{k}$ from $\beta$. The previous lemma tells us that $d(\alpha_{\bar{k}},\beta_{\bar{k}})\leq 5b+4\delta$. On the other hand, we may also apply the Morse lemma to the point $\hat{\beta}$ and get some point $\beta_{\bar{\ell}}$ on the geodesic $\beta,\beta^{\prime}]$, which is distance at most $M$ from $\hat{\beta}$, where $\bar{\ell}=d(\beta,\beta_{\bar{\ell}})$. Note that both points $\beta_{\bar{k}},\beta_{\bar{\ell}}$ lies on the geodesic $\beta,\beta^{\prime}]$, and thus $d(\beta_{\bar{k}},\beta_{\bar{\ell}})=|\bar{\ell}-\bar{k}|$. Furthermore $\bar{k}$ differ from $k$ by no more than $M$, and similarly for $\bar{\ell}$ and $\ell$, and thus $|\bar{\ell}-\bar{k}|\leq|\ell-k|+2M$. Now the result follows from the triangular inequality by relating $\hat{\alpha}$ to $\hat{\beta}$ successively through the points $\alpha_{\bar{k}},\beta_{\bar{k}},\beta_{\bar{\ell}}$. ∎
Corollary 3.4.
For every $\delta,K,b$ there is $M^{\prime}$ such that the following holds. Let $f,g$ be two isometries having invariant quasi-axes which are $K$-quasigeodesic paths respectively through points $\alpha,\beta$. Let $n,m$ be positive integers such that
$d(\alpha,\beta)\leq b,\quad d(f^{2n}\alpha,g^{2m}\beta)\leq b.$
Then $d(f^{n}\alpha,g^{m}\beta)\leq M^{\prime}$.
$\alpha$
$\beta$
$f^{2n}\alpha$
$g^{2n}\beta$
$f^{n}\alpha$
$g^{n}\beta$
Figure 3. Corollary 3.4
Proof.
Denote $M=M(K)$ the Morse constant. Let $k=d(\alpha,f^{n}\alpha)=d(f^{n}\alpha,f^{2n}\alpha)$, $\ell=d(\beta,g^{m}\beta)=d(g^{m}\beta,g^{2m}\beta)$. By the distance hypotheses, the lengths of the two geodesics $\alpha,f^{2n}\alpha]$, $\beta,g^{2m}\beta]$ differ by at most $2b$. Using this estimate and the closest point projection of $f^{n}\alpha$ on $\alpha,f^{2n}\alpha]$, and likewise for $\beta$, we get $|\ell-k|\leq 2M+b$. Now we may apply the previous lemma to get the wanted estimate, with $M^{\prime}=5b+4\delta+4M+(2M+b)$. ∎
We next exploit the fact that in hyperbolic space (quasi-)geodesics fellow-travel uniformly close before diverging. The toy example is the following situation. Suppose that $g_{1},g_{2}$ are two geodesics starting in a common point $p$. Now suppose that the geodesic $\alpha$ between $g_{1}(s),g_{2}(t)$ stays distance at least $L$ from $p$. Consider the first point $g_{1}(r)$ on $g_{1}$ which is not $\delta$–close to $g_{2}$. Such a point is (by $\delta$–hyperbolicity) $\delta$–close to the geodesic $\alpha$ and thus $r\geq L-\delta$. In other words, initial segments of $g_{1},g_{2}$ of length $L-\delta$ are $\delta$–fellow-traveling. In particular we have that $d(g_{1}(s),g_{2}(s))\leq 2\delta$ for all $s\leq L-\delta$.
The next lemma is a somewhat technical fellow-traveling statement for quasigeodesics generalizing this situation.
$g_{1}$
$g_{2}$
$g_{1}(s)$
$g_{1}(t)$
$\delta$
Figure 4. The motivating example for Lemma 3.5
Lemma 3.5.
Fix some $\delta,K,b_{0}$. There exists $b=b(K,\delta,b_{0})$ such that the following holds. For every $N$, there exists $L$ such that for every $K$-quasigeodesics $(\alpha_{i})_{i=0,\dots,i_{0}}$, $(\beta_{i})_{i=0,\dots,i_{0}}$, if
  1. (1)
    $d(\alpha_{0},\beta_{0})\leq b_{0}$,
  2. (2)
    $d(\alpha_{0},[\alpha_{i_{0}}\beta_{i_{0}}])\geq L$
then $N\leq i_{0}$, and there exists $N^{\prime}\leq i_{0}$ such that $d(\alpha_{N},\beta_{N^{\prime}})\leq b$.
Proof.
Let us fix $\delta,K,b_{0},N$, and consider $K$-quasigeodesics $(\alpha_{i})_{i=0,\dots,i_{0}}$, $(\beta_{i})_{i=0,\dots,i_{0}}$ with $d(\alpha_{0},\beta_{0})\leq b_{0}$. Extend $\beta_{i}$ by setting $\beta_{-1}=\alpha_{0}$. The resulting path is a $K^{\prime}$–quasigeodesic for a $K^{\prime}$ depending only on $K$ and $b_{0}$. Let $M=M(\delta,K,b_{0})$ be the Morse constant for $K^{\prime}$-quasigeodesic. Hence, $(\alpha_{i}),(\beta_{i}),i\leq i_{0}$ are $M$-close to geodesics $g_{1},g_{2}$ joining $\alpha_{0}$ to $\alpha_{i_{0}},\beta_{i_{0}}$. There is a constant $\hat{N}$ so that any $\alpha_{i},\beta_{i}$ with $i\leq N$ is $M$-close to a point on $g_{j}$ which has distance at most $\hat{N}$ from $\alpha_{0}$.
Choosing $L$ large enough, we can guarantee (as in the discussion before the lemma) that initial segments of length $\hat{N}$ of $g_{1},g_{2}$ stay $2\delta$–close. This also implies that any $\alpha_{N}$ is $2\delta+2M$ from some $\beta_{N^{\prime}}$. Thus $b=2\delta+2M$ has the desired property. ∎
3.3. Isometry types and translation length
We next turn to isometries of a $\delta$-hyperbolic space. The key quantity here is the (asymptotic) translation length of an isometry $\phi$:
$\mathrm{tl}(\phi)=\lim_{n\to\infty}\frac{d(x_{0},\phi^{n}x_{0})}{n}$
for some basepoint $x_{0}$. It is immediate that $\mathrm{tl}(\phi)$ does not depend on the choice of $x_{0}$. We then say that
  1. (1)
    $\phi$ is loxodromic (or hyperbolic) if $\mathrm{tl}(\phi)>0$.
  2. (2)
    $\phi$ is parabolic if $\mathrm{tl}(\phi)=0$, but orbits $\{\phi^{n}(x_{0}),n\in\mathbb{N}\}$ have infinite diameter, and
  3. (3)
    $\phi$ is elliptic if orbits $\{\phi^{n}(x_{0}),n\in\mathbb{N}\}$ have finite diameter.
If $\phi$ is loxodromic, then there is a bi-infinite quasigeodesic invariant under $\phi$ – in fact, one can simply take the images $\phi^{n}(x_{0}),n\in\mathbb{Z}$. We call any invariant quasigeodesic a (quasi)axis for $\phi$. Every isometry extends continuously to the Gromov boundary $\partial_{\infty}X$, and the extension of a loxodromic isometry $\phi$ has exactly two fixed points, denoted $\xi^{\pm}\phi$, such that every orbit $(\phi^{n}x)$ converges to $\xi^{-}\phi$ when $n$ tends to -$\infty$, and to $\xi^{+}\phi$ when $n$ tends to $\infty$.
Observe that there is no reason to expect that $\phi$ has an invariant geodesic axis. In fact, if the space $X$ is a graph, this would imply that $\mathrm{tl}(\phi)$ would be an integer, since $\phi$ would translate along this axis by an integral amount.
The following lemma shows that we can, however, find a quasi-axis of definite quality if the translation length is not too small. It is likely known to experts, but we were not able to find a discussion in the literature, so we include a proof.
In our application of the following lemma, the space $X$ will be a graph, where the minimizing point $x$ needed in the first point always exists. The hypothesis on $x$ may be weakened by only requiring that $x$ is a $\delta$-almost minimizing point, and this gives a meaningful statement in the general case. Details are left to the reader.
Lemma 3.6.
Let $X$ be a $\delta$-hyperbolic metric space $X$.
  1. (1)
    There are $K=K(\delta),L=L(\delta)$ such that the following holds. Let $f$ be a loxodromic element with $\mathrm{tl}(f)\geq L$. Choose $x\in X$ which minimizes $d(x,f(x))$, and let $[x,f(x)]$ denote any geodesic path from $x$ to $f(x)$. Then
    $\bigcup_{n\in\mathbb{Z}}f^{n}([x,f(x)])$
    is a $K$-quasigeodesic.
  2. (2)
    Let $K>0$, and assume $x$ is on a $K$-quasigeodesic which is invariant under $f$. Then for every positive $p,n$ we have
    $n\left(d(x,f^{p}(x))-2M\right)\leq d(x,f^{np}(x))$
    where $M=M(\delta,K)$ is the Morse constant. In particular,
    $\frac{1}{p}\left(d(x,f^{p}(x)\right)-2M)\leq\mathrm{tl}(f).$
Note that by sub-additivity we always have, for every $x\in X$ and $p>0$,
$\mathrm{tl}(f)\leq\frac{1}{p}(d(x,f^{p}(x))$
and together with the last part of the lemma, this gives the precise convergence speed towards the translation length, in terms of the quality of a quasigeodesic axis containing $x$.
Proof.
  1. (1)
    Assuming that $d(x,f(x))>14\delta$, we first show that
    $(f^{-1}x\cdot f(x))_{x}<4\delta.$
    Namely, assume that this would be false. Consider a geodesic triangle with corners $f^{-1}(x),x,f(x)$. By the assumption that $(f^{-1}x\cdot f(x))_{x}\geq 4\delta$, the sides $g=[x,f^{-1}(x)],h=[x,f(x)]$ are $2\delta$–fellow-traveling for initial subsegments of length $4\delta$. Let $y\in g$ be the endpoint of such a segment. Note that it is $2\delta$–close to a point $z_{1}\in h$ (which thus has $d(z_{1},x)\leq 6\delta$).
    $x$
    $f(x)$
    $f^{-1}(x)$
    $y$
    $f(y)$
    $z_{1}$
    $z_{2}$
    $g$
    $f(g)$
    $4\delta$
    Figure 5. The proof of Lemma 3.6 (1). The geodesic $f(g)$ is drawn dashed.
    The image $f(y)$ lies on the geodesic $f(g)$ which joins $f(x)$ to $x$, and is thus $\delta$–close to a point $z_{2}$ on $h$, and we have $d(z_{2},f(x))\leq 5\delta$. By our assumption that $d(x,f(x))>14\delta$, the point $z_{1}$ is between $x$ and $z_{2}$ on $h$, and since we have $d(z_{1},x)\geq 2\delta,d(z_{2},f(x))\geq 3\delta)$ we get
    $d(z_{1},z_{2})\leq d(x,f(x))-5\delta,$
    implying
    $d(y,f(y))\leq d(x,f(x))-5\delta+3\delta<d(x,f(x)),$
    which contradicts minimality of $x$.
    Given the estimate on $(f^{-1}x\cdot f(x))_{x}$, we apply the local-to-global Lemma 3.1 to $B=4\delta$, we obtain constants $K,L_{0}$. Setting $L=\max(L_{0},14\delta)$ this lemma then implies that
    $\bigcup_{n\in\mathbb{Z}}f^{n}([x,f(x)])$
    is a $K$-quasigeodesic.
  2. (2)
    Let $n,p>0$, and let us prove that $d(x,f^{np}(x))\geq n(d(x,f^{p}(x))-2M)$. By the Morse Lemma, there is a point $\pi$ on the geodesics $x,f^{np}x]$ which is distance less than $M$ from $f^{p}(x)$. We get
    $d(x,f^{np}(x))$ $=$ $d(x,\pi)+d(\pi,f^{np}(x))$
    $\geq$ $(d(x,f^{p}(x))-M)+(d(f^{p}(x),f^{np}(x))-M)$
    $=$ $d(x,f^{(n-1)p}(x))+d(x,f^{p}(x))-2M$
    and now the estimate follows by induction on $n$.
    Dividing by $np$ and letting $n$ tends to $\infty$ yields the estimate on the translation length.
Corollary 3.7.
For every $\delta,K,T$ there is some $B=B(\delta,K,T)$ with the following property. If $(\beta_{i})$ is a $K$-quasigeodesic which is invariant under the isometry $g$, with $\mathrm{tl}(g)\leq T$, then every $\beta_{i}$ is distance at most $B$ from some iterate $g^{n}(\beta_{0})$.
Proof.
Applying point (2) of Lemma 3.6 with $p=1$, we get $d(\beta_{0},g(\beta_{0}))\leq T+2M(\delta,K)$. Let $i_{0}$ be the index such that $\beta_{i_{0}}=g(\beta_{0})$. The $K$-quasi-geodesicity gives a bound on $i_{0}$ by some function of $K$ and $d(\beta_{0},g(\beta_{0}))$. Then the distance from any $\beta_{i}$ to the closest $\beta_{ni_{0}}=g^{n}(\beta_{0})$ is also bounded. ∎
3.4. Bestvina-Fujiwara equivalence relation
In this section we (very) briefly summarize results which allow to use actions on hyperbolic spaces to construct unbounded quasimorphisms on groups, or alternatively certify that elements have positive stable commutator length (see the introduction for the definition).
Let $G$ be a group acting on a $\delta$-hyperbolic space, and $f,g\in G$. Assume they act loxodromically, and denote $\xi^{\pm}(f),\xi^{\pm}(g)$ their end-points. Bestvina-Fujiwara ([BF02]) define a relation $f\sim g$ if there are copies of the axis of $g$ arbitrarily close to the axis of $f$. More precisely: if for every neighborhoods $V^{-},V^{+}$ of $\xi^{-}(f),\xi^{+}(f)$ there exists $h\in G$ such that $h(\xi^{-}(g))\in V^{-},h(\xi^{+}(g))\in V^{+}$.
We can now state the criterion:
Theorem ([BF02, Proposition 5]).
Suppose $G$ acts on a $\delta$–hyperbolic space by isometries. Assume $f$ acts loxodromically and $f\not\sim f^{-1}$. Then $\mathrm{scl}(f)>0$.
Let us end this section with an easy criterium for Bestvina-Fujiwara equivalence. Let $f,g$ be acting loxodromically as above, and assume $\xi^{+}(f)=\xi^{+}(g)$. Then for $n$ large enough $f^{-n}$ fixes $\xi^{+}(g)$ and sends $\xi^{-}(g)$ arbitrarily close to $\xi^{-}(f)$. We deduce:
Lemma 3.8.
If $\xi^{+}(f)=\xi^{+}(g)$ then $f\sim g$.
As an immediate consequence, we see that Theorem 4 (the Bestvina-Fujiwara class determines the scaled rotation set) entails Theorem 3 (the fixed point determines the scaled rotation set).
4. Fine curve graphs
In this section, we introduce the second main object of this article, the fine curve graph $\mathcal{C}^{\dagger}(S)$ of a closed, connected, oriented surface $S$. The vertices of $\mathcal{C}^{\dagger}(S)$ correspond to simple, essential curves on $S$. If $S$ has genus $g>1$, we join two vertices by an edge if the corresponding curves are disjoint. If $S$ is the torus, we also join vertices corresponding to curves which intersect transversely, and at most once. Note that there are by now many variants of the fine curve graph being studied which have slightly different edge relations (but all of which are quasi-isometric).
The first key property is the following.
Theorem 4.1 ([BHW22, Theorem 3.10]).
The fine curve graph is Gromov hyperbolic.
An important ingredient of the proof, that we will use again below, are bicorn surgery paths. Given simple closed curves $\alpha,\beta$, a bicorn defined by $\alpha$ and $\beta$ is a simple, essential closed curve made of an arc $a$ included in $\alpha$ and an arc $b$ included in $\beta$. Observe that not every subarc of $\alpha$ defines a bicorn. However, for every pair of transverse curves $\alpha,\beta$ one can find a quasi-path from $\alpha$ to $\beta$ in $\mathcal{C}^{\dagger}(S)$ made of bicorns defined by $\alpha$ and $\beta$, where the subarcs taken from $\alpha$ are strictly decreasing. We call such quasi-paths bicorn surgery paths. We warn the reader that, strictly speaking, these are not paths in the fine curve graph (but they are paths in the augmented graphs, where edges are put between curves intersecting at most $4$ times).
The key point is that bicorn surgery paths are geometrically efficient. This follows from work of Rasmussen [Ras20] on nonseparating curve graphs; see [BHW24, Section 4] for details.
Lemma 4.2 ([BHW24, Lemma 4.2]).
There is a constant $K>0$ with the following property: suppose $\alpha,\beta\subset S$ are two curves intersecting transversely. Then if $(a_{i})$ is a bicorn surgery path between $\alpha$ and $\beta$, there is a reparametrization $i(n)$ so that $n\mapsto a_{i(n)}$ is a KK–quasigeodesic.
The following lemma will allow us to consider bicorns for curves that do not intersect transversely. Here we consider vertices of $\mathcal{C}^{\dagger}(S)$ as continuous maps $\mathbb{S}^{1}\to S$ up to reparametrization, which allows to endow the set of curves with the quotient topology from the topology of uniform convergence of maps. (Note that this is stronger than the Hausdorff distance between sets; in the lemma the Hausdorff distance would be enough for us.)
Lemma 4.3 (Bicorn Perturbation Lemma).
Let $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$, and subarcs $a\subset\alpha,b\subset\beta$ such that $a\cup b$ is a bicorn defined by $\alpha$ and $\beta$. For every $\varepsilon>0$, there exists curves $\alpha^{\prime},\beta^{\prime}\in\mathcal{C}^{\dagger}(S)$ which intersect transversely, are respectively $\varepsilon$-close to $\alpha$ and $\beta$, and subarcs $a^{\prime}\subset\alpha^{\prime},b^{\prime}\subset\beta^{\prime}$ such that $a^{\prime}\cup b^{\prime}$ is a bicorn defined by $\alpha^{\prime}$ and $\beta^{\prime}$ which is $\varepsilon$-close to $a\cup b$. Furthermore we can assume that $\alpha^{\prime}=\alpha$.
We postpone the proof of the perturbation lemma to the end of this section. We will need the following consequence.
Corollary 4.4.
There is a constant $M>0$ such that any bicorn defined by any pair $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$ is within distance $M$ of any geodesic $\alpha,\beta]$.
Proof of the Corollary.
Let $M$ be the Morse constant associated to $K$-quasi-geodesics for the constant $K$ of Lemma 4.2. Let $a\cup b$ be a bicorn defined by curves $\alpha$ and $\beta$, and apply the Bicorn Perturbation Lemma to get curves $\alpha^{\prime},\beta^{\prime}$ intersecting transversely, arbitrarily close to $\alpha,\beta$, defining a bicorn $a^{\prime}\cup b^{\prime}$ arbitrarily close to $a\cup b$. Curves that are close have a common disjoint curve, and thus are within distance $2$ in the fine curve graph, thus we get
$d(\alpha,\alpha^{\prime})\leq 2,\ \ d(\beta,\beta^{\prime})\leq 2,\ \ d(a\cup b,a^{\prime}\cup b^{\prime})\leq 2.$
By Lemma 4.2 and the Morse Lemma, $a^{\prime}\cup b^{\prime}$ is within distance $M$ of any geodesic $\alpha^{\prime},\beta^{\prime}]$. On the other hand, by hyperbolicity, the geodesic $\alpha^{\prime},\beta^{\prime}]$ is included in the $(\delta+4)$-neighborhood of the geodesic $\alpha,\beta]$. Thus the corollary holds with constant $M+\delta+6$. ∎
We will also make use of the fine arc graph of the annulus $A=[0,1]\times S^{1}$, denoted $\mathcal{A}^{\dagger}(A)$, whose vertices are essential proper arcs in $A$, that is, images of $[0,1]$ under an embedding that maps $(0,1)$ into the interior of $A$ and 0 and $1$ to distinct boundary components of $A$, and edges correspond to disjointness. The distance between two vertices $a\neq b$ in this graph may be computed as follows (see [BHM+22, LRW24]).
If $a,b$ are arcs in $A$, let $\widetilde{a},\widetilde{b}$ be lifts of the two arcs to the universal cover of $A$. Define the width $\mathrm{wd}_{A}(a,b)$ of $a$ and $b$ in $A$ as the number of distinct lifts of $a$ which intersect one lift of $b$. We then have
Lemma 4.5.
For any arcs $a,b\subset A$ we have
$d_{\mathcal{A}^{\dagger}(A)}(a,b)=\mathrm{wd}_{A}(a,b)+1.$
In particular, the distance between two arcs that intersect exactly once is equal to $2$, which is just the observation that two essential proper arcs that intersect once have a common disjoint arc.
4.1. Proof of the perturbation lemma
We start with a short discussion of intersection points for general $C^{0}$ curves. Let $\alpha,\beta$ be two simple closed curves on our surface $S$. Call a point $z\in\alpha\cap\beta$ an unstable intersection point if there exist a neighborhood $V$ of $z$, and sequences of simple closed curves $(\alpha_{n}),(\beta_{n})$ converging to $\alpha,\beta$ such that for each $n$, $\alpha_{n}\cap\beta_{n}\cap V=\emptyset$. Here is a more concrete description of unstable intersection points. Assume $\alpha$ and $\beta$ meet at $z$. Consider coordinates $(x,y)$ around the intersection point $z$ such that $\alpha$ locally coincides with the line $y=0$. Then $z$ is an unstable intersection point if and only if there is an open subarc of $\beta$ containing $z$ which is included in $y\geq 0$ or in $y\leq 0$. The following lemma will allow us to replace unstable intersection points with stable ones (where stable means not unstable).
Lemma 4.6.
Let $\alpha,\beta$ be two simple closed curves on $S$, and $z\in\alpha\cap\beta$. Let $\varepsilon>0$. Then there exists a simple closed curve $\beta^{\prime}$, $\varepsilon$-close to $\beta$, equal to $\beta$ outside the $\varepsilon$-neighborhood of $z$ and such that $z$ is a stable intersection point of $\alpha$ and $\beta^{\prime}$.
Proof.
We may obviously assume that the intersection point $z$ is unstable. We consider local coordinates as above, and assume that $\beta$ is locally contained in the upper half-plane $y\geq 0$. The new curve $\beta^{\prime}$ is obtained as the image of $\beta$ under the composition $h_{2}h_{1}$ of two homeomorphisms with support arbitrarily close to $x$. Namely, we first apply $h_{1}$ to push $\beta$ vertically downwards, so that $h_{1}\beta$ meets $y<0$ while still having points in $y>0$. It is easy to see that $h_{1}\beta$ has a stable intersection point $z^{\prime}$ with $\alpha$ in the local chart. Now push $z^{\prime}$ horizontally by $h_{2}$ that preserves $\alpha$ and sends $z^{\prime}$ to $z$, so that $z$ is a stable intersection point of $\alpha$ and $h_{2}h_{1}\beta$. ∎
Proof of Lemma 4.3.
Let $\alpha,\beta,a,b,\varepsilon$ as in the lemma. Denote $x,y$ the common endpoints of $a$ and $b$. The first step is to get stable intersection points. By applying the above lemma independently in some small neighborhoods of $x$ and $y$, we may perturb $\beta$ into a curve $\beta^{\prime}$ which have stable intersection points with $\alpha$ at $x$ and $y$. Let $b^{\prime}$ be the subarc of $\beta^{\prime}$ with endpoints $x,y$ which is close to $b$. The closed curve $a\cup b^{\prime}$ may be not simple, but $a\cap b^{\prime}$ is included in the union of the neighborhoods of $x$ and $y$ on which the perturbation took place. Thus we may shorten $a$ and $b^{\prime}$ a little bit to get a simple closed curve $a^{\prime}\cup b^{\prime\prime}$ which is a bicorn defined by $\alpha$ and $\beta^{\prime}$, and which is still close to $a\cup b$. Note that the closeness ensures that $a^{\prime}\cup b^{\prime\prime}$ is homotopic to $a\cup b$, and thus is still an essential closed curve. The conclusion of this first step is that it suffices to prove the lemma in the case when the intersection points $x$ and $y$ are stable, which we assume now. We endow $S$ with a smooth structure. Thanks to the Schoenflies theorem, we may assume that $\alpha$ is smooth with respect to this structure. It is well known that smooth curves are $C^{0}$-dense in the set of continuous simple closed curves. Furthermore, among smooth curves, transversality to a given one is a generic property. Thus we may find $\beta^{\prime}$, arbitrarily $C^{0}$-close to $\beta$, which is smooth and transverse to $\alpha$. Since $x$ and $y$ were stable intersection points, $\alpha\cap\beta^{\prime}$ contains points $x^{\prime},y^{\prime}$ respectively arbitrarily close to $x$ and $y$. As above, we can now find a bicorn $a^{\prime}\cup b^{\prime}$ defined by $\alpha$ and $\beta^{\prime}$ and arbitrarily close to $a\cup b$. This completes the proof. ∎
5. Fine markings
In this section we restrict to the torus for simplicity. There are likely no serious issues generalising the notions and results discussed here to higher genus; however, we do not want to commit to a precise definition at this point, and we will explore the geometry of marking graphs in higher genus in future work.
Definition 5.1.
  1. i)
    A (fine) marking (of the torus) is a set $\{\alpha,\beta\}$ of simple closed curves which intersect once, and transversely. We say that $\beta$ is a dual to $\alpha$.
  2. ii)
    We say that a marking $\{\alpha^{\prime},\beta^{\prime}\}$ is obtained from $\{\alpha,\beta\}$ by an elementary move if one of the curves is unchanged, say $\alpha=\alpha^{\prime}$, and $\beta^{\prime}$ intersects $\beta$ at most once, and transversely.
  3. iii)
    The fine marking graph $\mathcal{M}^{\dagger}$ is the graph whose vertices are fine markings, and whose edges correspond to elementary moves.
  4. iv)
    If $m$ is a fine marking, we denote by $\mathcal{M}^{\dagger}(m)$ the subgraph spanned by all markings isotopic to $m$.
$\alpha$
$\beta$
$\alpha^{\prime}$
$\beta^{\prime}$
Figure 6. An elementary move of fine markings $\{\alpha,\beta\}$ to $\{\alpha,\beta^{\prime}\}$
Thus the vertices of the fine marking graph are edges of the fine curve graphs, and the edges corresponds to triangles in the fine curve graph. Note that, by classification of surfaces, the group $\mathrm{Homeo}(T^{2})$ acts transitively on the fine marking graph, and that the subgraph $\mathcal{M}^{\dagger}(m)$ is the orbit of $m$ under the subgroup $\mathrm{Homeo}_{0}(T^{2})$. Edges of this subgraph correspond to disjoint moves, that is, there is an edge from $\{\alpha,\beta\}$ to $\{\alpha,\beta^{\prime}\}$ if and only if $\beta$ and $\beta^{\prime}$ are disjoint. The geometry of the subgraph $\mathcal{M}^{\dagger}(m)$ captures the geometry of the homeomorphism group in the following sense.
Lemma 5.2.
For any $m_{0}$, the graph $\mathcal{M}^{\dagger}(m_{0})$ is connected and the orbit map
$\mathrm{Homeo}_{0}(T^{2})\to\mathcal{M}^{\dagger}(m_{0}),\quad f\mapsto f^{-1}(m_{0})$
is a quasi-isometry for the (unique) right-invariant quasi-isometric structure on $\mathrm{Homeo}_{0}(S)$ in the sense of Mann-Rosendal, and the (intrinsic) metric on $\mathcal{M}^{\dagger}(m_{0})$.
Proof of Lemma 5.2.
Let $m_{0}=(\alpha_{0},\beta_{0})$. We choose a covering of the surface by disks $U_{i}$ which are: the complement $U_{1}$ of $\alpha_{0}\cup\beta_{0}$, a small neighbourhood $U_{2}$ of $\alpha_{0}\cap\beta_{0}$, and neighbourhoods $U_{3},U_{4}$ of the arcs $\alpha_{0}\cup\beta_{0}\setminus U_{2}$. By the work of Mann-Rosendal ([MR18]), the geometry of $\mathrm{Homeo}_{0}(T^{2})$ is quasi-isometric to the fragmentation norm induced by the cover $\{U_{i}\}$. We denote this norm by $\|\cdot\|$. We denote the metric on $\mathcal{M}^{\dagger}(m_{0})$ for brevity by $d$. By right-invariance of the metric on $\mathrm{Homeo}_{0}(T^{2})$ and of the action by inverses of maps on $\mathcal{M}^{\dagger}(m_{0})$, it suffices to estimate $d(m_{0},\phi^{-1}m_{0})$ in terms of $\|\phi\|$, which we will do now.
First, observe that if $f$ is supported in some $U_{i}$, then
$d(m_{0},fm_{0})\leq 4,$
since for each possible support $U_{i}$ a marking $m_{i}$ can be obtained from $m_{0}$ by at most two elementary moves, which is disjoint from $U_{i}$ and hence fixed by $f$. If now $\phi\in\mathrm{Homeo}_{0}(T^{2})$ is arbitrary, with $\|\phi\|=n$, we can write
$\phi=f_{1}\cdots f_{n},$
with each $f_{i}$ supported in some $U_{m_{i}}$. We have by the triangle inequality
$d(m_{0},\phi m_{0})\leq\sum_{i=1}^{n}d(f_{1}\cdots f_{i-1}m_{0},f_{1}\cdots f_{i}m_{0}).$
Since the action on $\mathcal{M}^{\dagger}(m_{0})$ is isometric, and by the previous remark, we then have
$\sum_{i=1}^{n}d(f_{1}\cdots f_{i-1}m_{0},f_{1}\cdots f_{i}m_{0})=\sum_{i=1}^{n}d(m_{0},f_{i}m_{0})\leq 4n.$
Hence, we have
$d(m_{0},\phi^{-1}(m_{0}))=d(m_{0},\phi(m_{0}))\leq 4\|\phi\|.$
In particular, this shows that the graph is connected, since the action of $\mathrm{Homeo}_{0}(T^{2})$ on $\mathcal{M}^{\dagger}(m_{0})$ is transitive.
$m_{0}$
$f_{1}m_{0}$
$f_{1}f_{2}m_{0}$
$f_{1}\cdots f_{n}m_{0}$
Figure 7. Moving in the marking graph. The lines between consecutive $f_{1}\cdots f_{i}m_{0},f_{1}\cdots f_{i}m_{0}$ indicate distance at most $4$.
Now, on the other hand, suppose that $m_{1},m_{0}$ differ by a single elementary move. Then there is an element $f\in\mathrm{Homeo}_{0}(T^{2})$ with $f(m_{0})=m_{1}$ and $\|f\|\leq 4$; in fact $f$ can be written as a product of homeomorphisms supported on each $U_{i}$ at most once. Namely, assume that $\alpha_{0}$ is unchanged and $\beta_{0}$ changes to $\beta_{1}$, with $\beta_{0}\cap\beta_{1}=\emptyset$. By a modification in the neighbourhood of the intersection point followed by another one in the neighourhood of $\beta_{0}$, we get a curve $f_{1}f_{0}(\beta_{0})$ which is disjoint from $\beta_{0}$. By applying a homeomorphism $f_{2}$ supported in the neighborhood of $\alpha_{0}$, we get $f_{2}f_{1}f_{0}(\beta_{0})$ which additionaly coincides with $\beta_{1}$ near $\alpha_{0}$. Now a modification on the complement of $\alpha_{0}\cup\beta_{0}$ (i.e. a homeomorphism $f_{3}$ of fragmentation norm $1$) can send $f_{2}f_{1}f_{0}(\beta_{0})$ to $\beta_{1}$. The existence of $f_{2}$ and $f_{3}$ is an application of a Schoenflies-like theorem (see for instance [Sie05] for an account of the classical case, and [Hom53] for a very general case.
$\alpha_{0}$
$\beta_{0}$
$f_{1}f_{0}\beta_{0}$
$f_{2}f_{1}f_{0}\beta_{0}$
$\color[rgb]{1,0,0}f_{3}f_{2}f_{1}f_{0}\beta_{0}=\beta_{1}$
$U_{2}$
$U_{4}$
$U_{2}$
$U_{3}$
$U_{3}$
Figure 8. Getting one move in the marking graph with a homeomorphism of fragmentation norm 4.
Let now, $m^{\prime}=\Psi^{-1}(m_{0})$ be a marking obtained by applying an arbitrary element $\Psi\in\mathrm{Homeo}_{0}(T^{2})$. Denote $n=d(m_{0},m^{\prime})=d(m_{0},\Psi(m_{0}))$, and choose a geodesic
$m_{0},m_{1},\ldots,m_{n}=\Psi(m_{0})$
in the fine marking graph. By the above, there is a homeomorphism $\phi_{1}$ of fragmentation norm at most $4$ with $\phi_{1}(m_{0})=m_{1}$. Assume inductively that we have an element $\phi_{i}$ of fragmentation norm at most $4i$ so that $\phi_{i}(m_{0})=m_{i}$. Then, $\phi^{-1}_{i}(m_{i+1})$ is adjacent to $m_{0}$, and so there is an element $f_{i+1}$ of fragmentation norm at most $4$ so that $f_{i+1}(m_{0})=\phi^{-1}_{i}(m_{i+1})$. Hence, we have
$\phi_{i}f_{i+1}(m_{0})=m_{i+1},$
Hence, $\phi_{i+1}=\phi_{i}f_{i+1}$ satisfies the inductive hypothesis for $i+1$. We then have that
$\Psi^{-1}\phi_{n}(m_{0})=m_{0}.$
We leave it to the reader to check that a homeomorphism that preserves $m_{0}$ (but maybe does not fixes it pointwise) has fragmentation norm at most $5$. Therefore $\|\Psi^{-1}\phi_{n}\|\leq 5$. We thus have
$\|\Psi\|\leq 4n+5=4d(m_{0},\Psi^{-1}(m_{0}))+5.$
From now on, we identify the torus $T^{2}$ with $\mathbb{R}^{2}/\mathbb{Z}^{2}$, and we let $m_{0}$ be the marking defined by the horizontal and vertical curve through 0. For brevity, we write $\mathcal{M}^{\dagger}(m_{0})=\mathcal{M}^{\dagger}_{0}$, since the basepoint marking $m_{0}$ will be fixed throughout.
We then have the following useful consequences, connecting marking distance to $\mathcal{C}^{0}$–distance. Remember from the introduction that $\widetilde{d}$ denotes the distance on $\mathrm{Homeo}_{0}(T^{2})$ obtained by minimizing the usual $C^{0}$-distance on $\mathrm{Homeo}(\mathbb{R}^{2})$ over all pairs of lifts.
Corollary 5.3 ($C^{0}$ bound from markings).
For any $D>0$ there is a number $R>0$ with the following property. Suppose that $h\in\mathrm{Homeo}_{0}(T^{2})$ satisfies
$d_{\mathcal{M}_{0}^{\dagger}}(m_{0},hm_{0})\leq D.$
Then
$\widetilde{d}(\mathrm{id},{h})\leq R.$
Proof.
Cover the torus by finitely many disks $U_{i}$ which admit finite diameter lifts to $\mathbb{R}^{2}$, and let $r$ be the maximal diameter of such lifts. In particular, any homeomorphism supported on a disk $U_{i}$ admits a lift to $\mathbb{R}^{2}$ which has $\mathcal{C}^{0}$–distance at most $r$ from the identity. Then, if $h$ has fragmentation norm at most $n$ (with respect to $U_{i}$), it admits a lift $\widetilde{h}$ of distance at most $rn$ to the identity. Now, the corollary follows immediately from Lemma 5.2. ∎
We note that this corollary proves the easy direction of the following result due to Emmanuel Militon [Mil14]. Here $\mathrm{diam}$ denotes the diameter in the plane.
Theorem 5.4 ([Mil14, Theorem 2.4]).
For any choice of covering $\mathcal{U}$ of $T$ there are constants $C,C^{\prime}$ so that for any $f\in\mathrm{Homeo}_{0}(T)$ we have
$\frac{1}{C}\mathrm{diam}(\widetilde{f}[0,1]^{2})-C^{\prime}\leq\|f\|_{\mathcal{U}}\leq C\mathrm{diam}(\widetilde{f}[0,1]^{2})+C^{\prime}.$
It will be convenient for later use to interpret the distance in $\mathcal{M}_{0}^{\dagger}$ in a slightly different way. Given two homotopic simple closed curves $\delta,\delta^{\prime}$ we define the relative width $\mathrm{wd}_{\delta}(\delta^{\prime})$ as the number of distinct lifts of $\delta$ in $\mathbb{R}^{2}$ which a lift of $\delta^{\prime}$ intersects.
Given two markings $m=\{\alpha,\beta\},m^{\prime}=\{\alpha^{\prime},\beta^{\prime}\}$ in the isotopy class of $m_{0}$, we then define
$d_{w}(m,m^{\prime})=\mathrm{wd}_{\alpha}(\alpha^{\prime})+\mathrm{wd}_{\beta}(\beta^{\prime}).$
Lemma 5.5.
The metrics $d_{\mathcal{M}_{0}^{\dagger}}$ and $d_{w}$ on $\mathcal{M}_{0}^{\dagger}$ are equivalent.
Proof.
Note that both metrics are invariant under the action of $\mathrm{Homeo}_{0}(T)$, and thus it suffices to show that $d(m_{0},m),d_{w}(m_{0},m)$ are comparable for all $m$. To this end we consider the orbit map
$\mathrm{Homeo}_{0}(T)\to\mathcal{M}_{0}^{\dagger}(T),$
noting that we already know that this is a quasi-isometry for the metric $d$. Thus, it suffices to show that it also is a quasi-isometry for the metric $d_{w}$. Let $f\in\mathrm{Homeo}_{0}(T)$, denote $m_{0}=\{\alpha,\beta\}$ and $f(m_{0})=\{\alpha^{\prime},\beta^{\prime}\}$. From definition of relative width, it is clear that
$\mathrm{diam}(\widetilde{f}[0,1]^{2})\geq\max\{\mathrm{wd}_{\alpha}(\alpha^{\prime}),\mathrm{wd}_{\beta}(\beta^{\prime})\}$
Next, note that $\widetilde{f}[0,1]^{2}$ is contained in a (topological) strip bounded by two adjacent lifts of $\widetilde{f}(\alpha)$. By definition of the width, this strip is contained in a straight strip of width $\mathrm{wd}_{\alpha}(\alpha^{\prime})+2$ bounded by lifts of $\alpha$, and analogous for $\beta$. This shows
$\mathrm{diam}(\widetilde{f}[0,1]^{2})\leq\mathrm{wd}_{\alpha}(\alpha^{\prime})+\mathrm{wd}_{\beta}(\beta^{\prime})+4.$
Now, the desired claim follows from Theorem 5.4. ∎
Finally, we discuss the difference between the geometry of the subcomplex $\mathcal{M}^{\dagger}_{0}$ and the whole complex $\mathcal{M}^{\dagger}$.
Lemma 5.6.
The inclusion $\mathcal{M}^{\dagger}_{0}\to\mathcal{M}^{\dagger}$ is exponentially distorted; that is: there is a function $F$ of exponential growth type, so that
$d_{\mathcal{M}^{\dagger}_{0}}(m,m^{\prime})\leq F(d_{\mathcal{M}^{\dagger}}(m,m^{\prime}))$
for all $m,m^{\prime}\in\mathcal{M}^{\dagger}_{0}$, and every function with this property has at least exponential growth type.
Proof.
Denote by $R_{0}=[0,1]^{2}$ the fundamental domain corresponding to our basepoint marking $m_{0}$. Observe that there is a $k>0$ so that if $m$ is an arbitrary marking homotopic to $m_{0}$, and $R$ the corresponding fundamental domain, then we have
$\mathrm{diam}(R)\leq 3^{d_{\mathcal{M}^{\dagger}}(m,m_{0})}.$
This follows since each edge in $\mathcal{M}^{\dagger}$ corresponds to a move that can only change the corresponding fundamental domain to intersect three adjacent copies. This implies the “at most exponential” direction by Lemma 5.2 and Theorem 5.4.
To see the other direction, let $A$ be an Anosov map of the torus, and $\varphi$ a point push around a curve $\gamma$. Then the conjugate $A^{n}\varphi A^{-n}$ is a point push around $A^{n}\gamma$, and therefore
$\mathrm{diam}(A^{n}\varphi A^{-n})(R_{0})$
grows exponentially in $n$. Hence, by Lemma 5.2, the distance $d_{\mathcal{M}^{\dagger}_{0}}(m_{0},A^{n}\varphi A^{-n}m_{0})$ also grows exponentially. On the other hand, by the triangle inequality, the distance $d_{\mathcal{M}^{\dagger}}(m_{0},A^{n}\varphi A^{-n}m_{0})$ grows linearly. Hence, no subexponential bound of $d_{\mathcal{M}^{\dagger}_{0}}$ in terms of $d_{\mathcal{M}^{\dagger}}$ is possible. ∎
As an easy consequence we get the following properness, which says that a set of vertices in the subgraph $\mathcal{M}^{\dagger}_{0}$ is bounded if and only if it is bounded in the ambient graph $\mathcal{M}^{\dagger}$ . This will be crucial later.
Corollary 5.7.
For any $r$ there is an $R$ so that: for all $m,m^{\prime}\in\mathcal{M}^{\dagger}_{0}$
$d_{\mathcal{M}^{\dagger}}(m,m^{\prime})\leq r\quad\Rightarrow\quad d_{\mathcal{M}^{\dagger}_{0}}(m,m^{\prime})\leq R.$
The phenomenon we exhibited here is completely analogous to behavior seen in mapping class groups. Here, Broaddus–Farb–Putman show that the kernel of the Birman exact sequence is exponentially distorted [BFP11].
5.1. Outlook
To finish this section, we describe an outlook over the content of the next three sections. The eventual goal will be to prove a distance formula bounding the distance between two markings $m,m^{\prime}$ from above.
Our formula is inspired by Masur-Minsky’s hierarchy formula. We begin by describing the simplest case of this formula, in the case of the torus. In this case, the (non fine) curve graph is the classical Farey graph, whose vertices correspond to alternatively rational numbers, or isotopy classes of simple closed curves on the torus. Edges correspond to intersection $1$ up to isotopy, and thus a marking for the torus corresponds exactly to an edge in the Farey graph.
Suppose we are given a geodesic $\alpha_{1},\ldots,\alpha_{n}$ in the Farey graph. Then $\{\alpha_{1},\alpha_{2}\}$ and $\{\alpha_{2},\alpha_{3}\}$ are markings which share a curve $\alpha_{2}$. The duals $\alpha_{1},\alpha_{3}$ differ by a power of a Dehn twist about $\alpha_{2}$. Moving from $\{\alpha_{1},\alpha_{2}\}$ to $\{\alpha_{2},\alpha_{3}\}$ can be seen as “moving across a fan of triangles” in the Farey graph. The length of the fan is the power of the Dehn twist.
Continuing on, moving from $\{\alpha_{2},\alpha_{3}\}$ to $\{\alpha_{3},\alpha_{4}\}$ corresponds to applying a suitable Dehn twist power about $\alpha_{3}$; the power is the length of the fan around $\alpha_{3}$.
Hence, one obtains an upper (and, in fact, also a lower bound) for the distance between $\{\alpha_{1},\alpha_{2}\}$ and $\{\alpha_{n-1},\alpha_{n}\}$ as the sum of the lengths of all the fans.
What makes this approach powerful is that the “length of the fan” between, say, $\{\alpha_{i-1},\alpha_{i}\}$ and $\{\alpha_{i},\alpha_{i+1}\}$, can be computed, up to a bounded error, as a twist of $\alpha_{1}$ and $\alpha_{n}$ at $\alpha_{i}$ without needing to know the intermediate curves $\alpha_{k}$ in the geodesic. This is the celebrated “bounded geodesic image theorem” in this context: the amount of twisting of $\alpha_{1}$ and $\alpha_{i-1}$ about $\alpha_{i}$ is uniformly bounded (and similar for $\alpha_{n},\alpha_{i+1}$). In this setting, a reason this works is that all $\alpha_{i}$ are in different isotopy classes.
Combining everything one obtains a distance formula in the marking graph in terms of curve graph distance and a sum of twists.
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{3}$
$\alpha_{4}$
$\alpha_{5}$
$\beta_{2}$
$\beta_{3}$
$\beta_{4}$
$\beta_{5}$
Figure 9. From geodesics in the Farey graph to paths in the marking graph
We aim to prove a similar upper bound in our fine setting. There are three obstacles which will be addressed in the next sections. First, we need to define fine twist numbers, and prove fine bounded geodesic image theorems. This is done in Section 6. A major technical difficulty arises since geodesics in the fine curve graph of the torus can contain many curves in the same isotopy classes – which means that the bounded geodesic image theorem cannot directly be used to relate the twists of $\alpha_{1}$ and $\alpha_{i-1}$ about $\alpha_{i}$. To overcome this problem, in Section 7 we discuss paths in the fine curve graph which stay in the same isotopy class, and how to extract paths in the marking graphs from them. Finally, in Section 8, we put the pieces together to obtain the distance formula.
6. Fine twist numbers and bounded geodesic image theorems
In the beginning of this section and the next one we discuss a general (orientable) surface $S$ of genus $\geq 1$. We will need only the case of the torus, but the tools developed here are likely useful in the general setting. In higher genus, edges of the curve graph corresponds to disjointness, see [BHW22].
We study the metric properties of the natural map that projects the fine curve graph to the arc graph of an annulus embedded in $S$. The main result, stated here in the torus, is that the projection of the axis of a loxodromic element is bounded (Lemma 6.7).
6.1. Fine twist numbers
Definition 6.1.
Let us call almost embedded annulus a map $A:[0,1]\times S^{1}\to S$ which is an embedding on the interior of the annulus, and on each boundary component separately. The almost embedded annulus $A$ is called essential if the images of the boundary curves are essential.
A typical example on the torus is the almost embedded annulus $A_{\alpha}$ obtained by cutting along an essential simple closed curve $\alpha$. At some point below we will need to deform such annuli, which will lead to more complicated almost embedded annuli. We will often abusively identify $A$ and its image in $S$.
Let $A$ be an almost embedded essential annulus on $S$. The interior of $A$ is the image of the interior $(0,1)\times S^{1}$ under the almost embedding map, the boundary $\partial A$ is the image of the boundary of $0,1]\times S^{1}$. Likewise, we will call an essential proper arc of $A$ the image of an element of the fine arc graph $\mathcal{A}^{\dagger}([0,1]\times S^{1})$. We denote $\mathcal{A}^{\dagger}(A)$ the graph of essential proper arcs of $A$, with an edge between two essential arcs if they are the images of disjoint essential arcs of $0,1]\times S^{1}$: thus $A$ induces a graph isomorphism between $\mathcal{A}^{\dagger}([0,1]\times S^{1})$ and $\mathcal{A}^{\dagger}(A)$. As usual, we see the graph $\mathcal{A}^{\dagger}(A)$ as a metric space, and in particular we can consider the Hausdorff distance $d_{\mathrm{Hdff}}$ between finite subsets of $\mathcal{A}^{\dagger}(A)$. We then have a projection
$\pi_{A}:\mathcal{C}^{\dagger}(S)\to\mathbb{P}(\mathcal{A}^{\dagger}(A))$
where $\pi_{A}(\beta)$ is the set of all essential proper arcs of $A$ included in $\beta$. This is a finite set that could be empty; its diameter is at most $2$ (and $2$ may indeed happen when $\pi_{A}(\beta)$ contains two essential arcs sharing a boundary point that comes from an ”inner tangency” of $\beta$ with the boundary of $A$).
$\beta$
$\beta^{\prime}$
A
$\pi_{A}(\beta)$
$\pi_{A}(\beta^{\prime})$
Figure 10. Projection into annuli
Definition 6.2.
If $\beta,\beta^{\prime}$ both have nonempty projection in the almost embedded annulus $A$, we can define a twist number as the Hausdorff distance between the projections in the arc graph:
$\mathrm{tw}_{A}(\beta,\beta^{\prime})=d_{\mathrm{Hdff}}(\pi_{A}(\beta),\pi_{A}(\beta^{\prime})).$
Since both sets $\pi_{A}(\beta),\pi_{A}(\beta^{\prime})$ have diameter at most $2$, this quantity is coarsely equal to the distance between any two points of both sets, and to the diameter of the set $\pi_{A}(\beta)\cup\pi_{A}(\beta^{\prime})$ in the arc graph: the difference between any two of these quantities is bounded by $4$. More precisely, we have the following inequalities in any metric space: denoting $m$ and $M$ respectively the minimum and maximum distance between points of $A$ and $B$,
$m\leq d_{\mathrm{Hdff}}(A,B)\leq M\leq\mathrm{diam}(A\cup B)\leq m+\mathrm{diam}(A)+\mathrm{diam}(B).$
In particular, $\mathrm{tw}_{A}(\beta,\beta^{\prime})\leq\mathrm{diam}(\pi_{A}(\beta)\cup\pi_{A}(\beta^{\prime}))$.
We begin by observing some basic properties of twist numbers.
Lemma 6.3.
Let $A\subset S$ be an almost embedded essential annulus.
  1. i)
    The projection $\pi_{A}(\alpha)$ is empty if and only if there is a curve $\gamma$ which is included in the interior of $A$ and disjoint from $\alpha$. In particular, in this case $d_{\mathcal{C}^{\dagger}}(\delta,\alpha)\leq 2$, where $\delta$ denote one of the boundary curves of $A$, $\pi_{A_{\alpha}}(\delta)$ is also empty, and $\alpha$ is homotopic to $\delta$.
  2. ii)
    If $\pi_{A}(\alpha_{i})\neq\emptyset$ for $i=1,2$ and $d_{\mathcal{C}^{\dagger}}(\alpha_{1},\alpha_{2})\leq 1$, then
    $\mathrm{tw}_{A}(\alpha_{1},\alpha_{2})\leq 2.$
  3. iii)
    If $(g_{0},\dots,g_{\ell})$ is an edge path in $\mathcal{C}^{\dagger}(S)$ so that $\pi_{A}(g_{i})\neq\emptyset$ for all $i$, then
    $\mathrm{tw}_{A}(g_{0},g_{\ell})\leq 2\ell.$
  4. iv)
    If $\pi_{A}(\alpha_{i})\neq\emptyset$ for $i=1,2,3$ then there is a triangle inequality
    $\mathrm{tw}_{A}(\alpha_{1},\alpha_{3})\leq\mathrm{tw}_{A}(\alpha_{1},\alpha_{2})+\mathrm{tw}_{A}(\alpha_{2},\alpha_{3}).$
  5. v)
    The twist number is a topological invariant, in the sense that for every $h\in\mathrm{Homeo}(S)$ and all curves $\alpha,\alpha^{\prime}$ we have
    $\mathrm{tw}_{A}(\alpha,\alpha^{\prime})=\mathrm{tw}_{h(A)}(h\alpha,h\alpha^{\prime})$
    if one (hence both) sides are defined.
Proof.
If $\pi_{A}(\alpha)$ is empty, then every arc of $A$ included in $\alpha$ joins the same side of $A$ to itself, and defines a bigon with the boundary of $A$. Thus we may find a curve $\gamma$ in the interior of $A$ by pushing one side of $A$ off $\alpha$ along all bigons. This proves the direct implication in the first sentence of i). The reverse implication is obvious: if there is a common disjoint curve, then $\alpha$ cannot have an essential arc with respect to $\delta$. The remaining of i) is also obvious. Statements ii) and iii) follow from the observation that the projections of curves intersecting at most once also intersect at most once, and thus are distance at most two in the fine arc graph, by Lemma 4.5. Property iv) is simply the triangle inequality for the Hausdorff distance for finite subsets of the fine arc graph of $A$. The topological invariance v) is straightforward. ∎
Concerning point iii), we emphasize that the projection $\pi_{A}$ is not Lipschitz on its domain of definition. Indeed, given a curve $\alpha$, there are curves $\alpha^{\prime},\alpha^{\prime\prime}$ which both intersect $\alpha$ once and transversely, and with arbitrarily high twist $\mathrm{tw}_{A_{\alpha}}(\alpha^{\prime},\alpha^{\prime\prime})$. Then $(\alpha^{\prime},\alpha,\alpha^{\prime\prime})$ is a geodesic in $\mathcal{C}^{\dagger}(S)$ and $\pi_{A_{\alpha}}(\alpha^{\prime}),\pi_{A_{\alpha}}(\alpha^{\prime\prime})$ are both non-empty and arbitrarily far away. Nevertheless, it turns out that geodesics with non trivial projections have bounded images.
6.2. Bounded geodesic image theorems
We have the following bounded geodesic image theorems, optimizing either for minimizing the projection bound, or the distance to the projection annulus. Both of these are inspired by Masur-Minsky’s bounded geodesic image theorem, originally proved in [MM00]. Also compare [LT25] for a bounded geodesic image theorem in higher genus fine curve graphs projecting to non-annular subsurfaces. In the following lemmas, $\mathrm{diam}$ denotes the diameter of sets of vertices in ${\mathcal{A}^{\dagger}}(A)$, and $B_{r}(\alpha)$ denotes a ball in the fine curve graph.
Lemma 6.4 (First bounded geodesic image theorem).
For any $K>0$ there is a number $r>0$ so that the following is true. Let $A\subset S$ be an almost embedded essential annulus, and $\alpha$ be any of the boundary curves of $A$. Suppose that $(c_{i})$ is a KK–quasigeodesic in $\mathcal{C}^{\dagger}(S)$ which is disjoint from $B_{r}(\alpha)$. Then
$\mathrm{diam}(\cup_{i}\pi_{A}(c_{i}))\leq 2.$
Proof of Lemma 6.4.
We argue by contraposition. Suppose that
$\mathrm{diam}(\cup_{i}\pi_{A}(c_{i}))\geq 3.$
Since each $\pi_{A}(c_{i})$ has diameter at most $2$, there are $i\neq j$ and arcs $a\subset c_{i},b\subset c_{j}$ so that
  1. i)
    $a,b$ are essential proper arcs in $A$,
  2. ii)
    in the universal cover of $A$, any lift of $b$ meets two lifts of $a$, that is, $a\cup b$ fills the annulus $A$.
Hence, there are subarcs $a^{\prime}\subset a,b^{\prime}\subset b$ so that $a^{\prime}\cup b^{\prime}$ is a bicorn defined by $c_{i}$ and $c_{j}$ If $a^{\prime}\cup b^{\prime}$ is actually included in the interior of $A$ then it is distance $1$ from $\alpha$. If $a^{\prime}\cup b^{\prime}$ meets the boundary of $A$, then we can still find a curve which intersects exactly once, and transversely, both $\alpha$ and $a^{\prime}\cup b^{\prime}$, and thus the distance between the two curves is equal to $2$.
By Corollary 4.4, there a universal constant $M$ such that a geodesic $c_{i}c_{j}]$ enters the $M$-neighborhood of $a^{\prime}\cup b^{\prime}$, and thus also the $M+2$-neighborhood of $\alpha$. Since $(c_{i})$ is a KK–quasigeodesic, this implies (by the Morse lemma) that there is a constant $r=r(K,M,\delta)$ depending only on the quasigeodesic constants and the hyperbolicity constant of the fine curve graph, so that some $c_{k}\in B_{r}(\alpha)$. ∎
The second version below provides a non-explicit bound on the diameter of the projection, for any quasigeodesic path under the only condition that every element has a nonempty projection. Note that this hypothesis is satisfied as soon as the path does not enter the ball of radius 2 around the boundary curve of the annulus (by point i) of lemma 6.3).
Lemma 6.5 (Second bounded geodesic image theorem).
For any $K>0$ there is a $B>0$ so that the following is true. Let $A\subset S$ be an almost embedded essential annulus, and suppose that $(c_{0},\dots,c_{N})$ is an edge-path which is a KK–quasigeodesic in $\mathcal{C}^{\dagger}(T^{2})$ so that $\pi_{A}(c_{i})$ is never empty. Then
$\mathrm{diam}(\cup_{i}\pi_{A}(c_{i}))\leq B.$
We thank Katie Mann for pointing out to us that this second version follows easily from the first one; our initial proof was an adaptation of Richard Webb’s argument in the usual curve graph ([Web15]).
Proof.
Let $r$ be given by the previous lemma, and $\alpha$ denotes one of the boundary curve of $\partial A$. The case when the path $(c_{i})$ does not enter the ball $B_{r}(\alpha)$ follows at once from the previous lemma (with $B=2$), so let us assume there is some $i_{0}$ such that $c_{i_{0}}\in B_{r}(\alpha)$. Note that since $(c_{i})$ is a K-quasigeodesics, there exists some $L=L(K,r)$ such that whenever $|i-j|\geq L$, the distance between $c_{i}$ and $c_{j}$ in the fine curve graph is more than $2r$. Thus the path decomposes as the concatenation of the three sub-paths $\gamma_{1}\cup\gamma_{2}\cup\gamma_{3}$ with $\gamma_{1},\gamma_{3}$ (maybe empty) entirely outside $B_{r}(\alpha)$, and $\gamma_{2}$ containing $c_{i_{0}}$ of length no more than $2L$. By Lemma 6.3, iii), the projection of $\gamma_{2}$ has diameter no more than $4L$, and by the first Bounded Geodesic Image Theorem the projections of $\gamma_{1},\gamma_{3}$ have diameter at most 2. Thus we get the lemma with $B=4L+4$. ∎
In Section 7 we will prove a third bounded geodesic image theorem for a (slightly more technical) class of paths.
6.3. Twist bounds along axes
Here, we restrict to the torus again. First, we define projections and twist numbers of markings in the obvious way.
Namely, if $m=(\alpha,\beta)$ is a fine marking, we define
$\pi_{A}(m)=\pi_{A}(\alpha)\cup\pi_{A}(\beta).$
The key advantage is that since $\alpha$ and $\beta$ are not homotopic, at least one of them has non-empty projection to any annulus $A$, so this set is never empty. Furthermore it has diameter at most $2$, since $\alpha$ and $\beta$ intersect only once. The twist number between two markings $m,m^{\prime}$ is then defined as
$\mathrm{tw}_{A}(m,m^{\prime})=d_{\mathrm{Hdff}}(\pi_{A}(m),\pi_{A}(m^{\prime})).$
Note that since $\pi_{A}(m)$ and $\pi_{A}(m^{\prime})$ both have diameter at most $2$, for every $\alpha\in m,\alpha^{\prime}\in m^{\prime}$ we have $\mathrm{tw}_{A}(\alpha,\alpha^{\prime})=\mathrm{tw}_{A}(m,m^{\prime})\pm 4$. We then have the following lemma which includes a version of Lemma 6.3, iii) for markings, which removes the requirement of nonempty projection.
Lemma 6.6.
Let $A$ be an almost embedded essential annulus in $T^{2}$.
  1. i)
    The projection
    $\pi_{A}:\mathcal{M}^{\dagger}\to(\mathbb{P}\mathcal{A}^{\dagger}(A),d_{\mathrm{Hdff}})$
    is a 22–Lipschitz map.
  2. ii)
    If the two markings $m=(\alpha,\beta)$ and $m^{\prime}=(\alpha,\beta^{\prime})$ share the curve $\alpha$, then
    $d_{\mathcal{M^{\dagger}}}(m,m^{\prime})\leq\mathrm{tw}_{A_{\alpha}}(\beta,\beta^{\prime}).$
  3. iii)
    There is a triangle inequality
    $\mathrm{tw}_{A}(m_{1},m_{3})\leq\mathrm{tw}_{A}(m_{1},m_{2})+\mathrm{tw}_{A}(m_{2},m_{3}),$
  4. iv)
    The twist number is a topological invariant, in the sense that for every $h\in\mathrm{Homeo}(T^{2})$ and every markings $m,m^{\prime}$ we have
    $\mathrm{tw}_{A}(m,m^{\prime})=\mathrm{tw}_{h(A)}(hm,hm^{\prime}).$
Proof.
To prove point i), it suffices to check that the twist number between any two adjacent markings $m=\{\alpha,\beta\},m^{\prime}=\{\alpha,\beta^{\prime}\}$ is at most 2. First consider the case when $\pi_{A}(\alpha)\neq\emptyset$. Then the sets $\pi_{A}(m),\pi_{A}(m^{\prime})$ have some commun element; since their diameter is at most $2$, the Hausdorff distance between them is bounded by $2$, as wanted. It remains to consider the case when $\pi_{A}(\alpha)=\emptyset$. Then $\pi_{A}(m)=\pi_{A}(\beta),\pi_{A}(m^{\prime})=\pi_{A}(\beta^{\prime})$, both sets are non-empty, and since $\beta$ and $\beta^{\prime}$ are distance $1$ in the fine curve graph, the wanted inequality is exactly Lemma 6.3, point ii).
Let us prove ii). Denote $A=A_{\alpha}$ the almost embedded annulus determined by $\alpha$. Here $\pi_{A}(\beta)$, $\pi_{A}(\beta^{\prime})$ are singletons and $r=\mathrm{tw}_{A}(\beta,\beta^{\prime})$ is by definition equal to the distance between $\pi_{A}(\beta)$ and $\pi_{A}(\beta^{\prime})$ in the fine arc graph $\mathcal{A}^{\dagger}(A)$. Let $(b_{0},\dots,b_{r})$ be a geodesic in $\mathcal{A}^{\dagger}(A)$ from $\pi_{A}(\beta)$ to $\pi_{A}(\beta^{\prime})$. One can modify $b_{i}$ inductively so that the almost embedding map $A$ maps both endpoints of $b_{i}$ to the same point, so that the image $\beta_{i}$ of $b_{i}$ under the map $A$ is a simple closed curve. Now the sequence $(\alpha,\beta_{0}),\dots,(\alpha,\beta_{r})$ is a path in the fine marking graph from $m$ to $m^{\prime}$, showing that $d_{\mathcal{M^{\dagger}}}(m,m^{\prime})\leq r$.
Points iii) and iv) are immediate. ∎
As a consequence of the Bounded Geodesic Image theorem, we obtain the following lemma which shows that hyperbolic elements have uniformly bounded twists along their axes.
First, we observe that given any base marking $m_{0}$, and any $f\in\mathrm{Homeo}_{0}(T^{2})$ a homeomorphism acting hyperbolically on the fine curve graph, the marking $m_{0}$ can be extended to an ff–invariant path $(m_{i})$ in the marking graph. Also observe that for any such invariant path there is a constant $K$, depending on $m_{0}$ and $f$, so that $(m_{i})$ defines a (parametrized) bi-infinite KK–quasigeodesic in the fine curve graph.
Lemma 6.7 (Twist bounds along axes).
Suppose $f\in\mathrm{Homeo}_{0}(T^{2})$ acts hyperbolically on the fine curve graph, and let $(m_{i})$ be an ff–invariant path in the marking graph.
Then there is a number $T$ so that for every almost embedded essential annulus $A$ and every $i,j$ we have
$\mathrm{tw}_{A}(m_{i},m_{j})\leq T.$
Proof.
Let $A$ be any almost embedded annulus. Let $N$ be so that $m_{i+N}=fm_{i}$ for all $i$. We can choose some $L=L(K)$ and indices $i_{0}<i_{1}$ with the following properties:
  1. (1)
    for every $i$, if either $i\leq i_{0}$ or $i\geq i_{1}$, then $\pi_{A}(c_{i})\neq\emptyset$, and
  2. (2)
    $i_{1}-i_{0}\leq L$.
This is due to the fact that $\pi_{A}(\beta)=\emptyset$ implies that the curve graph distance of $\beta$ to a boundary curve of $A$ is at most $2$.
Now, the second Bounded Geodesic Image Theorem (Lemma 6.5) implies that
$\mathrm{diam}(\cup\{\pi_{A}(m_{i}),i\leq i_{0}\})\leq B,\quad\mathrm{diam}(\cup\{\pi_{A}(m_{i}),i\geq i_{1}\})\leq B,$
while the fact that twists are Lipschitz in the marking graph implies that
$\mathrm{diam}(\cup\{\pi_{A}(m_{i}),i_{0}\leq i\leq i_{1}\})\leq 2L+2.$
Together this shows that $T=2B+2L+2$ satisfies the desired bound, since $\mathrm{tw}_{A}(m_{i},m_{j})\leq\mathrm{diam}(\pi_{A}(m_{i})\cup\pi_{A}(m_{j}))\leq\mathrm{diam}(\cup\{\pi_{A}(m_{k}),i\leq k\leq j\})$. ∎
6.4. Width in annuli
In this section we discuss the notion of width of a curve relative to some annulus $A$. Roughly speaking, the width generalizes the twist number for arcs and curves that do not project in $A$. This will be useful in the next section when constructing duals in order make efficient moves in the fine marking graph.
In this section we also consider any orientable surface $S$ of genus $\geq 1$.
Definition 6.8.
Consider an almost embedded essential annulus $A$ in the surface $S$. If $a,b$ are any embedded (compact) arcs in $A$, then the relative width $\mathrm{wd}_{A}(a,b)$ is defined as the number of lifts of $a$ met by one lift of $b$ in the universal cover of $A$. If $\alpha,\beta$ are any sets in $S$, e.g. simple closed curves, then $\mathrm{wd}_{A}(\alpha,\beta)$ is defined as the maximum of $\mathrm{wd}_{A}(a,b)$ where $a,b$ are arcs included in $A$ and respectively in $\alpha$ and $\beta$. Note that here we do not assume any kind of transversality between our curves and the boundary of $A$. Furthermore the definition is compatible with the definition already given in $\mathcal{A}^{\dagger}(A)$ (see Lemma 4.5). This is clearly a symmetric quantity, namely $\mathrm{wd}_{A}(\alpha,\beta)=\mathrm{wd}_{A}(\beta,\alpha)$ for every $\alpha,\beta$.
The notions of the previous section readily extend to arcs. Namely, given an arc $a$ and an almost embedded essential annulus $A$ in the surface $S$, we define the projection $\pi_{A}(a)$ to be the set of all essential proper arcs of $A$ included in $a$. Now the twist $\mathrm{tw}_{A}(\alpha,\beta)=d_{\mathrm{Hdff}}(\pi_{A}(\alpha),\pi_{A}(\beta))$ makes sense whenever $\alpha,\beta$ are arcs or curves with non-empty projection in $A$.
Lemma 6.9 (Width is a twist).
Suppose $A$ is an almost embedded essential annulus in a surface $S$, $b_{0}\subset A$ an essential arc, and $\alpha$ a simple closed curve.
  1. i)
    if $\pi_{A}(\alpha)\neq\emptyset$, then
    $\mathrm{wd}_{A}(b_{0},\alpha)-1\leq\mathrm{tw}_{A}(b_{0},\alpha)\leq\mathrm{wd}_{A}(b_{0},\alpha)+1.$
  2. ii)
    if $\pi_{A}(\alpha)=\emptyset$ but $\alpha$ meets the interior of $A$, then there is an almost embedded essential annulus $A_{0}\subset A$ so that $\pi_{A_{0}}(\alpha)\neq\emptyset$ and
    $\mathrm{wd}_{A}(b_{0},\alpha)-1\leq\mathrm{tw}_{A_{0}}(b_{0},\alpha)\leq\mathrm{wd}_{A}(b_{0},\alpha)+1.$
Proof.
$b_{0}$
$\alpha$
Figure 11. Width is a twist
Let us prove i). Since $b_{0}$ is an essential arc, $\pi_{A}(b_{0})=\{b_{0}\}$. Given the definition of the twist and the formula relating the distance in $\mathcal{A}^{\dagger}(A)$ with the width (lemma 4.5), we have
$\mathrm{tw}_{A}(b_{0},\alpha)=\max_{a}d_{\mathcal{A}^{\dagger}(A)}(b_{0},a)=\max_{a}\mathrm{wd}_{A}(b_{0},a)+1$
where $a$ runs over the set $\pi_{A}(\alpha)$ made of all essential proper arcs of $A$ included in $\alpha$. On the other hand, $\mathrm{wd}_{A}(b_{0},\alpha)+1$ is defined by the same rightmost formula, except that now $a$ runs over all arcs $a\subseteq\alpha$ of $A$, not necessarily essential. Thus the upper bound in i) is clear. For the lower bound, consider an essential arc $a_{0}$ of $A$ included in $\alpha$, and denote $\Delta_{0}$ the closed fundamental domain cut by $a_{0}$ in the universal cover of the annulus. If $a$ is any other arc included in $\alpha$, any lift of $a$ is included in the union of two consecutive copies of $\Delta_{0}$ (the worst case happens when $\alpha$ has a ”tangency” at both ends of $_{0}, aa$ contains $a_{0}$ and continues at both ends of $a_{0}$ on opposite sides of $a_{0}$). Thus it meets at most two more lifts of $b_{0}$ than $a_{0}$ does. Thus $\mathrm{wd}_{A}(b_{0},a)\leq\mathrm{wd}_{A}(b_{0},a_{0})+2$. By choosing $a_{0}$ such that $\mathrm{tw}_{A}(b_{0},\alpha)=\mathrm{wd}_{A}(b_{0},a_{0})+1$, this gives the lower bound.
Let us prove ii). If every connected component of $\alpha\cap A$ meets $b_{0}$ at most once, then any essential embedded annulus $A_{0}\subset A$ in which $\alpha$ has non-empty projection satisfies the conclusion of the lemma.
From now on we assume that some component of $\alpha\cap A$ meets $b_{0}$ at least twice. We will obtain $A_{0}$ by pushing $A$ along $b_{0}$, taking care that the width does not decrease too much. By definition of width, there is a connected component $\bar{\alpha}$ of $\alpha\cap A$ such that $\mathrm{wd}_{A}(b_{0},\alpha)=\mathrm{wd}_{A}(b_{0},\bar{\alpha})$, and we can choose $\bar{\alpha}$ that meets $b_{0}$ at least twice. We build $A_{0}\subseteq A$ by pushing both boundary components of $A$ along $b_{0}$ until they both touch $\bar{\alpha}$, each at a single point on $b_{0}$. By construction we have $\bar{\alpha}\subset A_{0}$, and $b_{0}^{\prime}=b_{0}\cap A_{0}$ is connected. The universal cover $\widetilde{A_{0}}$ of $A_{0}$ naturally sits inside the universal cover $\widetilde{A}$ of $A$. Unrolling the definitions we see that $\mathrm{wd}_{A_{0}}(b_{0},\alpha)=\mathrm{wd}_{A_{0}}(b_{0},\bar{\alpha})=\mathrm{wd}_{A}(b_{0},\bar{\alpha})=\mathrm{wd}_{A}(b_{0},\alpha)$. Now the estimate follows from point i) applied to $A_{0}$ and $b_{0}^{\prime}$. ∎
7. The isotopy-constrained fine curve graph
This section is concerned with a subgraph of the fine curve graph, which will play a central role in our construction of paths in the marking graph. We fix an essential simple closed curve $\alpha_{0}$ in the torus throughout. Let
$\mathcal{C}^{\dagger}_{[\alpha_{0}]}(T^{2})=\mathcal{C}^{\dagger}_{0}(T^{2})\subset\mathcal{C}^{\dagger}(T^{2})$
be the full subgraph of the fine curve graph spanned by those curves in the isotopy class of $\alpha_{0}$. We denote the induced path-distance on this graph by $d_{[\alpha]}$, or $d_{0}$ if the base curve $\alpha_{0}$ is understood.
Note that in general $d_{0}$ is much bigger than $d$: if $\beta$ is a curve not in the isotopy class of $\alpha_{0}$, then the projection of any curve isotopic to $\alpha_{0}$ to the complementary annulus $A_{\beta}$ of $\beta$ is nonempty, and therefore
$\mathcal{C}^{\dagger}_{0}(T^{2})\to\mathbb{Z},\quad\alpha\mapsto\mathrm{tw}_{A_{\beta}}(\alpha_{0},\alpha)$
is a Lipschitz function. Hence, there is a curve $\alpha$ at distance two of $\alpha_{0}$ in $\mathcal{C}^{\dagger}(T^{2})$ whose distance to $\alpha_{0}$ in $\mathcal{C}^{\dagger}_{0}(T^{2})$ can be arbitrarily large. For a similar reason $\mathcal{C}^{\dagger}_{0}(T^{2})$ is not hyperbolic (one can find a quasi-flat by considering disjoint annuli).
The graph $\mathcal{C}^{\dagger}_{0}(T^{2})$ is maybe more similar to the fine arc graph of an annulus than the full fine curve graph of the torus. Namely, let us choose coordinates so that $T^{2}\simeq\mathbb{R}^{2}/\mathbb{Z}^{2}$ and $\alpha_{0}=\{0\}\times\mathbb{R}/\mathbb{Z}$. Consider the infinite cyclic cover $A\simeq\mathbb{R}\times\mathbb{S}^{1}\to T^{2}$, where curves isotopic to $\alpha_{0}$ lift to simple closed curves. We denote the deck transformation of this cover by $T$ (which, in our coordinates is simply the shift by $1$ in the $\mathbb{R}$-direction).
Lemma 7.1 (see [LRW24], Section 4.3).
Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$ be two distinct curves, and let $\widehat{\alpha},\widehat{\alpha}^{\prime}$ be lifts to $A$. Then
$d_{0}(\alpha,\alpha^{\prime})=\#\{n\in\mathbb{Z},\widehat{\alpha}\cap T^{n}\widehat{\alpha}^{\prime}\neq\emptyset\}+1,$
that is: the distance in the graph $\mathcal{C}^{\dagger}_{0}(T^{2})$ is equal to the number of lifts to $A$ of $\alpha$ met by one lift to $A$ of $\alpha^{\prime}$, plus one.
We remark that $d_{0}$ is also the minimal degree of a finite cover of the torus in which the curves have disjoint lifts.
7.1. Statement of the main proposition
The rest of this section is concerned with stating and proving Proposition 7.3, which is a key tool we need in the next section in order to estimate fine marking graph distance. The proposition will construct markings which in some sense “follow” an efficient path in $\mathcal{C}^{\dagger}_{0}$. At first reading, it might be useful to skip the proof to see how this proposition is used later.
The statement of the central proposition relies on the following notion.
Definition 7.2.
Let $\sigma,\tau\in\mathcal{C}^{\dagger}_{0}(T^{2})$, and $n>0$. We say that $\tau$ is almost distance $n$ from $\sigma$ if it belongs to the $C^{0}$-closure of the ball of radius $n$ centered at $\sigma$ in $\mathcal{C}^{\dagger}_{0}(T^{2})$. A quasi-path from $\alpha$ to $\alpha^{\prime}$ is a sequence $(\sigma_{0},\dots,\sigma_{n})$ in $\mathcal{C}^{\dagger}_{0}(T^{2})$, such that consecutive curves $\sigma_{i},\sigma_{i+1}$ are almost distance $1$, for each $i=0,\dots,n-1$.
Remember that we say that $\beta$ is a dual to $\alpha$ if the curves intersect transversely in one point (i.e. $\{\alpha,\beta\}$ is a fine marking). The following proposition provides an efficient sequence of fine markings connecting two given markings $\{\alpha,\beta\}$ and $\{\alpha,^{\prime},\beta^{\prime}\}$ where $\alpha$ and $\alpha^{\prime}$ are in the same isotopy class. All distances between two consecutive markings are bounded, up to constants, by twists in some annuli between the duals $\beta$ and $\beta^{\prime}$.
Proposition 7.3 (Efficient isotopic markings).
There exists constants $B,C$ such that the following holds. Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$, and denote $n=d_{0}(\alpha,\alpha^{\prime})$ the distance in the graph. Let $\beta,\beta^{\prime}$ be duals respectively to $\alpha,\alpha^{\prime}$.There exist
  • a quasi-path $\sigma_{0}=\alpha,\dots,\sigma_{n}=\alpha^{\prime}$ in $\mathcal{C}^{\dagger}_{0}(T^{2})$,
  • curves $\tau_{0}^{-}=\beta$, $\tau_{0}^{+},\tau_{1}^{-},\tau_{1}^{+},\dots,\tau_{n-1}^{-},\tau_{n-1}^{+},\tau_{n}^{-}$, $\tau_{n}^{+}=\beta^{\prime}$,
  • and annuli $A_{i}\subseteq A_{\sigma_{i}}$
such that:
  1. (1)
    each $\tau_{i}^{\pm}$ is dual to $\sigma_{i}$, we denote $m_{i}^{-}=\{\sigma_{i},\tau_{i}^{-}\}$ and $m_{i}^{+}=\{\sigma_{i},\tau_{i}^{+}\}$,
  2. (2)
    $d(m_{i}^{+},m_{i+1}^{-})\leq C$ for all $i=0,\dots,n-1$,
  3. (3)
    $d(m_{0}^{-},m_{0}^{+})\leq\mathrm{tw}_{A_{0}}(\beta,\beta^{\prime}){+B+1}$,
  4. (4)
    $d(m_{1}^{-},m_{1}^{+})\leq\mathrm{tw}_{A_{1}}(\beta,\beta^{\prime})+\mathrm{tw}_{A_{0}}(\beta,\beta^{\prime})+{4+2B}$,
  5. (5)
    $d(m_{i}^{-},m_{i}^{+})\leq\mathrm{tw}_{A_{i}}(\beta,\beta^{\prime})+2B+6$, for all $i=2,\dots,n-2$,
  6. (6)
    $d(m_{n-1}^{-},m_{n-1}^{+})\leq\mathrm{tw}_{A_{n-1}}(\beta,\beta^{\prime})+\mathrm{tw}_{A_{n}}(\beta,\beta^{\prime})+{4+2B}$,
  7. (7)
    $d(m_{n}^{-},m_{n}^{+})\leq\mathrm{tw}_{A_{n}}(\beta,\beta^{\prime})+{B+1}$,
where all the distances are in the fine marking graph. Furthermore, (8) for each $i=0,\dots,n$, the curve $\sigma_{i}$ is almost distance $i$ from $\alpha$, and almost distance nin-i from $\alpha^{\prime}$.
$m_{0}^{-}$
$m_{0}^{+}$
$m_{1}^{-}$
$m_{1}^{+}$
$\sigma_{0}$
$\sigma_{1}$
$\sigma_{n-1}$
$\sigma_{n}$
$(3)$
$(2)$
$(4)$
$m_{2}^{-}$
$m_{2}^{+}$
$(5)$
$(2)$
$\sigma_{2}$
$\tau_{0}^{-}$
$\tau_{0}^{+}$
$\tau_{2}^{-}$
$\tau_{1}^{-}$
$\tau_{2}^{+}$
$\tau_{1}^{+}$
$\tau_{n}^{-}$
$\tau_{n}^{+}$
$\dots$
$(6)$
$(7)$
$\color[rgb]{1,0,0}\alpha=$
$\color[rgb]{1,0,0}\beta=$
$\color[rgb]{1,0,0}=\alpha^{\prime}$
$\color[rgb]{1,0,0}=\beta^{\prime}$
Figure 12. Efficient sequence of fine markings connecting two curves in the same isotopy class. Numbers refer to upper bounds in Proposition 7.3.
7.2. Quasi-paths
Let us first give a more concrete, geometrical characterization of being at almost distance at most $1$, which will show in particular that this is a symmetric notion. From now on we fix an orientation on our base curve $\alpha_{0}$, and orient each curve isotopic to $\alpha_{0}$ accordingly, so that the oriented curves are isotopic. Consider the infinite cyclic cover $A=\mathbb{R}\times\mathbb{S}^{1}$ as above where $\alpha_{0}$ lifts to a closed curve, and let $T:A\to A$ denote the covering transformation $(x,y)\mapsto(x+1,y)$. We may choose the orientation so that for each lift $\widetilde{\alpha}$ of $\alpha\in\mathcal{C}^{\dagger}_{0}(T^{2})$, the lift $T\widetilde{\alpha}$ is on the right-hand side of $\widetilde{\alpha}$. Given two essential simple closed curves $a,b$ in $A$, we write $a<b$ if $b$ is included in the open domain bounded by $a$ on its right-hand side, and $a\leq b$ if $b$ is included in the closure of this domain.
Lemma 7.4.
Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$. Fix a lift $\widetilde{\alpha}$ of $\alpha$ in $A$.
  1. i)
    $\alpha^{\prime}$ and $\alpha$ are disjoint exactly if $\alpha^{\prime}$ has a lift $\widetilde{\alpha}^{\prime}$ such that $\widetilde{\alpha}<\widetilde{\alpha}^{\prime}<T\widetilde{\alpha}$.
  2. ii)
    The curve $\alpha^{\prime}$ is almost distance $1$ from $\alpha$ if and only if $\alpha^{\prime}$ has a lift $\widetilde{\alpha}^{\prime}$ such that $\widetilde{\alpha}\leq\widetilde{\alpha}^{\prime}\leq T\widetilde{\alpha}$.
Proof.
Part i) is obvious. It also immediately implies that being almost distance $1$ implies the criterion ii) in the lemma. Conversely, if $\widetilde{\alpha}^{\prime}$ satisfies condition ii), there are arbitrarily small deformations so that $\widetilde{\alpha}<\widetilde{\alpha}^{\prime}<T\widetilde{\alpha}$, and thus $\alpha^{\prime}$ has almost distance $1$. ∎
Assuming the situation of the lemma, the curves are distance one if and only if $\widetilde{\alpha}<\widetilde{\alpha}^{\prime}<T\widetilde{\alpha}$. In the opposite case, we will say that $\alpha^{\prime}$ touches $\alpha$ only from the right if $\widetilde{\alpha}^{\prime}<T\widetilde{\alpha}$, $\alpha^{\prime}$ touches $\alpha$ only from the left if $\widetilde{\alpha}<\widetilde{\alpha}^{\prime}$, and $\alpha^{\prime}$ touches $\alpha$ only from both sides if none of the two inequalities hold. Rephrasing the lemma, a curve is almost distance $1$ from $\alpha$ if and only if it either touches $\alpha$ from the right, or from the left, or from both sides, or is disjoint from it.
Note that $\alpha^{\prime}$ touches $\alpha$ from one side if and only if $\alpha$ touches $\alpha^{\prime}$ from the other side; and ”touching from both sides” is a symmetric relation. Also note that if $\alpha^{\prime}$ touches $\alpha$ from both sides then the projection $\pi_{\alpha}(\alpha^{\prime})$ in the annulus $A_{\alpha}$ is not empty.
We also need a version of the bounded geodesic image theorem for quasi-paths, which uses the following lemma.
Lemma 7.5.
Suppose $A$ is an almost embedded essential annulus in $T^{2}$, and $(\alpha_{i}),i=0,\ldots,n$ a quasi-path. If $\pi_{A}(\alpha_{i})\neq\emptyset$ for all $i$, then
$\mathrm{diam}(\cup_{i}\pi_{A}(\alpha_{i}))\leq 2n.$
Proof.
Observe that by Lemma 7.4, consecutive $\alpha_{i},\alpha_{i+1}$ do not intersect transversely, and thus the same is true for the projections $\pi_{A}(\alpha_{i}),\pi_{A}(\alpha_{i+1})$. This implies that
$\mathrm{tw}_{A}(\alpha_{i},\alpha_{i+1})\leq 2$
for all $n$, and thus the lemma by induction. ∎
Given the previous lemma, the proof of the second Bounded Geodesic Image Theorem applies verbatim to give the following.
Lemma 7.6 (Quasi-path bounded geodesic image theorem).
For any $K>0$ there is a $B>0$ so that the following is true. Let $A\subset S$ be an almost embedded essential annulus, and suppose that $(c_{0},\dots,c_{N})$ is a quasi-path in $\mathcal{C}^{\dagger}_{0}(T^{2})$ which is also a KK–quasigeodesic in $\mathcal{C}^{\dagger}(T^{2})$, so that $\pi_{A}(c_{i})$ is never empty. Then
$\mathrm{diam}(\cup_{i}\pi_{A}(c_{i}))\leq B.$
7.3. Proof of Proposition 7.3, Strategy
We will split our main proposition above into three statements. The first is concerned with constructing the quasi-path which will form the $\sigma_{i}$, and the other two provide the nice duals.
Proposition 7.7.
Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$, and denote $n=d_{0}(\alpha,\alpha^{\prime})$ the distance in the subgraph. Then there exists a quasi-path $\sigma_{0}=\alpha,\dots,\sigma_{n}=\alpha^{\prime}$ in $\mathcal{C}^{\dagger}_{0}(T^{2})$ with the following properties.
  1. (1)
    $\sigma_{1}$ touches $\sigma_{0}$ from the left, $\sigma_{n-1}$ touches $\sigma_{n}$ from the left,
  2. (2)
    for each $i=1,\dots,n-1$, the curves $\sigma_{i+1}$ touches $\sigma_{i}$ from both sides,
  3. (3)
    for each $i=0,\dots,n$, the curve $\sigma_{i}$ is almost distance $i$ from $\alpha$, and almost distance nin-i from $\alpha^{\prime}$.
In particular, note that $\pi_{A_{\sigma_{i}}}(\sigma_{j})\neq\emptyset$ if $|i-j|\geq 2$. Furthermore, all these curves are included in $\alpha\cup\alpha^{\prime}$.
Proposition 7.8.
Let $\sigma,\sigma^{\prime}\in\mathcal{C}^{\dagger}(T^{2})$ be isotopic curves that touch each other from one side. Let $\tau$ be a dual to $\sigma$. Then there exists a curve $\tau^{\prime}$ which is a common dual to $\sigma$ and $\sigma^{\prime}$, and an almost embedded annulus $A\subset A_{\sigma}$, in which $\sigma^{\prime}$ has non-empty projection, and such that $\mathrm{tw}_{\sigma}(\tau,\tau^{\prime})\leq\mathrm{tw}_{A}(\tau,\sigma^{\prime})$. Furthermore we can choose $\tau^{\prime}$ isotopic to $\tau$.
Proposition 7.9.
There exists a constant $C$ such that the following holds. Let $\sigma,\sigma^{\prime}\in\mathcal{C}^{\dagger}(T^{2})$ be isotopic curves that touch each other from both sides. Then:
  1. (1)
    There exist curves $\tau^{+},\tau^{\prime-}$ respectively dual to $\sigma,\sigma^{\prime}$ such that
    $\mathrm{tw}_{\sigma}(\tau^{+},\sigma^{\prime})\leq 2,\ \ \mathrm{tw}_{\sigma^{\prime}}(\sigma,\tau^{\prime-})\leq 2.$
  2. (2)
    Denoting $m=\{\sigma,\tau^{+}\}$, $m^{\prime}=\{\sigma^{\prime},\sigma^{\prime-}\}$, we have $d_{\mathcal{M}_{0}^{\dagger}}(m,m^{\prime})\leq C$.
The proofs are contained in the next subsections. We now deduce our main proposition.
Proof of Proposition 7.3.
Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$, $n=d(\alpha,\alpha^{\prime})$, $\beta$ a dual to $\alpha^{\prime}$. Proposition 7.7 provides a quasi-path $(\sigma_{0},\dots,\sigma_{n})$ from $\alpha$ to $\alpha^{\prime}$, with distance to the end-points satisfying the wanted property (8).
Let us construct the duals. We set $\tau_{0}^{-}=\beta$, and apply Proposition 7.8 with $\sigma=\sigma_{0},\sigma^{\prime}=\sigma_{1}$ which touch each other from one side, and dual $\tau_{0}^{-}$. We get an annulus $A_{0}\subseteq A_{\sigma_{0}}$ and a common dual $\tau^{\prime}$ for $\sigma_{0}$ and $\sigma_{1}$, and set $\tau_{0}^{+}=\tau_{1}^{-}=\tau^{\prime}$. Now we start estimating distances in the fine marking graph, using silently the basic properties of twists given by Lemma 6.6. Since $\sigma_{0}$ and $\sigma_{1}$ touch each other from one side and $\tau^{\prime}$ is a common dual, we can find a curve $\sigma^{\prime}$ which is disjoint from both $\sigma_{0},\sigma_{1}$ and still dual to $\tau^{\prime}$. This entails $d(m_{0}^{+},m_{1}^{-})\leq 2$, which is point (2) for $i=0$. According to Proposition 7.8 we have
$d(m_{0}^{-},m_{0}^{+})\leq\mathrm{tw}_{\sigma_{0}}(\tau_{0}^{-},\tau_{0}^{+})\leq\mathrm{tw}_{A_{0}}(\tau_{0}^{-},\sigma_{1})=\mathrm{tw}_{A_{0}}(\beta,\sigma_{1}).$
Now, $\sigma_{2},\ldots,\sigma_{n}$ is a quasi-path which is also a $2$-quasigeodesic in $\mathcal{C}^{\dagger}(T^{2})$, and by Proposition 7.7 the projections $\pi_{A_{0}}(\sigma_{i}),i\geq 2$ are non empty. Thus the Quasi-path bounded geodesic image theorem 7.6, together with the triangular inequality for twists, implies that
$\mathrm{tw}_{A_{0}}(\beta,\sigma_{1})\leq\mathrm{tw}_{A_{0}}(\beta,\sigma_{n})+B.$
Since $\sigma_{n},\beta^{\prime}$ intersect once, and both project to $A_{0}$, the definition of twists gives $\mathrm{tw}_{A_{0}}(\sigma_{n},\beta^{\prime})\leq 1$ and thus
$d(m_{0}^{-},m_{0}^{+})\leq\mathrm{tw}_{A_{0}}(\beta,\beta^{\prime})+B+1,$
which is point (3).
Next we apply Proposition 7.9 with $\sigma,\sigma^{\prime}=\sigma_{i},\sigma_{i+1}$ for each $i=1,\dots,n-2$. We get curves $\tau_{i}^{+},\tau_{i+1}^{-}$ which are duals respectively to $\sigma_{i},\sigma_{i+1}$. By the proposition we have $d(m_{i}^{+},m_{i+1}^{-})\leq C$, which gives point (2) for all $i$ except n1n-1.
Let us check (4), setting $A_{1}=A_{\sigma_{1}}$. By the proposition we have $\mathrm{tw}_{\sigma_{1}}(\tau_{1}^{+},\sigma_{2})\leq 2$, and thus
$\begin{array}[]{rcl}d(m_{1}^{-},m_{1}^{+})&\leq&\mathrm{tw}_{\sigma_{1}}(\tau_{1}^{-},\tau_{1}^{+})\\ &\leq&\mathrm{tw}_{\sigma_{1}}(\tau_{1}^{-},\beta)+\mathrm{tw}_{\sigma_{1}}(\beta,\sigma_{2})+2.\end{array}$
The first term $\mathrm{tw}_{\sigma_{1}}(\tau_{1}^{-},\beta)=\mathrm{tw}_{\sigma_{1}}(\tau_{0}^{+},\tau_{0}^{-})$ is bounded by the ”universal twist” between the isotopic curves $\tau_{0}^{+},\tau_{0}^{-}$, namely the number of lifts of $\tau_{0}^{+}$ met by one lift of $\tau_{0}^{-}$ in the universal cover of the torus. Since both curves are dual to $\sigma_{0}$, this coincides with $\mathrm{tw}_{\sigma_{0}}(\tau_{0}^{+},\tau_{0}^{-})$. We have proved above that this is bounded by $\mathrm{tw}_{A_{0}}(\beta,\sigma_{1})$. Finally we get
$d(m_{1}^{-},m_{1}^{+})\leq\mathrm{tw}_{A_{0}}(\beta,\sigma_{1})+\mathrm{tw}_{\sigma_{1}}(\beta,\sigma_{2})+2.$
As before, we note that the projections of $\sigma_{i}$ to $A_{\sigma_{0}},A_{\sigma_{1}}$ are nonempty for all $i>2$ by Proposition 7.7, and thus the Quasi-path bounded geodesic image theorem 7.6 again implies that we can replace $\sigma_{1},\sigma_{2}$ in the above formula by $\beta^{\prime}$ for an additional cost of $B+1$ each, yielding
$d(m_{1}^{-},m_{1}^{+})\leq\mathrm{tw}_{A_{0}}(\beta,\beta^{\prime})+\mathrm{tw}_{\sigma_{1}}(\beta,\beta^{\prime})+4+2B,$
which is (4).
Let $i\in\{2,\dots,n-2\}$, and let $A_{i}=A_{\sigma_{i}}$. To prove (5), we first note that the two markings $m_{i}^{-},m_{i}^{+}$ have the curve $\sigma_{i}$ in common, thus their distance is bounded by $\mathrm{tw}_{A_{i}}(\tau_{i}^{-},\tau_{i}^{+})$. Now the triangular inequality for twists and the bounds in point (1) of Proposition 7.9 gives
$d(m_{i}^{-},m_{i}^{+})\leq\mathrm{tw}_{A_{i}}(\sigma_{i-1},\sigma_{i+1})+4.$
The same argument with the bounded geodesic image theorem allows to replace the twist term by $\mathrm{tw}_{A_{i}}(\beta,\beta^{\prime})$ at the cost of $2B+2$, showing (5). We apply again Proposition 7.8 symmetrically with $\sigma=\alpha^{\prime},\sigma^{\prime}=\sigma_{n-1},\tau^{\prime}=\beta^{\prime}$, and get a curve $\tau_{n-1}^{+}=\tau_{n}^{-}$ which is a common dual to $\sigma_{n-1}$ and $\sigma_{n}$. Finally we get property (6) exactly as (4), and (7) exactly as (3). ∎
7.4. Operations on curves in the infinite annulus
In order to construct our quasi-path, we will mainly work, as before, in the infinite annulus $A\simeq\mathbb{R}\times\mathbb{S}^{1}$, remembering the covering transformation $T(x,y)=(x+1,y)$. This annulus has a natural compactification by adding two points denoted $\pm\infty$. Remark that $A\cup\{+\infty\}$ is homeomorphic to the plane, and a sequence tends to $\infty$ if and only if its first coordinate tends to $\infty$. Given a set $K$ whose first projection is bounded from above, we denote $R(K)$ the unique connected component of $A\setminus K$ whose closure contains $\infty$.
Denote $\mathcal{C}^{\dagger}(A)$ the set of essential simple closed curve in $A$, which corresponds to curves in the plane that surrounds $\infty$. Given two curves $\alpha,\beta\in\mathcal{C}^{\dagger}(A)$, we denote $\alpha\wedge\beta$ the boundary of $R(\alpha\cup\beta)$. A theorem of Kerekjarto (see [vK23]) asserts that $\alpha\wedge\beta$ is an element of $\mathcal{C}^{\dagger}(A)$, the remarkable fact being that this holds without any regularity nor transversality assumption. Thus we have an operation $\wedge:\mathcal{C}^{\dagger}(A)\times\mathcal{C}^{\dagger}(A)\to\mathcal{C}^{\dagger}(A)$. Note the following caracterization, which is an immediate consequence of the Schoenflies theorem: a point $x$ belongs to $\alpha\wedge\beta$ if and only if (1) it belongs to $\alpha\cup\beta$, and (2) it is accessible from $\infty$ in $R(\alpha)\cap R(\beta)$, that is, there is an embedding with $\rho(0)=x$, $\rho(+\infty)=+\infty$, and $\rho((0,1))\subset R(\alpha)\cap R(\beta)$. We call $\rho$ a witness ray for $x$ in $\alpha\wedge\beta$.
Let us recall the following relations on $\mathcal{C}^{\dagger}(A)$:
  • $\alpha<\beta$ if $\beta\subset R(\alpha)$,
  • $\alpha\leq\beta$ if $\beta\subset\mathrm{Clos}(R(\alpha))$.
The following lemma collects some classical properties of the wedge operation. The proof is left to the reader.
Lemma 7.10 (Algebra of curves).
For every $\alpha,\alpha^{\prime},\beta,\gamma$ the following holds:
  1. (1)
    $\alpha\wedge\beta=\beta\wedge\alpha$.
  2. (2)
    The relations $\leq$ are transitive, and $\alpha<\beta\leq\gamma\Rightarrow\alpha<\gamma$.
  3. (3)
    $\alpha\leq\alpha^{\prime}\Rightarrow\alpha\wedge\beta\leq\alpha^{\prime}\wedge\beta$.
  4. (4)
    $\alpha<\gamma,\beta<\gamma\Rightarrow\alpha\wedge\beta<\gamma$.
  5. (5)
    $\alpha\leq\alpha\wedge\beta$, and $\alpha<\alpha\wedge\beta\iff\alpha<\beta$, and in this case $\alpha\wedge\beta=\beta$.
  6. (6)
    $\alpha\cap T\alpha=\emptyset\Rightarrow\alpha<T\alpha$.
  7. (7)
    $T(\alpha\wedge\beta)=(T\alpha)\wedge(T\beta)$.
  8. (8)
    $\alpha<\beta\iff T\alpha<T\beta$, and likewise for $\leq$.
Lemma 7.11.
If $\alpha\cap T\alpha=\emptyset$ and $\beta\cap T\beta=\emptyset$ then $(\alpha\wedge\beta)\cap T(\alpha\wedge\beta)=\emptyset$.
Proof.
By transitivity (2) and (5) of the lemma we have $\alpha<T\alpha\leq(T\alpha)\wedge(T\beta)$, and likewise $\beta<(T\alpha)\wedge(T\beta)$, thus by (4) we get $(\alpha\wedge\beta)\cap T(\alpha\wedge\beta)=\emptyset$. ∎
7.5. Construction of the quasi-path
Coming back to the torus, let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$, and denote $n=d_{0}(\alpha,\alpha^{\prime})$ as above. We want to prove Proposition 7.7, and we may assume $n\geq 2$, otherwise there is nothing to prove. Choose lifts $\widetilde{\alpha},\widetilde{\alpha}^{\prime}$ such that the lifts met by $\widetilde{\alpha}^{\prime}$ are $T^{-(n-1)}\alpha,\dots,T^{-1}\alpha$. Consider
$\tilde{\sigma}_{i}=\tilde{\alpha}\wedge T^{i}\tilde{\alpha}^{\prime},\ \ i=0,\dots,n.$
Lemma 7.12 (Properties of the quasi-path).
The following holds:
  1. (i)
    $\widetilde{\sigma}_{0}=\widetilde{\alpha}$, $\widetilde{\sigma}_{n}=T^{n}\widetilde{\alpha}^{\prime}$, and each $\widetilde{\sigma}_{i}$ is an essential simple closed curve isotopic to $\widetilde{\alpha}$ and disjoint from its image under $T$.
  2. (ii)
    $\widetilde{\sigma}_{0}\leq\widetilde{\sigma}_{1}<T\widetilde{\sigma}_{0}$, and $\widetilde{\sigma}_{0}$ meets $\widetilde{\sigma}_{1}$;
  3. (iii)
    $\widetilde{\sigma}_{i}\leq\widetilde{\sigma}_{i+1}\leq T\widetilde{\sigma}_{i}$, and $\widetilde{\sigma}_{i+1}$ meets $\widetilde{\sigma}_{i}$ and $T\widetilde{\sigma}_{i}$, for each $1\leq i\leq n-2$;
  4. (iv)
    $\widetilde{\sigma}_{n-1}<\widetilde{\sigma}_{n}\leq T\widetilde{\sigma}_{n-1}$, and $\widetilde{\sigma}_{n}$ meets $T\widetilde{\sigma}_{n-1}$.
  5. (v)
    $\widetilde{\alpha}\leq\widetilde{\sigma}_{i}<T^{i}\widetilde{\alpha}$, and $T^{i}\widetilde{\alpha}^{\prime}\leq\widetilde{\sigma}_{i}<T^{n}\widetilde{\alpha}^{\prime}$, for $1\leq i\leq n-1$.
More geometrically, one can think of the $\sigma_{i}$’s the following way. Let $0<i<n$. Call $(a,b)$ an $i$-bigon if $a\cup b$ is a simple closed curve with $a$ a sub-arc of $\widetilde{\alpha}$, and $b$ a sub-arc of $T^{i}\widetilde{\alpha}^{\prime}$ on the right of $\widetilde{\alpha}$. Call an $i$-bigon maximal if it is not included in the disk bounded by any other $i$-bigon. Note that for each $i$ there are maximal $i$-bigons meeting $T^{i-1}\widetilde{\alpha}$. Then $\sigma_{i}$ is obtained from $\alpha$ by changing $a$ into $b$ for every maximal $i$-bigon $(a,b)$. Before proving the lemma, let us first check that it entails the wanted properties for the quasi-path.
Proof of Proposition 7.7.
Let $\alpha,\alpha^{\prime}\in\mathcal{C}^{\dagger}_{0}(T^{2})$, denote $n=d_{0}(\alpha,\alpha^{\prime})$. Consider the curves $\widetilde{\sigma}_{i}$, and let $\sigma_{i}$ denote their projection in the torus. By (i), these are simple closed curves in the right isotopy class. Each curve touches the next one from one or two sides, by points (ii), (iii), (iv) of the lemma, as required by points (1) and (2) of the proposition. For point (3), note that due to property (v) of the lemma, pushing $\sigma_{i}$ slightly to the right yields a curve $\alpha^{\prime}_{i}$ arbitrary close to $\sigma_{i}$ with a lift $\widetilde{\alpha}^{\prime}_{i}$ such that
$\widetilde{\alpha}<\widetilde{\alpha}^{\prime}_{i}<T^{i}\widetilde{\alpha}\text{ and }T^{i}\widetilde{\alpha}^{\prime}<\widetilde{\alpha}^{\prime}_{i}<T^{n}\widetilde{\alpha}^{\prime}$
and thus $d_{0}(\alpha,\alpha^{\prime}_{i})\leq i$ and $d_{0}(\alpha^{\prime}_{i},\alpha^{\prime})\leq n-i$. ∎
Proof of Lemma 7.12.
[…]
Let us prove (i). We have $\widetilde{\alpha}\leq\widetilde{\sigma}_{1}$, and the two curves meet, by (5). Since $\widetilde{\alpha}^{\prime}<\widetilde{\alpha}$ we have $T\widetilde{\alpha}^{\prime}<T\widetilde{\alpha}$, and since $\widetilde{\alpha}<T\widetilde{\alpha}$ we get $\widetilde{\sigma}_{1}<T\widetilde{\alpha}$ by (4). Point (4) of the lemma is similar.
Let us prove (ii). Since $T^{i}\widetilde{\alpha}^{\prime}<T^{i+1}\widetilde{\alpha}^{\prime}$ we get $\widetilde{\sigma}_{i}\leq\widetilde{\sigma}_{i+1}$ by (3). Likewise, from $\widetilde{\alpha}<T\widetilde{\alpha}$ we get $\widetilde{\sigma}_{i+1}\leq T\widetilde{\sigma}_{i}$. It remains to prove that the curve $\widetilde{\sigma}_{i+1}$ meet $\widetilde{\sigma}_{i}$ and $T\widetilde{\sigma}_{i}$. For this we need a claim.
Claim 7.13.
Let $0<i<n$, and consider a connected component $b$ of $\widetilde{\alpha}_{i}\setminus\widetilde{\alpha}$. Then $Tb$ is included in some connected component $B^{\prime}$ of $\widetilde{\alpha}_{i+1}\setminus\widetilde{\alpha}$.
The claim entails immediately that $\widetilde{\sigma}_{i+1}$ meets $T\widetilde{\sigma}_{i}$, as asserted in point (iii) of the lemma.
Proof.
Let $x\in b$ with $b$ as in the claim. By the characterization of $\tilde{\sigma}_{i}=\tilde{\alpha}\wedge T^{i}\tilde{\alpha}^{\prime}$, there is a witness ray $\rho$ for $x$. The curve $T\rho$ is clearly on the right-hand side of $\widetilde{\alpha}$ and $T^{i+1}\widetilde{\alpha}^{\prime}$, thus it is a witness ray for $Tx$. Thus $Tx$ belongs to $\widetilde{\sigma}_{i+1}$, and since it does not belong to $\widetilde{\alpha}$, it belongs to $\widetilde{\alpha}_{i+1}\setminus\widetilde{\alpha}$. This proves that $Tb$ is included in $\widetilde{\alpha}_{i+1}\setminus\widetilde{\alpha}$, and thus in some connected component $b^{\prime}$ of this set. ∎
The fact (iv) that $\widetilde{\sigma}_{i+1}$ meets $\widetilde{\sigma}_{i}$ has a similar proof, here is a sketch. Consider some point $x$ in $\widetilde{\sigma}_{i+1}\setminus T^{i+1}\widetilde{\alpha}^{\prime}$. A witness ray $\rho$ for $x$ in $\widetilde{\sigma}_{i+1}=\tilde{\alpha}\wedge T^{i+1}\tilde{\alpha}^{\prime}$ is on the right-hand side of both $\widetilde{\alpha}$ and $T^{i+1}\widetilde{\alpha}^{\prime}$, and thus also on the righ-hand side of $T^{i}\widetilde{\alpha}^{\prime}$, thus it is also a witness ray for $x$ in $\widetilde{\sigma}_{i}$. This proves that $\widetilde{\sigma}_{i+1}\setminus T^{i+1}\widetilde{\alpha}^{\prime}\subset\widetilde{\sigma}_{i}$, and thus both curves meet.
Let us prove (v). Let $1\leq i\leq n-1$. First $\widetilde{\alpha}\leq\widetilde{\alpha}\wedge T^{i}\widetilde{\alpha}^{\prime}=\widetilde{\sigma}_{i}$ by (5). For the second inequality, we have $\widetilde{\alpha}^{\prime}<\widetilde{\alpha}$, so $T^{i}\widetilde{\alpha}^{\prime}<T^{i}\widetilde{\alpha}$, and obviously $\widetilde{\alpha}<T^{i}\widetilde{\alpha}$, so $\widetilde{\sigma}_{i}<T^{i}\widetilde{\alpha}$ by (4). Likewise, we have $T^{i}\widetilde{\alpha}^{\prime}<\widetilde{\sigma}_{i}$ by (5), and from $T^{-n}\widetilde{\alpha}<\widetilde{\alpha}^{\prime}$ we get $\widetilde{\sigma}_{i}<T^{n}\widetilde{\alpha}^{\prime}$. ∎
7.6. Constructing duals
Proof of Proposition 7.8.
We work in the annulus $A_{\sigma}$ cut by $\sigma$. Let $c$ be an arc inside $A$ that starts on the side which is not touched by $\sigma^{\prime}$, is disjoint from the dual $\tau$, and ends at a point $x$ of $\sigma^{\prime}$; we may choose $c$ such that it is disjoint from $\sigma^{\prime}$ except at its endpoint (see Figure 13).
$\tau$
$\sigma^{\prime}$
$\sigma$
$x$
$D$
$c^{\prime}$
$c$
Figure 13. Constructing a dual in Proposition 7.8
Let $t\subset\sigma^{\prime}$ be the connected component of $\sigma^{\prime}\setminus\sigma$ containing $x$. Let $s\subset\sigma$ be the arc such that $s\cup t$ bounds an open disk $D$ in the annulus $A_{\sigma}$. Note that since $\sigma^{\prime}$ touches $\sigma$ on one side, it is disjoint from $D$. Let $c^{\prime}$ be any arc that starts at $x$ and ends at a point $y$ on $s$, and is included in $D$ but at its endpoints. Now we may modify $c$ close to $\sigma$ so that its starting point, seen in the torus, coincides with $y$; note that this can be done with $c$ still not meeting $\tau$. Now the curve $c\cup c^{\prime}$ defines a simple closed curve $\tau^{\prime}$ in the torus, which meets $\sigma$ only at $y$ and $\sigma^{\prime}$ only at $x$, and this is a common dual. (This is not our final curve $\tau^{\prime}$ since it is not isotopic to $\tau$; we will remedy for this at the very end.)
Let us check the inequality about twists. Remember that $c$ is disjoint from $\tau$. The arc $c^{\prime}$ is included in a bigon $D$ of $\sigma^{\prime}$ and $\sigma$ bounded by $s\cup t$. In the universal cover of the annulus $A_{\sigma}$, consider lifts $\widetilde{s}$, $\widetilde{t}$ of $s$ and $t$ whose union bounds a lift $\widetilde{D}$ of $D$, and the lift $\widetilde{c}^{\prime}$ of $c^{\prime}$ included in $\widetilde{D}$. Each lift of $\tau$ that meets $\widetilde{c}^{\prime}$ must also meet $\widetilde{t}$, since it is a connected set that meets both $\widetilde{D}$ and its complement. Thus we have
$\mathrm{tw}_{\sigma}(\tau,\tau^{\prime})=\mathrm{wd}_{\sigma}(\tau,c^{\prime})\leq\mathrm{wd}_{\sigma}(\tau,\sigma^{\prime}).$
Now we apply the ”width is a twist” Lemma 6.9 to find an annulus $A$ included in $A_{\sigma}$ such that $\mathrm{wd}_{\sigma}(\tau,\sigma^{\prime})=\mathrm{tw}_{A}(\tau,\sigma^{\prime})$, and thus satisfies the wanted inequality on twists.
It remains to modify $\tau^{\prime}$ to get a curve in the same isotopy class as $\tau$. Since $\tau$ is disjoint from a small enough tubular neighborhood of $\sigma$ on one side of $\sigma$, we can modify it anyhow in this neighborhood while keeping a common dual to $\sigma,\tau$. We do such a modification in order to change $\tau^{\prime}$ into a curve isotopic to $\tau$. This can be done, in the universal cover of $A_{\sigma}$, by modifying only the $c$ part of $\tau^{\prime}$, without crossing any lift of $\tau$ which is not already crossed by $c^{\prime}$. Thus the twist number $\mathrm{tw}_{\sigma}(\tau,\tau^{\prime})$ does not increase under the modification. This completes the proof. ∎
Proof of Proposition 7.9.
Let $\sigma,\sigma^{\prime}$ be as in Proposition 7.9. In order to evaluate the marking distance, we will use the distance $d_{w}$, which was defined by
$d_{w}(\{\alpha,\beta\},\{\alpha^{\prime},\beta^{\prime}\})=\mathrm{wd}_{\alpha}(\alpha^{\prime})+\mathrm{wd}_{\beta}(\beta^{\prime})$
(see Lemma 5.5). Note that $\mathrm{wd}_{\sigma}(\sigma^{\prime})\leq 3$ since they touch from both sides.
We claim that there is a curve $\gamma$ which is an “approximate dual” to both $\sigma,\sigma^{\prime}$. By this we mean that there are duals $\tau^{+},\tau^{-}$ to $\sigma,\sigma^{\prime}$ which are arbitrarily $C^{0}$-close to $\gamma$. To find such a curve, consider a dual $\tau$ to $\sigma$ which intersects $\sigma^{\prime}$ with no triple points. Then replace the segment after the first intersection with $\sigma^{\prime}$ until the last intersection with $\sigma^{\prime}$ with a subarc of $\sigma^{\prime}$.
$\sigma$
$\sigma^{\prime}$
$\tau$
$\gamma$
Figure 14. Constructing an approximate dual in Proposition 7.9
Now, let $m=\{\sigma,\tau^{+}\},m^{\prime}=\{\sigma^{\prime},\tau^{-}\}$ be the two markings for $\tau^{+},\tau^{-}$ close enough to $\gamma$. We then have that the relative width of $\tau^{-},\tau^{+}$ is at most $1$, and thus $d_{w}(m,m^{\prime})\leq 3+1=4$. By Lemma 5.5 this implies that a constant $C$ as in (2) exists.
For the twist bound (1) we note
$\mathrm{tw}_{\sigma}(\tau^{+},\sigma^{\prime})\leq\mathrm{tw}_{\sigma}(\tau^{+},\gamma)+\mathrm{tw}_{\sigma}(\gamma,\sigma^{\prime})\leq 2,$
and similar for the other case. ∎
8. The marking graph estimate
The aim of this section is to prove an upper bound for the distance $d(m,m^{\prime})$ in the marking graph inspired by foundational Masur-Minsky distance formula [MM00] in the “usual” marking graph.
This bound is in terms of (1) the distance between $m$ and $m^{\prime}$ in the fine curve graph, and (2) the sum over the twist numbers between $m$ and $m^{\prime}$ in suitable annuli.
Before we phrase the result, note that given two markings $m=\{\alpha,\beta\},m^{\prime}=\{\alpha^{\prime},\beta^{\prime}\}$, we have $d_{\mathcal{C}^{\dagger}}(\alpha,\beta)=d_{\mathcal{C}^{\dagger}}(\alpha^{\prime},\beta^{\prime})=1$ , and thus we may define $d_{\mathcal{C}^{\dagger}}(m,m^{\prime})=d_{\mathcal{C}^{\dagger}}(\alpha,\alpha^{\prime})$, up to an error of $2$.
Lemma 8.1 (Fine marking distance estimate).
There is a constant $K>0$ so that the following holds: Suppose that $m,m^{\prime}\in\mathcal{M}^{\dagger}$ are two markings. Then there is a number $N\leq d_{\mathcal{C}^{\dagger}}(m,m^{\prime})$ and curves $\delta_{i},i=0,\ldots,N$ so that
$\frac{1}{K}d_{\mathcal{M}^{\dagger}}(m,m^{\prime})-K\leq d_{\mathcal{C}^{\dagger}}(m,m^{\prime})+\sum_{i=0}^{N}\mathrm{tw}_{\delta_{i}}(m,m^{\prime}).$
The proof of the marking distance bound uses the following class of geodesics, which seem to be an analog of Masur-Minsky’s tight geodesics in our setting.
Definition 8.2.
A geodesic $(\alpha_{i}),i=0,\ldots,n$ in $\mathcal{C}^{\dagger}(T)$ is called steady if the number of indices $i$ so that $\alpha_{i}\cap\alpha_{i+1}=\emptyset$ is maximal among all geodesics joining $\alpha_{1}$ to $\alpha_{n}$.
For such a geodesic we define the disjointness intervals as the maximal intervals $I_{k}=[i,j]\cap\mathbb{Z}$ so that $\alpha_{r},\alpha_{r+1}$ are disjoint for all $r,r+1\in I_{k}$.
Note that every curve on a steady geodesic lies in a disjointness interval, which might have length 0.
The reason for the usefulness of these geodesics is the next lemma, which shows that these geodesics behave like tight geodesics in the usual curve graph: they ”don’t loose projection”. Remember that a curve $_{a}lpha$ is almost distance $k$ from a curve $\beta$ if it may be $C^{0}$-approximated by curves at distance at most $k$ from $\beta$. Note that if $\alpha$ is almost distance 0 from $\beta$ then $\alpha=\beta$.
Lemma 8.3.
Suppose that $(\alpha_{i}),i=0,\ldots,n$ is steady, $\beta_{0}$ is a dual to $\alpha_{0}$, $\beta_{n}$ is a dual to $\alpha_{n}$ and $I_{k}=[i,j]\cap\mathbb{Z}$ is a disjointness interval for $i>1,j<n$.
Suppose that $A$ is an annulus one of whose boundary curves $\delta$ has almost distance $k$ from $\alpha_{i}$ and almost distance (ji)k(j-i)-k from $\alpha_{j}$.
Then $\pi_{A}(\alpha_{r})\neq\emptyset$ for all $r$ with $r<i$ or $r>j$.
Proof.
We prove that $\pi_{A}(\alpha_{r})\neq\emptyset$ for $r<i$; the other case is similar. Thus, suppose that $\pi_{A}(\alpha_{r})=\emptyset$, and thus $d_{0}(\alpha_{r},\delta)\leq 2$. Since $\delta$ has almost distance (ji)k(j-i)-k from $\alpha_{j}$, and upper distance bounds are an open condition, we thus find that
$j-r=d(\alpha_{r},\alpha_{j})\leq d(\alpha_{r},\delta^{\prime})+d(\delta^{\prime},\alpha_{j})\leq 2+(j-i)-k\leq 2+j-i,$
where $\delta^{\prime}$ is a curve close enough to $\delta$. Hence, we obtain
$r\geq i-2+k.$
Since we assume $r<i$, this implies r=i1r=i-1 or r=i2r=i-2 (and, in this latter case, $k=0$).
Note that by the definition of disjointess intervals, $\alpha_{i-1}$ is in a different isotopy class than any curve inside the disjointess interval, and therefore $\pi_{A}(\alpha_{i-1})\neq\emptyset$.
Thus, the only remaining case is r=i2,k=0r=i-2,k=0, and thus $\pi_{A}(\alpha_{i-2})=\emptyset$ where a boundary of $A$ has almost distance 0 from $\alpha_{i}$, which means that $A=A_{\alpha_{i}}$. However, in this case, there is a curve $\gamma$ which is disjoint from $\alpha_{i},\alpha_{i-2}$. Thus, replacing $\alpha_{i-1}$ by $\gamma$ yields a geodesic between $\alpha_{1},\alpha_{n}$ with more edges corresponding to disjointness – contradicting the definition of steadyness. ∎
Combining this lemma with the main result of the previous section, we obtain the following crucial estimate.
Corollary 8.4.
There is a constant $K>0$ so that the following holds. Suppose that $\alpha_{k},k=0,\ldots,N$ is a steady geodesic, and $\alpha_{-1}$ is dual to $\alpha_{0}$, and $\alpha_{N+1}$ is dual to $\alpha_{N}$.
Suppose that $I=[i,j]$ is a disjointness interval of the steady geodesic. Then $m_{I}^{-}=(\alpha_{i},\alpha_{i-1}),m_{I}^{+}=(\alpha_{j},\alpha_{j+1})$ are markings and there are annuli $A_{k},k=0,\ldots,j-i$ so that
$d(m_{I}^{-},m_{I}^{+})\leq K(j-i)+2\sum_{k=0}^{j-i}\mathrm{tw}_{A_{k}}(\alpha_{-1},\alpha_{N+1})$
Proof.
The fact that $m_{I}^{-},m_{I}^{+}$ are markings is immediate from the definition of disjointness interval and the choice of $\alpha_{-1},\alpha_{N+1}$. To obtain the estimate, apply Proposition 7.3 to $\alpha=\alpha_{i},\beta=\alpha_{i-1},\alpha^{\prime}=\alpha_{j},\beta^{\prime}=\alpha_{j+1},n=j-i$ to obtain markings $m_{k}^{\pm}$ and annuli $A_{k}$ whose boundary have almost distance $k$ from $\alpha$ and almost distance nkn-k from $\alpha^{\prime}$ so that
[…]
$\leq$ $\sum_{k=0}^{n}d(m_{k}^{-},m_{k}^{+})+\sum_{k=0}^{n-1}d(m_{k}^{+},m_{k+1}^{-})$
We now use the estimates (2) to (7) from Proposition 7.3. By (2) the second sum is bounded by $nC$. The estimates (3) to (7) provide a bound for each term of the first sum by one or two twist numbers and a constant. Putting this together, we get
$d(m_{I}^{-},m_{I}^{+})$ $\leq$ $2\mathrm{tw}_{A_{0}}(\beta,\beta^{\prime})+\mathrm{tw}_{A_{1}}(\beta,\beta^{\prime})+\sum_{k=2}^{n-2}(\mathrm{tw}_{A_{k}}(\beta,\beta^{\prime})+2B+6)$
$+\mathrm{tw}_{A_{n-1}}(\beta,\beta^{\prime})+2\mathrm{tw}_{A_{n}}(\beta,\beta^{\prime})+6B+10+nC$
$\leq$ $(C+3B+6)n+2\sum_{i=0}^{n}\mathrm{tw}_{A_{k}}(\beta,\beta^{\prime}).$
Now, by the previous Lemma 8.3, we have that $\pi_{A_{k}}(\alpha_{k^{\prime}})\neq 0$ for all $k^{\prime}<i,k^{\prime}>j$ and $k=0,\dots,n$. Thus, by the bounded geodesic image theorem and the triangular inequality for twists we have for every $i$ that
$\mathrm{tw}_{A_{k}}(\beta,\beta^{\prime})\leq\mathrm{tw}_{A_{k}}(\alpha_{-1},\alpha_{N+1})+2B.$
Together, these imply the lemma. ∎
We are now ready for the proof of the main estimate.
Proof of Lemma 8.1.
Let $m=(\alpha,\beta)$ and $m^{\prime}=(\alpha^{\prime},\beta^{\prime})$ be given. We begin by choosing a steady geodesic $(\alpha_{i}),i=1,\ldots,N$ between $\alpha=\alpha_{1}$ and $\alpha^{\prime}=\alpha_{n}$. Denote by $I_{1},\ldots,I_{K}$ the disjointness intervals, and put $I_{k}=[i_{k},j_{k}]$. Let $m_{k}^{\pm}$ be the markings obtained from Corollary 8.4 applied to $I_{k}$. Note that by construction for any $k$ we have $m_{k}^{+}=m_{k+1}^{-}$, and thus
$d(m,m^{\prime})\leq\sum d(m_{k}^{-},m_{k}^{+}).$
Summing the estimate from Corollary 8.4 then gives the desired formula, noting that the sum of the lengths of the disjointness intervals is the length of the steady geodesic. ∎
9. Metric WPD property, $C^{0}$ distance estimates
In this short section we prove the metric WPD property, by mixing all our ingredients: the twist bound along axes (Lemma 6.7), the marking distance estimate in terms of twists (Lemma 8.1), and the bound on $\tilde{d}$ from marking distance (Corollary 5.3).
9.1. A strong version
For clarity we prove that the metric WPD property is equivalent to another, seemingly stronger, property. We do not really need this lemma here since in our context the proof of the stronger version will not be more difficult. We consider a topological group $G$ acting by isometries on a Gromov-hyperbolic metric space $X$.
Lemma 9.1.
A loxodromic element $f$ satisfies the metric WPD property if and only if for every $x\in X$ and every $r>0$, there exists $n_{0}$ and $D$ such that for every $n\geq n_{0}$, the set $Z_{r}(x,f^{n}x)$ is bounded by $D$ for the maximal metric in $G$.
Proof.
The reverse implication is clear, so let us prove the direct one. Assume $f$ satisfies the metric WPD property. Let $x\in X$ and $r>0$, and let us choose an invariant quasigeodesic path $A$ for $f$ containing $x$. Let $M$ be the associated Morse constant, and let $M^{\prime}=5r+4\delta+4M$ be the constant of Lemma 3.3. The metric WPD property gives a number $n_{0}\geq 0$ such that the set $Z_{M^{\prime}}(x,f^{n_{0}}x)$ is bounded by some $D$ for the maximal metric. Let $n\geq n_{0}$. Now the proof will be complete if we show that $Z_{r}(x,f^{n}x)\subseteq Z_{M^{\prime}}(x,f^{n_{0}}x)$.
To see this, let $h\in Z_{r}(x,f^{n}x)$, that is, we have $d(x,hx)\leq r$ and $d(f^{n}x,hf^{n}x)\leq r$. Denote $\ell=d(x,f^{n_{0}}x)$ and observe that $\ell=d(hx,hf^{n_{0}}x)$. We are in the situation of Lemma 3.3, with $\alpha,\alpha^{\prime},\beta,\beta^{\prime},\hat{\alpha},\hat{\beta}$ respectively set to $x,f^{n}x,hx,hf^{n}x,f^{n_{0}}x,hf^{n_{0}}x$, and $k=\ell$. We get that $d(f^{n_{0}}x,hf^{n_{0}}x)\leq M^{\prime}$, that is, $h\in Z_{M^{\prime}}(x,f^{n_{0}}x)$. ∎
9.2. Proof of the metric WPD property
This subsection is dedicated to the proof of the metric WPD property (Proposition 1.3).
Lemma 9.2 (Twist bound).
Suppose $f\in\mathrm{Homeo}_{0}(T^{2})$ acts hyperbolically on the fine curve graph,and let $m_{0}=(\alpha,\beta)$ be a marking.
Then for every $r>0$, there exists $n_{0}$ and $T$ such that for every $n\geq n_{0}$, for every $h\in Z_{r}(\alpha,f^{n}\alpha)$ and every almost embedded annulus $A$ we have
$\mathrm{tw}_{A}(m_{0},hm_{0})\leq T.$
Proof.
$m_{0}$
$f^{n}(m_{0})$
$h(m_{0})$
$hf^{n}(m_{0})$
$\partial A$
Figure 15. Lemma 9.2
Let $m_{0}$ as in the statement, and let $r>0$. Complete $m_{0}$ into a path of markings which is invariant under $f$. Lemma 6.7 provides a bound $T$ for twists numbers along this path. Choose
$n_{0}>\frac{2r+2}{\mathrm{tl}(f)}$
so that for every $n\geq n_{0}$, $d_{\mathcal{C}}^{\dagger}(\alpha,f^{n}\alpha)>2r+2$. Let $h\in Z_{r}(\alpha,f^{n}\alpha)$, and $A$ be an annulus. Let $B$ be given by the Bounded Geodesic Image Theorem 6.5. Let us prove that $\mathrm{tw}_{A}(m_{0},hm_{0})\leq B+2T$. Assume $\mathrm{tw}_{A}(m_{0},hm_{0})>B$, otherwise there is nothing to do. Then the Bounded Geodesic Image Theorem says that a geodesic $\alpha,h\alpha]$ enters the $2$-neighborhood of a boundary curve of $A$. By the choice of $n_{0}$ and triangular inequality we get that any geodesic $f^{n}\alpha,hf^{n}\alpha]$ is disjoint from the $2$-neighborhood of the boundary curve of $A$, and so the twist number $\mathrm{tw}_{A}(f^{n}m_{0},hf^{n}m_{0})$ is bounded by $B$. We now apply the triangular inequality for twist numbers (Lemma 6.6, iii) to get
$\mathrm{tw}_{A}(m_{0},hm_{0})\leq\mathrm{tw}_{A}(m_{0},f^{n}m_{0})+\mathrm{tw}_{A}(f^{n}m_{0},hf^{n}m_{0})+\mathrm{tw}_{A}(hf^{n}m_{0},hm_{0}).$
The first term is bounded by $T$, the second by $B$. By topological invariance of the twists numbers, the last one equals $\mathrm{tw}_{h^{-1}A}(f^{n}m_{0},m_{0})$ and thus is also bounded by $T$. Thus we get the wanted bound. ∎
Proof of Proposition 1.3.
Let $f$ be an element of $\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically, $\alpha\in\mathcal{C}^{\dagger}(S)$, and $r>0$. We will prove the strong version of the metric WPD expressed in Lemma 9.1, that is, we look for some $n_{0}$ such that for every $n\geq n_{0}$, the set $Z_{r}(\alpha,f^{n}\alpha)$ is bounded.
Complete $\alpha$ to get a marking $m_{0}=(\alpha,\beta)$. Apply lemma 9.2 to get some $n_{0}$ and the twist bound $T$. Let $n\geq n_{0}$ and $h\in Z_{r}(\alpha,f^{n}\alpha)$. We want to bound the $\tilde{d}$ distance from the identity to $h$. We know that $d(\alpha,h\alpha)\leq r$, and for every annulus $A$ we have $\mathrm{tw}_{A}(m_{0},hm_{0})\leq T$. Lemma 8.1 provides a number $D=D(r,T)$ such that $d_{\mathcal{M}^{\dagger}}(m_{0},hm_{0})\leq D$. By properness of the inclusion $\mathcal{M}_{0}^{\dagger}\subset\mathcal{M}^{\dagger}$ (Corollary 5.7), and $C^{0}$ bound from markings (Corollary 5.3), we get some $R^{\prime}$, depending only on $D$, such that $\tilde{d}(h,\mathrm{Id})\leq R^{\prime}$. ∎
The proof of the semi-continuity theorem in the next section will make use of the following form of the metric WPD property.
Corollary 9.3.
Let $b>0$, let $f$ be an element of $\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically, and let $\alpha\in\mathcal{C}^{\dagger}(S)$. Then there exists $n_{0}>0$, $D>0$ such that for every $g\in\mathrm{Homeo}_{0}(S)$, for every $\beta\in\mathcal{C}^{\dagger}(S)$, for every $n\geq n_{0}$ and $m\in\mathbb{Z}$, if
$d(\alpha,\beta)\leq b,\quad d(f^{n}\alpha,g^{m}\beta)\leq b,\quad d(f^{2n}\alpha,g^{2m}\beta)\leq b$
then $\tilde{d}(f^{-n},g^{-m})\leq D$.
Note that here we do not assume that $g$ is loxodromic.
Proof.
In short, we get the result by showing that $h=g^{-m}f^{n}$ belongs to $Z_{2b}(\alpha,f^{n}\alpha)$, and then applying the WPD. Here are the details. Let $b,f,\alpha$ be as in the statement. By Proposition 1.3, $f$ satisfies the metric WPD property. Applying this to $\alpha$ and $b$ provides a positive integer $n_{0}$ and a positive number $D$. Let $g\in\mathrm{Homeo}_{0}(S)$, $\beta\in\mathcal{C}^{\dagger}(S)$, $n\geq n_{0}$, $m\in\mathbb{Z}$, and consider the map $h=g^{-m}f^{n}$. By the triangular inequality and the fact that $f$ and $g$ act by isometries, we deduce that $h\in Z_{2b}(\alpha,f^{n}\alpha)$: indeed,
$d(\alpha,h\alpha)\leq d(\alpha,\beta)+d(\beta,h\alpha)=d(\alpha,\beta)+d(g^{m}\beta,f^{n}\alpha)\leq 2b,$
$d(f^{n}\alpha,hf^{n}\alpha)\leq d(f^{n}\alpha,g^{m}\beta)+d(g^{2m}\beta,f^{2n}\alpha)\leq 2b.$
Then the metric WPD property in the form of Lemma 9.1 tells us that $\tilde{d}(f^{-n},g^{-m})=\tilde{d}(h,\mathrm{Id})\leq D$. ∎
10. Axes and rotation sets
In this section we prove our main results relating the axis of a loxodromic element and the rotation sets, namely (1) Each end-point of the axis determines the scaled rotation set (Theorem 3), (2) Two elements that are equivalent in the sense of Fujiwara have the same scaled rotation set (Theorem 4), and (3) The rotation set depends semi-continuously on each end-point (Theorem 10.2 below). Each result follows easily from the next one: we have already explained how to deduce Theorem 3 from Theorem 4 at the end of Section 3, and we explain below how to deduce Theorem 4 from the semi-continuity. We nevertheless include a direct proof of Theorem 3, firstly because we will need the main lemma of the proof to get the metric characterization of the axis end-points (in section 10.4) below), and secondly as a warm-up for the more difficult proof of the semi-continuity. For the convenience of the reader we also include a sketch of a direct proof for Theorem 4. Then we will prove the full semi-continuity. We end the section by proving the metric characterization of the attracting fixed points for loxodromic elements.
10.1. Each axis end-point determines the rotation set
Let us recall Theorem 3 from the introduction.
Theorem.
Consider $f,g\in\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically on the fine curve graph, and assume $\xi^{+}(f)=\xi^{+}(g)$. Then the scaled rotation sets
$\frac{1}{\mathrm{tl}(f)}\mathrm{Rot}(f)\text{ and }\frac{1}{\mathrm{tl}(g)}\mathrm{Rot}(g)$
are translates of each other.
For $f$ as in the theorem, the scaled rotation set $\mathrm{SRot}(f)$ of $f$ is defined as an element of the set $\mathcal{K}(\mathbb{R}^{2})$ of compact convex subsets of the plane up to translations. The theorem says that the map $\mathrm{SRot}$ factorizes through $\xi^{+}$ and a map $\mathrm{srot}:X\to\mathcal{K}(\mathbb{R}^{2})$, where $X$ is the subset of the points in $\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})$ that are fixed points of some loxodromic element. In sub-section 10.3 we will explore the continuity of the map $\mathrm{srot}$ with respect to the distance induced by the Hausdorff distance on the set $\mathcal{K}(\mathbb{R}^{2})$, for the usual semi-continuity associated with the Hausdorff distance.
Proof.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ as in the theorem, with $\xi^{+}(f)=\xi^{+}(g)$. We begin by proving the following boundedness property in $\mathrm{Homeo}_{0}(T^{2})$.
Lemma 10.1.
There exists a sequence $(m_{n})$ tending to $\infty$, such that the sequences
$(d_{\mathcal{C}^{\dagger}}(f^{n}\alpha_{0},g^{m_{n}}\alpha_{0}))_{n\geq 0},\ \ (\widetilde{d}(f^{-n},g^{-m_{n}}))_{n\geq 0}$
are bounded.
Proof.
Fix some $\alpha_{0}\in\mathcal{C}^{\dagger}(T^{2})$. The hypothesis $\xi^{+}(f)=\xi^{+}(g)$ exactly means that the Gromov product
$(f^{n}(\alpha_{0}),g^{m}(\alpha_{0}))_{\alpha_{0}}$
tends to $\infty$ when $n,m$ tends to $\infty$. We apply Lemma 3.5 to the quasigeodesic axes $(f^{2n}(\alpha_{0}))_{n\in\mathbb{Z}}$, $(g^{2n}(\alpha_{0}))_{n\in\mathbb{Z}}$. This provides some sequence $(m_{n})_{n\geq 0}$ of positive numbers such that the sequence $(d_{\mathcal{C}^{\dagger}}(f^{2n}(\alpha_{0}),g^{2m_{n}}(\beta_{0})))_{n\geq 0}$ is bounded by some number $b$.
Corollary 3.4 provides some $M^{\prime}$ so that for every $n$, $d_{\mathcal{C}^{\dagger}}(f^{n}\alpha_{0},g^{m_{n}}\alpha_{0})\leq M^{\prime}$, which provides our first bound. We now apply the metric WPD property as expressed by Corollary 9.3, with $b^{\prime\prime}=\max(b,M^{\prime})$. We get some $D$ and some $n_{0}$, depending on $b,\alpha_{0},f$, so that, assuming $n\geq n_{0}$, we get $\tilde{d}(f^{-n},g^{-m_{n}})\leq D$. ∎
We can now complete the proof of the theorem. We first note that since $\mathrm{Rot}(f^{-1})=-\mathrm{Rot}(f)$ and $\mathrm{tl}(f^{-1})=\mathrm{tl}(f)$, it suffices to prove that the scaled rotation sets of $f^{-1}$ and $g^{-1}$ coincides. Let $(m_{n})$ be the sequence given by the lemma: the sequence $(d_{\mathcal{C}^{\dagger}}(f^{n}\alpha_{0},g^{m_{n}}\alpha_{0}))_{n\geq 0}$ is bounded. From the definition of translation lengths it follows easily that
$\lim\frac{m_{n}}{n}=\frac{\mathrm{tl}(f)}{\mathrm{tl}(g)}.$
Denote this limit by $\lambda$. Since, by the above lemma, the sequence $(f^{-n}g^{m_{n}})$ is bounded in the topological group $\mathrm{Homeo}_{0}(T^{2})$, we may apply the ”boundedness and the rotation” Lemma 2.7, ii) to the maps $f^{-1}$ and $g^{-1}$. We conclude that the sets $\mathrm{Rot}(f^{-1})$ and $\lambda\mathrm{Rot}(g^{-1})$ are translates of each other. This completes the proof. ∎
10.2. The Bestvina-Fujiwara class determines the rotation set
The above proof of Theorem 3 may be adapted to prove Theorem 4. We now sketch such a proof, for a full indirect proof see the end of the next subsection.
Sketch of proof of Theorem 4.
Let $f,g\in\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically on the fine curve graph. Assume $f\sim g$: there exists a sequence $(g_{k})=(h_{k}gh_{k}^{-1})$ of maps conjugate to $g$ such that $\lim\xi^{-}(g_{k})=\xi^{-}(f),\lim\xi^{+}(g_{k})=\xi^{+}(f)$. Choose some axes $A=(f^{n}\alpha_{0})$ for $f$, and $B=(g^{n}\beta_{0})$ for $g$. Let $n\geq 0$. If $k$ is large enough, then the end-points of $h_{k}(B)$ are close to that of $A$, and thus $h_{k}(B)$ fellow travels with the part of $A$ from $\alpha_{0}$ to $f^{2n}\alpha_{0}$. As before, using Corollary 3.4 and 9.3, we get a bound $D$ (not dependent on $n$) on the distance $\widetilde{d}(f^{-n},g_{k_{n}}^{-m_{n}})$. Thus we get the following property, which plays the role of Lemma 10.1 in the previous proof:
$\alpha_{0}$
$f^{2n}(\alpha_{0})$
AA
$f^{n}(\alpha_{0})$
$h\beta_{0}$
$hg^{2n}(\beta_{0})$
hBhB
$hg^{n}(\beta_{0})$
Figure 16. Theorem 4
There are sequences of positive integers $(k_{n})_{n\geq 0}$, $(m_{n})_{n\geq 0}$ such that the sequence $\widetilde{d}((f^{-n},g_{k_{n}}^{-m_{n}})_{n\geq 0}$ is bounded by some $D$. Furthermore, $\lim\frac{m_{n}}{n}=\frac{\mathrm{tl}(f)}{\mathrm{tl}(g)}$.
Now let $\varepsilon>0$. With the emphasized property we can apply Lemma 2.7, i), and we get some $n_{0}$ so that for every $n\geq n_{0}$, the set
$\frac{1}{n}\mathrm{Rot}(g_{k_{n}}^{-m_{n}})=\frac{m_{n}}{n}\mathrm{Rot}(g^{-1})$
has a translate included in the $\varepsilon$-neighborhood of $\mathrm{Rot}(f^{-1})$. Letting $n$ goes to $\infty$, and then $\varepsilon$ to 0, we get that $\lambda\mathrm{Rot}(g^{-1})$ has a translate included in $\mathrm{Rot}(f^{-1})$, where $\lambda$ denotes as above the quotient of the translation lengths. We get the reverse inclusion by symmetry of the hypotheses. This completes the proof. ∎
10.3. Semi-continuity
We now turn to the more general version.
Theorem 10.2.
Let $f\in\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically on the fine curve graph. Let $(g_{k})$ be a sequence in $\mathrm{Homeo}_{0}(T^{2})$ of elements acting loxodromically such that
  1. (1)
    the sequence $(\mathrm{tl}(g_{k}))$ is included in $\eta_{0},\eta_{1}]$ for some $0<\eta_{0}<\eta_{1}$;
  2. (2)
    $(\xi^{-}(g_{k}))_{k\geq 0}$ has no subsequence converging to $\xi^{+}(f)$;
  3. (3)
    $(\xi^{+}(g_{k}))_{k\geq 0}$ converges to $\xi^{+}(f)$.
Then for every $\varepsilon>0$ and every $k$ large enough, some translate of $\frac{1}{\mathrm{tl}(g_{k})}\mathrm{Rot}(g_{k})$ is included in the $\varepsilon$-neighborhood of $\frac{1}{\mathrm{tl}(f)}\mathrm{Rot}(f)$.
Let us make explicit the semi-continuity property. Let $\mathcal{N}$ be any neighborhood of the diagonal in $\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})^{2}$, and define $\mathcal{X}_{\mathcal{N}}$ to be the set of $\xi\in\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})$ such that there exists $f\in\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically such that $\xi^{+}(f)=\xi$ and $(\xi^{-}(f),\xi^{+}(f))\not\in\mathcal{N}$. Then the theorem says that the map $\mathrm{srot}$, defined in section 10.1, is semi-continuous when restricted to the set $\mathcal{X}_{\mathcal{N}}$.
We are currently unable to prove a full continuity result; in other words:
Question 10.3.
Under the hypotheses of Theorem 10.2, do we have, up to translation,
$\lim\frac{1}{\mathrm{tl}(g_{k})}\mathrm{Rot}(g_{k})=\frac{1}{\mathrm{tl}(f)}\mathrm{Rot}(f)?$
Compared to the proofs of Theorems 3 and 4, which essentially involved only the two maps $f$ and $g$, the main difficulty here is that we need to control the quality of the quasigeodesic axes for the $g_{k}$’s. This will be achieved via Lemma 3.6.
Proof of theorem 10.2.
Let $f$ and $(g_{k})$ be as in the statement of the theorem. Since all translation lengths are assumed to be larger than $\eta_{0}$, up to changing $f$ to $f^{N}$ for some large $N$, and each $g_{k}$ to $g_{k}^{N}$, which enlarge all involved rotation sets by the same factor $N$, we may assume that $\mathrm{tl}(f)$ and all $\mathrm{tl}(g_{k})$’s are larger than the constant $L(\delta)$ from Lemma 3.6. The lemma provides a constant $K=K(\delta)$, and invariant paths
[…]
which are all $K$-quasigeodesics. We also observe that by Corollary 3.7, every $\beta_{i}^{(k)}$ is distance at most $B=B(\delta,K,\eta_{1})$ from some $g_{k}^{2m}(\beta_{0}^{(k)})$.
By hypotheses (2) and (3), the sequence $((\xi^{-}(g_{k}),\xi^{+}(g_{k}))_{\alpha_{0}})_{k\geq 0}$ is bounded: indeed, otherwise there is a subsequence of $(g_{k})$ such that the Gromov product tends to $\infty$, and since, by (3), the sequence $(\xi^{+}(g_{k}))$ tends to $\xi^{+}(f)$ we would have a subsequence of $(\xi^{-}(g_{k}))$ also converging to $\xi^{+}(f)$, contrary to (2). Using Morse lemma, we get some $b_{0}$ such that for every $k$, the quasigeodesic $(\beta_{i}^{(k)})_{i\in\mathbb{Z}}$ enter the $b_{0}$-neighborhood of $\alpha_{0}$. Up to shifting the indices, we may (and will) assume that $d(\alpha_{0},\beta_{0}^{k})\leq b^{\prime}$ for every $k$.
Let $b=b(\delta,K,b_{0})$ be given by Lemma 3.5. Fix some $n\geq 0$, and let $N$ be such that $\alpha_{N}=f^{2n}\alpha_{0}$. Applying the lemma with our $N$, we get some $L$. By the convergence hypothesis (3), for $k$ large enough we have $(\xi^{+}(g_{k}),\xi^{+}(f))_{\alpha_{0}}>L$. Let us consider such a $k$. Then $(\alpha_{i_{0}},\beta^{k}_{i_{0}})_{\alpha_{0}}\geq L$ for every large enough $i_{0}$. Now Lemma 3.5 provides some point on the quasigeodesic $(\beta_{i}^{(k)})$ which is distance no more than $b$ from $f^{2n}\alpha_{0}$. By the above observation from Corollary 3.7, up to increasing $b$ by $B$, we may assume that this point is $g_{k}^{2m}\beta^{k}_{0}$ for some $m=m_{k,n}$, and we get
$d(f^{2n}\alpha_{0},g_{k}^{2m}\beta^{k}_{0})\leq b.$
We now apply Corollary 3.4 and get some $M^{\prime}=M^{\prime}(\delta,K,\max(b,b^{\prime}))$ so that $d(f^{n}\alpha_{0},g_{k}^{m}\beta^{k}_{0})\leq M^{\prime}$. Now Corollary 9.3 may be applied with $b^{\prime\prime}=\max(b,b^{\prime},M^{\prime})$ and provides some $D$ and some $n_{0}$, depending on $b,\alpha_{0},f$, so that, assuming $n\geq n_{0}$, we get $\tilde{d}(f^{-n},g_{k}^{-m})\leq D$. For clarity, let us summarize what we have proved until now.
Claim 10.4.
There is some $n_{0}$ and $D$ such that for every $n\geq n_{0}$, there exists $k_{n}$ such that for every $k\geq k_{n}$, there is some $m_{k,n}$ such that
$\tilde{d}(f^{-n},g_{k}^{-m_{k,n}})\leq D.$
Now let $\varepsilon>0$, and $n_{0}$ and $D$ as above. Let $n\geq n_{0}$ and $k\geq k_{n}$ as in the claim. Since $\widetilde{d}(f^{-n},g_{k}^{-m_{n,k}})\leq D$, we may apply Lemma 2.7, part i), with $f^{-1}$ instead of $f$ and $g=g_{k}^{-m_{k,n}}$. Remembering that $\mathrm{Rot}(f^{-1})=-\mathrm{Rot}(f)$, we get that up to enlarging $n_{0}$,
$\frac{m_{n,k}}{n}\mathrm{Rot}(g_{k})\subseteq N_{\varepsilon}(\mathrm{Rot}(f)).$
To get the inclusion asserted by the Theorem, it remains to see that if $n$ is large enough then the fraction on the left-hand side is arbitrarily close to $\frac{\mathrm{tl}(f)}{\mathrm{tl}(g_{k})}$, uniformly in $k$. Let us check this last fact. On the one hand, since all our quasigeodesic axes come from Lemma 3.6, the second part of the lemma yields
(1) $d(\alpha_{0},f^{n}\alpha_{0})$ $=$ $n\mathrm{tl}(f)+\Delta$
(2) $d(\beta^{(k)}_{0},g_{k}^{m_{k,n}}\beta^{(k)}_{0})$ $=$ $m_{k,n}\mathrm{tl}(g_{k})+\Delta^{\prime}$
with $|\Delta|,|\Delta^{\prime}|$ bounded by the Morse constant $M(K)$ from the lemma. On the other hand, we have
$d(\alpha_{0},\beta_{0}^{k})\leq b^{\prime},\ \ d(f^{n}\alpha_{0},g_{k}^{m_{k,n}}\beta^{(k)}_{0})\leq M^{\prime},$
and thus (1) and (2) differ by at most $M^{\prime}+b^{\prime}$. This gives the wanted convergence. ∎
Proof of Theorem 4 from Theorem 10.2.
Assume $f\sim g$: there exists a sequence $(g_{k})$ of maps conjugate to $g$ such that $\lim\xi^{-}(g_{k})=\xi^{-}(f)$, $\lim\xi^{+}(g_{k})=\xi^{+}(f)$. Note that the $g_{k}$’s, begin conjugate to $g$, have the same translation length and the same rotation set. We may apply Theorem 10.2 to $(g_{k})$. Thus we get that $\mathrm{Rot}(g)$ has some translate included in
$\frac{\mathrm{tl}(g)}{\mathrm{tl}(f)}\mathrm{Rot}(f)$
Exchanging the roles of $f$ and $g$ gives the reverse inclusion. We conclude by noting that if each of both compact subsets $C,C^{\prime}$ has some translate included in the other one, then some translate of $C^{\prime}$ is equal to $C$. ∎
10.4. Metric characterization of fixed points of loxodromic
The following statement contains Corollary 1.4 from the introduction.
Proposition 10.5.
Let $f,g$ be two elements of $\mathrm{Homeo}_{0}(T^{2})$ acting loxodromically on the fine curve graph. Then the following are equivalent.
  1. (1)
    $\xi^{+}(f)=\xi^{+}(g)$,
  2. (2)
    There exists a sequence $(m_{n})$ tending to $\infty$, such that the sequence
    $(f^{-n}g^{m_{n}}))_{n\geq 0}$
    is bounded in $\mathrm{Homeo}_{0}(T^{2})$,
  3. (3)
    There exist sequences $(n_{i}),(m_{i})$ tending to $\infty$, such that the sequence
    $(f^{-n_{i}}g^{m_{i}}))_{i\geq 0}$
    is bounded in $\mathrm{Homeo}_{0}(T^{2})$.
Note that the formulation of the logical equivalence between (2) and (3) does not involve the fine curve graph, but we are not aware of any argument that does not go through property (1).
Proof.
Let $f,g$ be loxodromic elements. The implication (1) $\Rightarrow$ (2) is exactlty Lemma 10.1. The implication (2) $\Rightarrow$ (3) is obvious, so it remains to prove that (3) implies (1). We first make the following observation.
Claim 10.6.
Let $h\in\mathrm{Homeo}_{0}(T^{2})$, and $\alpha_{0}$ denote a horizontal curve in $T^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. Then
$d_{\mathcal{C}^{\dagger}(T^{2})}(h\alpha_{0},\alpha_{0})\leq 2\widetilde{d}(h,\mathrm{Id})+2.$
Proof of the claim.
Let $D=\widetilde{d}(h,\mathrm{Id})$, and let us fix a lift $\widetilde{\alpha}_{0}$ of $\alpha_{0}$ in the universal cover. From the hypothesis it follows that $h\alpha_{0}$ has a lift which is between $\widetilde{\alpha}_{0}-(0,D)$ and $\widetilde{\alpha}_{0}+(0,D)$. This lifts meets at most $2D+1$ lifts of $\alpha_{0}$. Now the claim follows from [BHM+22, Lemma 4.5]. ∎
Assume (3), and let $(m_{i}),(n_{i}),D$ be such that for every $i$, $\widetilde{d}(f^{-n_{i}}g^{m_{i}},\mathrm{Id})\leq D$. Let $\alpha_{0}$ be a horizontal simple closed curve in the torus. Let $i\geq 0$. By the claim and the fact that $f$ acts by isometry on the fine curve graph, we get that the distance in $\mathcal{C}^{\dagger}(T^{2})$ between $f^{n_{i}}\alpha_{0}$ and $g^{m_{i}}\alpha_{0}$ is bounded by $2D+2$. This implies at once that $\xi^{+}(f)=\xi^{+}(g)$. ∎
Note that the proof also shows, under hypothesis (3), that the quotient $m_{i}/n_{i}$ tends to the ratio of the translation lengths.
11. (Counter)examples
In this section we describes examples of torus homeomorphisms that illustrate our results. In section 11.1 we give formulas for homeomorphisms having a given parallelogram as a rotation set. For triangles we are not aware of any formula, thus instead we provide a geometric construction in the spirit of Kwapisz [Kwa92], in section 11.2. In section 11.3 we define the twist numbers between two points at infinity, and the twist number of a loxodromic element, and prove some elementary properties. We illustrate this notion in section 11.4 by providing an example of computation for a nice family of examples. In section 11.5 we describe a free subgroup of $\mathrm{homeo}_{0}(T^{2})$ made of loxodromic elements, whose scaled rotation sets behave as badly as possible: some boundary points in $\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})$ have no associated (scaled) rotation sets, and the scaled rotation set associated to the attracting fixed points of loxodromic element does not depend continuously on the element.
11.1. Parallelograms
Lemma 11.1.
Every parallelogram with rational vertices is the rotation set of a homeomorphism $f$ some power of which is conjugate to its inverse, and thus such that $\mathrm{scl}(f)=0$.
Proof.
We first explain how to realize the square $0,1]^{2}$ by a homeomorphism $h$ which is conjugate to its inverse, from which we will deduce the general case. Let $\tilde{f}(x,y)=(x+\sin^{2}(\pi y),y)$, $\tilde{g}(x,y)=(x,y+\sin^{2}(\pi x))$, $\tilde{h}=\tilde{g}\tilde{f}$. Then the rotation set of $\tilde{h}$ is the square $0,1]^{2}$ (see for instance [LPSW24]). Furthermore, denoting $T_{u}$ the translation by vector $u$, we note that the map $T_{(\frac{1}{2},0)}$ commutes with $\tilde{f}$ and conjugate $\tilde{g}$ to $(x,y+1-\sin^{2}(\pi x))$. Denoting $T=T_{(\frac{1}{2},\frac{1}{2})}$, we get
$T\tilde{h}T^{-1}=T_{(1,1)}\tilde{h}^{-1}.$
Hence the torus homeomorphism $h$ induced by $\tilde{h}$ is conjugate to its inverse under the translation $T$.
Now let us deduce the general case. Any rational parallelogram $P$ is the image of the square $0,1]^{2}$ under a rational affine transformation $\Phi$. Then Lemma 2.3 provides a torus homeomorphism $h^{\prime}$ whose rotation set is $P$, with a lift of the form $\tilde{h}^{\prime}=T_{v}A^{-1}\tilde{h}^{p}A$, where $\tilde{h}$ is as before, for some rational vector $v$ and integer matrix $A$. Denoting $T^{\prime}=A^{-1}TA$, we get $T^{\prime}\tilde{h}^{\prime}T^{\prime-1}=T_{A^{-1}(p,p)}\tilde{h}^{\prime-1}$. If we choose $q$ such that the matrix $pqA^{-1}$ has integer coefficients, then the map $T^{\prime}$ conjugate $h^{\prime q}$ to its inverse.
Let us now explain the well-known fact that a map $h$ having some power conjugate to its inverse has vanishing stable commutator length. Since $\mathrm{scl}(h^{p})=p\mathrm{scl}(h)$, it suffices to show that the relation $h=\Phi h^{-1}\Phi^{-1}$ implies $\mathrm{scl}(h)=0$. From this relation we get $h^{n}=\Phi h^{-n}\Phi^{-1}$, and multiplying by $h^{n}$ displays $h^{2n}$ as the commutator of $h^{-n}$ and $\Phi$. In particular $\mathrm{cl}(h^{2n})=1$, and thus $\mathrm{scl}(h)=0$. ∎
Motivated by this, we ask
Question 11.2.
Is every symmetric rational polygon realized as the rotation set of a torus homeomorphism which has a power conjugate to its inverse?
Up to our knowledge, there is no known example of a rotation set which is a finite-sided polygon that is not rational.
11.2. Triangles
We now describe a torus homeomorphism (which can be realised as a $C^{\infty}$-diffeomorphism) whose rotation set is the triangle $\mathcal{T}$ with vertices $(0,0)$, $(1,0)$, $(0,1)$. From this particular example we get all rational triangles by applying Lemma 2.3.
To get our example we follow the construction by Kwapisz ([Kwa92]), as exposed by Béguin in [Bé07], except that we do not require $f$ to be Axiom A (we observe that Markov partitions are not needed for computing the rotation set). We include the construction since the more general construction of Kwapisz, which allows him to realize every rational polygon, is much more difficult to follow.
$R$
$H$
$V$
AA
SHSH
TVTV
$\widetilde{f}V$
$\widetilde{f}H$
Figure 17. Construction of $f$
Let us partition the square $0,1]^{2}$ into four squares $\widetilde{H},\widetilde{A},\widetilde{R},\widetilde{V}$ as on Figure 17, and $H,A,V,R$ their projection in the torus $\mathbb{R}^{2}/\mathbb{Z}^{2}$. We want to define $\widetilde{f}\in\mathrm{Homeo}(\mathbb{R}^{2},\mathbb{Z}^{2})$ such that, roughly speaking, $A$ will be an attractor, $R$ a repulsor, points in $H$ will go to the right, and points of $V$ will go up. More precisely, denote $S,T$ the unit horizontal and vertical translations respectively. We demand the following condition (see Figure 17):
  1. (1)
    $\widetilde{f}(\widetilde{A})\subset\operatorname{Inte}\widetilde{A}$,
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
[…]
  1. (3b)
    $f$ has a fixed point in $H$ whose rotation vector is $(1,0)$,
  2. (4b)
    $f$ has a fixed point in $V$ whose rotation vector is $(0,1)$.
With these conditions, by convexity the rotation set of $\widetilde{f}$ contains the triangle $\mathcal{T}$. It remains to check that it is also included in $\mathcal{T}$. Let $x\in T^{2}$, and consider the positive orbit $(f^{n}(x))_{n\geq 0}$. If it is entirely included in $R$, then it has rotation vector 0. If it enters $A$ then it is trapped in $A$, and again it has rotation vector 0. Let us examine the remaining case, namely when the orbit is not entirely in $R$ and is disjoint from $A$. Since $R$ is a repulsor, up to changing $x$ to some iterate we may assume that the orbit is included in $H\cup V$. The following observation is not crucial but makes the argument easier: we have
$TV=H+\left(\frac{1}{2},\frac{1}{2}\right),SH=V+\left(\frac{1}{2},\frac{1}{2}\right).$
Now if $y$ is a point of $C=H\text{ or }V$ whose image $f(y)$ also belongs to $H\cup V$, then the image of a lift $\widetilde{y}$ belongs to some translate $\widetilde{C}$ of $\widetilde{H}$ or $\widetilde{V}$, and according to properties (3) and (4),
$\widetilde{f}(\widetilde{y})\in\widetilde{C}+v,\text{ with }v\in E:=\left\{(0,0),(1,0),(0,1),\left(\frac{1}{2},\frac{1}{2}\right)\right\}.$
Choose a lift $\widetilde{x}$ of $x$ in $\widetilde{C}=\widetilde{H}\text{ or }\widetilde{V}$. An immediate induction yields, for every $n\geq 0$,
$\widetilde{f}^{n}(\widetilde{x})\in C+nv$
for some $v$ which is a convex combination of elements of $E$, and thus also of the three vertices of the triangle $\mathcal{T}$. We deduce that the $\mathrm{Rot}(\widetilde{f})\subset\mathcal{T}$, and conclude that $\mathrm{Rot}(\widetilde{f})=\mathcal{T}$.
11.3. Twists numbers for boundary points, and for maps
In this subsection we work in the fine curve graph $\mathcal{C}^{\dagger}(S)$ of any surface $S$, with the only assumption that this graph is unbounded and hyperbolic. We will use twist numbers to define a new conjugacy invariant for loxodromic maps, and thus in particular for elements of our group $\mathrm{Homeo}_{0}(T^{2})$ with rotation set having non-empty interior. Elements with sufficiently distinct twist numbers are not Bestvina-Fujiwara-equivalent, and thus this also provides a tool for getting positive scl. In the next subsection we will compute this invariant on relevant examples, and show that it is not a function of the rotation set and the translation length.
Let us denote $r\geq 2$ the constant given by the first Bounded Geodesic Image Theorem, Lemma 6.4, for geodesics: for every curve $\alpha$, for any two points $x,y$ that may be joined by a geodesic that does not enter the $r$-ball around $\alpha$, the diameter of $\pi_{\alpha}(x)\cup\pi_{\alpha}(y)$ in the arc graph is at most $2$, where $\pi_{\alpha}$ denotes the projection in the annulus $A_{\alpha}$. Furthermore, none of the projections $\pi_{\alpha}(x),\pi_{\alpha}(y)$ is empty. This implies that every arc in $\pi_{\alpha}(x)$ is distance at most two of every arc in $\pi_{\alpha}(y)$. Thus the number $\mathrm{tw}_{\alpha}(x,y)$ is well defined and no more than $2$. Note that this situation happens in particular if the Gromov product $(x\cdot y)_{\alpha}$ is larger than $r$ (see section 3.1).
We now fix two boundary points $\xi,\xi^{\prime}\in\partial\mathcal{C}^{\dagger}(S)$. Let $(x_{i}),(x^{\prime}_{i})$ be admissible sequences that respectively converge to $\xi,\xi^{\prime}$. Fix another curve $\alpha$. We can find $i_{0}$ such that for all $i,j\geq i_{0}$, $(x_{i}\cdot x_{j})_{\alpha}\geq r$, and $(x^{\prime}_{i}\cdot x^{\prime}_{j})_{\alpha}\geq r$. Thus
  1. (1)
    $\forall i,j\geq i_{0}$, $\mathrm{tw}_{\alpha}(x_{i},x_{j})\leq 2$ and $\mathrm{tw}_{\alpha}(x^{\prime}_{i},x^{\prime}_{j})\leq 2$.
Using the triangle inequality for twist numbers (Lemma 6.3, iv)), we get
  1. (2)
    the sequence $(\mathrm{tw}_{\alpha}(x_{i},x^{\prime}_{i}))$ is bounded by $\mathrm{tw}_{\alpha}(x_{i_{0}},x^{\prime}_{i_{0}})+4$.
Therefore the formula
$\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})=\sup\{t,\exists(x_{i})\to\xi,\exists(x^{\prime}_{i})\to\xi^{\prime},\forall i,\operatorname{tw}_{\alpha}(x_{i},x^{\prime}_{i})=t\}$
defines a natural number. Furthermore, for every admissible sequences $(x_{i}),(x^{\prime}_{i})$ converging to $\xi,\xi^{\prime}$, for every $i$ large enough, $\mathrm{tw}_{\alpha}(x_{i},x^{\prime}_{i})$ differs from $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})$ by at most $4$.
Definition 11.3.
The twist number between $\xi$ and $\xi^{\prime}$ is
$\mathrm{tw}(\xi,\xi^{\prime})=\sup_{\alpha\in\mathcal{C}^{\dagger}(S)}\operatorname{tw}_{\alpha}(\xi,\xi^{\prime}).$
This is either a natural number or $\infty$.
We have the following useful triangular inequalities.
Lemma 11.4 (Triangular inequalities).
  1. (1)
    For all $\xi,\xi^{\prime},\xi^{\prime\prime},\alpha$, we have $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime\prime})\leq\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})+\mathrm{tw}_{\alpha}(\xi^{\prime},\xi^{\prime\prime})$.
  2. (2)
    For all $\xi,\xi^{\prime},\xi^{\prime\prime}$, $\mathrm{tw}(\xi,\xi^{\prime\prime})\leq\mathrm{tw}(\xi,\xi^{\prime})+\mathrm{tw}(\xi^{\prime},\xi^{\prime\prime})$.
Proof.
The first one is proved by taking $x,x^{\prime\prime}$ close to $\xi,\xi^{\prime\prime}$ where the sup is (almost) attained, and any $x^{\prime}$ close to $\xi^{\prime}$. The second follows. ∎
Let us briefly explore the infinite case; we want to show that the twist number $\mathrm{tw}(\xi,\xi^{\prime})$ being infinite is a property of the individual points at infinity $\xi$ and/or $\xi^{\prime}$ rather than of the pair $\{\xi,\xi^{\prime}\}$. Fix $\xi,\xi^{\prime}$. Fix some $\alpha_{0}$ and some radius $R$. We claim that $\{\mathrm{tw}_{\alpha}(\xi,\xi^{\prime}),\alpha\in B_{R}(\alpha_{0})\}$ is bounded, where $B_{R}(\alpha_{0})$ denotes the $R$ -ball around $\alpha_{0}$. To prove the claim, take small neighborhoods $N,N^{\prime}$ of $\xi,\xi^{\prime}$ such that any geodesic segment joining two points in $N$ or two points in $N^{\prime}$ is outside $B_{R+r}(\alpha_{0})$. Fix some curves $x,x^{\prime}$ in $N,N^{\prime}$. Up to replacing them by nearby curves, we may assume they have finitely many intersection points. Let $\alpha\in B_{R}(\alpha_{0})$. Then $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})\leq\mathrm{tw}_{\alpha}(x,x^{\prime})+4$. On the other hand, $\mathrm{tw}_{\alpha}(x,x^{\prime})$ is bounded by the cardinal of $x\cap x^{\prime}$, since the number of intersection points is clearly an upper bound for the distance in the arc graph. Thus we get a bound on $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})$ for all $\alpha\in B_{R}(\alpha_{0})$.
Now let $(x_{i})_{i\in\mathbb{Z}}$ be a $K$-quasigeodesic path joining $\xi$ to $\xi^{\prime}$ for some $K$ (for the existence of quasigeodesics connecting boundary points, see for instance [LT25], Proposition 2.3). For every $\alpha$ outside the $r(K)$-neighborhood of this path, $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})\leq 2$; here $r(K)$ is the constant from the Bounded Geodesic Image Theorem. We now assume that $\mathrm{tw}(\xi,\xi^{\prime})=+\infty$: by definition there is a sequence $(\alpha_{i})$ such that $\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime})$ tends to $\infty$. According to the previous observations, such a sequence lies inside the $r(K)$ neighborhood of our path, and escape every ball. Up to extracting, such a sequence converges either to $\xi$ or $\xi^{\prime}$. There are two (overlapping) cases: (i) there are such sequences converging to $\xi$, or (ii) there are such sequences converging to $\xi^{\prime}$.
Definition 11.5.
We say that $\xi$ has infinite twist (relative to $\xi^{\prime}$) if there exists a sequence $(\alpha_{i})$, converging to $\xi$, such that the twist numbers $\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime})$ tends to $\infty$.
So the preceding discussion proves that if $\mathrm{tw}(\xi,\xi^{\prime})=+\infty$ then at least one of the two points $\xi,\xi^{\prime}$ has infinite twist. Furthermore, it turns out that $\xi$ having infinite twist (relative to $\xi^{\prime}$) has nothing to do with $\xi^{\prime}$.
Lemma 11.6.
If $\xi$ has infinite twist relative to $\xi^{\prime}$, then it has infinite twist relative to all other boundary points.
Proof.
Fix a quasigeodesic sequence $\gamma$ from $\xi^{\prime}$ to another boundary point, distinct from $\xi$, denoted $\xi^{\prime\prime}$. We want to check that if $\xi$ does not have infinite twist relative to $\xi^{\prime\prime}$, then it does not have infinite twist relative to $\xi^{\prime}$. Let $(\alpha_{i})$ be a sequence converging to $\xi$. For every $i$ large enough, the quasigeodesic $\gamma$ is distance more that $r(K)$ of $\alpha_{i}$, and thus $\mathrm{tw}_{\alpha_{i}}(\xi^{\prime},\xi^{\prime\prime})\leq 2$. Now by the triangular inequality, if the sequence $(\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime\prime}))$ is bounded, then so is $(\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime}))$. ∎
Now we want to define the twist number of a map. Let $f$ acts loxodromically on $\mathcal{C}^{\dagger}(S)$, and denote $\xi^{-}(f),\xi^{+}(f)$ its repulsive and attracting boundary fixed points. We can set, for each $\alpha\in\mathcal{C}^{\dagger}(S)$, $\mathrm{tw}_{\alpha}(f)=\mathrm{tw}_{\alpha}(\xi^{-}(f),\xi^{+}(f))$.
Definition 11.7.
The twist of $f$ is defined by
$\mathrm{tw}(f)=\mathrm{tw}(\xi^{-}(f),\xi^{+}(f))=\sup_{\alpha\in\mathcal{C}^{\dagger}(S)}\operatorname{tw}_{\alpha}(\xi^{-}(f),\xi^{+}(f)).$
Note that, at least on the torus (but this is most probably more general), $\mathrm{tw}(f)$ cannot be infinite, since twist numbers are bounded along axes (Lemma 10.6). In other words, boundary points with infinite twist cannot be fixed by any loxodromic element.
The Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ is equipped with a topology (see e.g. [BH99] for a general discussion of boundaries, and [BHW24, LT25] specifically for the fine curve graph).
Lemma 11.8.
The map $f\mapsto(\xi^{-}(f),\xi^{+}(f))$ which is defined on the subset of $\mathrm{Homeo}(S)$ of elements acting loxodromically, is continuous.
Proof.
Let $(f_{m})$ be a sequence of loxodromic isometries in some Gromov-hyperbolic metric space $X$, and $f$ be another loxodromic isometry. Assume the following ”simple” convergence: for every $x\in X$, the sequence $(f_{m}x)$ converges to $fx$. Then the sequence of boundary fixed points $(\xi^{+}(f_{m}))$ converges to $(\xi^{+}(f))$. This is a classical property.
Let us go back to our specific context. Let $(f_{m})$ be a sequence in $\mathrm{Homeo}(S)$ which converges to $f$ for the $C^{0}$ distance. In our context, the space $X=\mathcal{C}^{\dagger}(S)$ is a graph, the distance takes only integer values, and the ”simple” convergence does not hold. Instead, we have the following ”coarse simple convergence” (cf [BHM+22, Lemma 3.5]): for every $\alpha$, and every $n$ large enough, we have $d(f_{m}\alpha,f\alpha)\leq 2$. With this property, the convergence of $(\xi^{+}(f_{m}))$ to $(\xi^{+}(f))$ may be proved along the same lines as the continuity of the translation length ([BHM+22, Theorem 3.4]).
Here is a sketch of proof, details are left to the reader. According to [BHM+22, Lemma 3.8], up to replacing each map with a fixed power, we may assume the following. There is a curve $\alpha$ such that for $m$ large enough, for every $n$, and for every $i$ with $n\leq i\leq n$, the point $f_{m}^{i}\alpha$ is within $4\delta$ of the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$. Fix $i_{0}$ so that $f^{i_{0}}\alpha$ is close to $\xi^{+}(f)$, and $m$ large enough so that the previous property holds, and $d(f_{m}^{i_{0}}\alpha,f^{i_{0}}\alpha)\leq 2$: then the set $N=\{x\in\mathcal{C}^{\dagger}(S),(x,f_{m}^{i_{0}}\alpha)_{\alpha}\geq d(\alpha,f_{m}^{i_{0}}\alpha)-100\delta\}$ is a small neighborhood of $\xi^{+}(f)$. For every $n\geq i_{0}$, the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$ passes close to $f_{m}^{i_{0}}\alpha$, and thus the point $f_{m}^{n}\alpha$ belongs to $N$. We conclude that $\xi^{+}(f_{m})$ belongs to the closure of $N$, which is a small neighborhood of $\xi^{+}(f)$ in the $\mathcal{C}^{\dagger}(S)\cup\partial\mathcal{C}^{\dagger}(S)$. ∎
Lemma 11.9.
Let $\xi,\xi^{\prime}$ be two distinct boundary points.
Assume $\mathrm{tw}(\xi,\xi^{\prime})$ is finite. Then if $(\eta,\eta^{\prime})$ is sufficiently close to $(\xi,\xi^{\prime})$ then $\mathrm{tw}(\eta,\eta^{\prime})\geq\mathrm{tw}(\xi,\xi^{\prime})-4$.
Assume $\mathrm{tw}(\xi,\xi^{\prime})$ is infinite. Then $\lim_{(\eta,\eta^{\prime})\to\xi,\xi^{\prime})}\mathrm{tw}(\eta,\eta^{\prime})=+\infty$.
Proof.
Assume $\mathrm{tw}(\xi,\xi^{\prime})$ is finite, and let $\alpha$ so that $\mathrm{tw}(\xi,\xi^{\prime})=\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})$. For $(\eta,\eta^{\prime})$ sufficiently close to $(\xi,\xi^{\prime})$ we have $\mathrm{tw}_{\alpha}(\xi,\eta)\leq 2$ and $\mathrm{tw}_{\alpha}(\eta^{\prime},\xi^{\prime})\leq 2$. By the triangular inequality we get $\mathrm{tw}_{\alpha}(\eta,\eta^{\prime})\geq\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})-4$, and thus $\mathrm{tw}(\eta,\eta^{\prime})\geq\mathrm{tw}(\xi,\xi^{\prime})-4$. The infinite case is analogous. ∎
Remember that the map $f\mapsto\mathrm{tl}(f)$ is also continuous ([BHM+22, Theorem 3.4]). In particular being loxodromic is an open property. By chaining the two lemmas we immediately get the following.
Corollary 11.10 (Semi-continuity of twist numbers).
Given $f\in\mathrm{Homeo}(S)$ acting loxodromically on $\mathcal{C}^{\dagger}(S)$, if $g\in\mathrm{Homeo}(S)$ is sufficiently close to $f$, then $\mathrm{tw}(g)\geq\mathrm{tw}(f)-4$.
In the next subsection we will see that there are examples with arbitrarily large twist number. Together with the semi-continuity, this implies the following non density result. Remember that $\mathrm{Homeo}_{0}(S)$ acts minimally on the boundary, see [BH25, LT25].
Corollary 11.11.
If $\mathrm{tw}(\xi,\xi^{\prime})<+\infty$ then the orbit of $(\xi,\xi^{\prime})$ under the action of $\mathrm{Homeo}(S)$ on $\partial\mathcal{C}^{\dagger}(S)\times\partial\mathcal{C}^{\dagger}(S)$ is not dense.
Furthermore, the following consequence may be useful, e.g. to construct non trivial quasi-morphisms.
Corollary 11.12.
Assume that $|\mathrm{tw}(f)-\mathrm{tw}(g)|>4$. Then $f\not\sim g$ and $f\not\sim g^{-1}$ in the sense of Bestvina-Fujiwara. In particular, the four points $\xi^{-}(f),\xi^{+}(f),\xi^{-}(g),\xi^{+}(g)$ are distinct.
Proof.
Up to exchanging $f$ and $g$ we may assume that $\mathrm{tw}(g)<\mathrm{tw}(f)-4$, and thus by Lemma 11.9, the couple $(\xi^{-}(g)),\xi^{+}(g))$ may not be sent arbitrarily close to $(\xi^{-}(f)),\xi^{+}(f))$ by another element $h$. In other words, $f\not\sim g$. Since $\mathrm{tw}(g^{-1})=\mathrm{tw}(g)$, the second inequivalence follows. The last part of the corollary follows from Lemma 3.8. ∎
11.4. Computation of twist and translation length
Let us get back to the torus. We will illustrate twist numbers by showing how computations can be done on our favorite family of examples. The idea is that first a lower bound is obtained by using explicit annuli. Then we can use a particular quasigeodesic axis and the Bounded Geodesic Image Theorem to show that all annuli in which the twist $\mathrm{tw}_{A}(f)$ is big have their boundary close to the quasi-axis, thus close to one of our explicit annulus, and then use this information to give an upper bound. As a by-product, we are able to show that this quasi-axis is actually a geodesic axis, and thus to compute the precise value of the translation length on these examples.
We consider the maps $f$ and $g$ from the proof of Lemma 11.1. Let us fix some $p,q>0$ and denote $h=h_{p,q}=g^{q}f^{p}$. Let $\alpha,\beta$ be respectively the horizontal and vertical curves that bound the fundamental domain $0,1]^{2}$. Roughly speaking, $h$ twists with length $p$ in the complement of $\alpha$, and then twists with length $q$ in the complement of $\beta$. Note that $f\alpha=\alpha,g\beta=\beta$, $h\alpha=g^{q}\alpha$ intersects $\beta$ exactly once and transversely, so that the bi-infinite path
$\gamma=(\dots,h^{-1}\alpha,h^{-1}\beta,\alpha,\beta,h\alpha,h\beta,\dots)$
is an invariant quasi-axis for $h$. This already tells us that $\mathrm{tl}(h)\leq 2$, as noted in [LPSW24].
$\alpha$
$\beta$
$h(\alpha)$
$h(\beta)$
$h^{-1}(\alpha)$
$h^{-1}(\beta)$
$g^{q}$
$f^{p}$
Figure 18. Building an axis for $h$. The homeomorphisms $f^{p},g^{q}$ act as elliptic isometries which “rotate” around $\alpha,\beta$.
Let $B$ denotes the constant from the second Bounded Geodesic Image Theorem for geodesics. We may assume that $B\geq 2$.
Proposition 11.13.
Assume $p,q>B$. Then the path $\gamma$ is an invariant geodesic for $h$.
As an immediate corollary we get the translation length.
Corollary 11.14.
$\mathrm{tl}(h)=2$.
We denote $\tilde{\alpha}=\mathbb{R}\times\{0\},\tilde{\beta}=\{0\}\times\mathbb{R}$ our favorite lifts of $\alpha,\beta$. At this point, as a motivation for the next definitions, we suggest that the reader draws a picture of $h\beta$ (and maybe $h^{2}\beta$; compare also Figure 19). For every $k\geq 1$, let us call an arc in the plane a good $\beta$-arc of size $k$ if it has one end-point in $(0,0)$ and the other on the top side on the rectangle $0,1]\times[0,k]$ and is otherwise included in the interior of this rectangle. We define the good $\alpha$-arcs by exchanging the $x$ and $y$ coordinates. Note that every good $\beta$-arc contains a good $\beta$-arc of size $1$. Also note that $f$ fixes $\alpha$ pointwise, $g$ fixes $\beta$ pointwise, and in particular the point $(0,0)$ is fixed by both $f$ and $g$. The reader can easily check, by drawing pictures, the following properties:
  1. (1)
    $h(\alpha)=g^{q}\alpha$ contains a good $\beta$-arc of size $q$;
  2. (2)
    the image of a good $\beta$-arc under $f^{p}$ contains a good $\alpha$-arc of size $p$;
  3. (3)
    the image of a good $\alpha$-arc under $g^{q}$ contains a good $\beta$-arc of size $q$.
By an induction chaining these properties, we get that for every $n>0$, $h^{n}\alpha$ contains a good $\beta$-arc of size $q$. We can be a little more precise.
Lemma 11.15.
For every $n>0$, $h^{n}\alpha$ contains an arc included in the rectangle $0,1]\times[0,q+1]$, connecting the fixed point $(0,0)$ to the right-hand side of the rectangle, and containing a $\beta$-arc of size $q$.
Proof.
$b_{0}$
$f^{p}(b_{0})$
$a_{0}$
$g^{q}(a_{0})$
Figure 19. The proof of Lemma 11.15. $b_{0}\subset h^{n}\alpha$ is the $\beta$-arc of size $1$. The middle picture shows the image under $f^{p}$, with the initial segment $a_{0}$ which is an $\alpha$-arc of size $1$. The right picture shows the image of that subarc under $g^{q}$.
For $n=1$ this is easy. Let us assume $n>2$. We know that $h^{n-1}\alpha$ contains a $\beta$-arc of size $q$, and in particular a $\beta$-arc of size $1$. The image of this arc under $f^{p}$ contains an $\alpha$-arc of size $1$, and the image of such an arc under $g^{q}$ contains the wanted curve. ∎
The arc provided by the lemma is distance $q+2$ from $\alpha$ in the arc graph $\mathcal{A}^{\dagger}(A_{\beta})$, and no arc disjoint from it is distance $q+3$. Thus we get the twist number $\mathrm{tw}_{\beta}(\alpha,h^{n}\alpha)=q+2$.
Analogous considerations give, for every $n>0$, $\mathrm{tw}_{\beta}(\alpha,h^{-n}\alpha)=2$. Applying the triangular inequality for twists gives the first part of the following lemma. The second part follows from analogous considerations.
Lemma 11.16.
  1. (1)
    For every $m\geq 0,n>0$, $\mathrm{tw}_{\beta}(h^{-m}\alpha,h^{n}\alpha)\geq q$.
  2. (2)
    For every $m,n>0$, the curves $h^{-m}\alpha,h^{n}\alpha$ have non-empty projection under the map $\pi_{A_{\alpha}}$, and $\mathrm{tw}_{\alpha}(h^{-m}\alpha,h^{n}\alpha)\geq p$.
Proof of Proposition 11.13.
Let $n>0$. Consider a geodesic path $\gamma^{\prime}$ from $\alpha$ to $h^{n}\alpha$. For every $k=0,\dots,n-1$, we have, by point (1) of the lemma,
$\mathrm{tw}_{h^{k}\beta}(\alpha,h^{n}\alpha)=\mathrm{tw}_{\beta}(h^{-k}\alpha,h^{n-k}\alpha)\geq q>B$
and likewise, for $k=1,\dots,n-1$, $\mathrm{tw}_{h^{k}\alpha}(\alpha,h^{n}\alpha)\geq p>B$. By the Bounded Geodesic Image Theorem, the geodesic $\gamma^{\prime}$ has to ”lose projection” on each of the annuli determines by the following curves, which constitutes a segment of our bi-infinite path $\gamma$:
$\beta,h\alpha,h\beta,\dots,h^{n-1}\alpha,h^{n-1}\beta.$
This means that we can find a curves $c_{0},d_{1},\dots,d_{n-1},c_{n-1}$ which are vertices of the geodesic $\gamma^{\prime}$ and such that $\pi_{h^{k}\beta}(c_{k})=0$, $\pi_{h^{k}\alpha}(d_{k})=0$.
Claim 11.17.
The $c_{k}$’s and $d_{k}$’s are pairwise distinct.
Let us prove the claim. The curve $c_{k}$ does not project in the annulus determined by $h^{k}\beta$, thus it is homotopic to $h^{k}\beta$, which is homotopic to $\beta$. Likewise, each $d_{j}$ is homotopic to $\alpha$. Thus the $c_{k}$’s are certainly distinct from the $d_{j}$’s. Let us check that the $d_{j}$’s are also pairwise disjoint (the proof for the $c_{k}$’s is similar). By applying a power of $h$ it suffices to prove the following second claim.
Claim 11.18.
Let $k>0$. If $q>2$ then there is no curve $\delta$ such that both $\pi_{A_{\alpha}}(\delta)=\emptyset$ and $\pi_{h^{k}(A_{\alpha})}(\delta)=\emptyset$.
Assume for a contradiction that such a $\delta$ exists. By assumption $\delta$ does not cross the annulus determined by $\alpha$. Thus there is a lift $\tilde{\delta}$ which is between $\tilde{\alpha}$ and $\tilde{\alpha}+(0,2)$, which we denote
$\tilde{\alpha}<\tilde{\delta}<\tilde{\alpha}+(0,2).$
Likewise, there is an integer $\tau$ such that $\tilde{h}^{k}\tilde{\alpha}<\tilde{\delta}+(0,\tau)<\tilde{h}^{k}\tilde{\alpha}+(0,2)$, and thus
$\tilde{h}^{k}\tilde{\alpha}-(0,\tau)<\tilde{\delta}<\tilde{h}^{k}\tilde{\alpha}+(0,2-\tau).$
We first deduce that $\tilde{\alpha}<\tilde{h}^{k}\tilde{\alpha}+(0,2-\tau)$ which implies $2-\tau>0$ since $\tilde{h}^{k}\tilde{\alpha}$ meets $\tilde{\alpha}$ at the point $(0,0)$. And also $\tilde{h}^{k}\tilde{\alpha}-(0,\tau)<\tilde{\alpha}+(0,2)$, and since $\tilde{h}^{k}\tilde{\alpha}$ meets $\tilde{\alpha}+(0,q)$, this implies $2+\tau>q$. We conclude that $q-2<\tau<2$, and since all these numbers are integers, $q\leq 2$. This contradicts the assumption on $q$. This completes the proof of both claims.
From the first claim we conclude that the length of $\gamma^{\prime}$ is at least $2n$. Since this is exactly the length of the path $\gamma$ which has the same end-points, we conclude that $d(\alpha,h^{n}\alpha)=2n$. Thus $\gamma$ is a geodesic. ∎
In the above proof we seemingly were lucky to find exactly the good number of distinct annuli in which our geodesic $\gamma$ does not project. Actually the proof is a little more robust than this would make you believe: indeed, assume we had only found, say, 2n102n-10 such vertices; then we first conclude that $d(x,h^{n}(x))\geq 2n-9$, but then $\mathrm{tl}(h)\geq 2$, and from this it follows that $d(x,h^{n}(x))=2n$, since any smaller number would give a smaller translation length.
We finally search for an estimate on the twist number of $h$, showing that it is coarsely equal to $\max(p,q)$. We denote $B$ the constant from the second Bounded Geodesic Image Theorem, as above.
Proposition 11.19.
Assume $p,q>B$. Then
$\mathrm{tw}(h)\in[\max(p,q),\max(p,q)+4].$
Proof.
From Lemma 11.16 we get that
$\mathrm{tw}_{\beta}(\xi^{-}(h),\xi^{+}(h))\geq q,\mathrm{tw}_{\alpha}(\xi^{-}(h),\xi^{+}(h))\geq p$
and thus $\mathrm{tw}(h)\geq\max(p,q)$.
It remains to prove the upper bound. Assume that $\mathrm{tw}(h)>B+4$, otherwise there is nothing to prove. It suffices to prove that under this assumption, $\mathrm{tw}(h)\leq\max(p,q)+4$. By definition, there is some annulus $A=A_{c}$ for some curve $c$ such that $\mathrm{tw}_{A}(\xi^{-}(h),\xi^{+}(h))>B+4$. By the usual argument we may pick $n$ large enough so that $\mathrm{tw}_{A}(h^{-n}\alpha,h^{n}\alpha)\geq\mathrm{tw}_{A}(\xi^{-}(h),\xi^{+}(h))-4>B$. By the Bounded Geodesic Image Theorem, the geodesic $\gamma$ must lose projection in $A$: there is some $k$ such that $c$ does not project either in the annulus bounded by $h^{k}\alpha$ or in the annulus bounded by $h^{k}\beta$. We treat the first case, the second one is similar. The curve $c^{\prime}=h^{-k}c$ does not project in $A_{\alpha}$, and we have $\mathrm{tw}_{c^{\prime}}(h^{-n-k}\alpha,h^{n-k}\alpha)>B$. We can have chosen $n$ larger than $|k|+1$, so that nk-n-k and nkn-k are still respectively negative and positive. Since $\pi_{\alpha}(c^{\prime})=\emptyset$, there is a lift $\tilde{c^{\prime}}$ which is between $\tilde{\alpha}$ and $\tilde{\alpha}+(0,2)$. The strip $\tilde{A^{\prime}}$ bounded by $\tilde{c^{\prime}}$ and $\tilde{c^{\prime}}+(0,1)$ is a universal cover of the annulus $A^{\prime}=A_{c}$, and it is included in the strip $\tilde{A}$ bounded by $\tilde{\alpha}$ and $\tilde{\alpha}+(3,0)$. Let $A$ denote the annulus obtained by quotienting $\tilde{A}$ by the horizontal unit translation. Consider the torus $\mathbb{R}^{2}/(\mathbb{Z}\times 3\mathbb{Z})$, which is a degree 3 cover of our initial torus, and denote $\bar{h}$ a lift of $h$, and $\bar{\alpha}$ the projection of $\tilde{\alpha}$ in this cover. The annulus cut by $\bar{\alpha}$ in this cover identifies with $A$. The twist number $\mathrm{tw}_{A}(\bar{h}^{-n-k}\bar{\alpha},\bar{h}^{n-k}\bar{\alpha})$ is, by definition, the Hausdorff distance in the fine arc graph $\mathcal{A}^{\dagger}(A)$ between the projections of the curves $\bar{h}^{-n-k}\bar{\alpha}$ and $\bar{h}^{n-k}\bar{\alpha}$. The inclusion of strips implies that $\mathrm{tw}_{A^{\prime}}(h^{-n-k}\alpha,h^{n-k}\alpha)\leq\mathrm{tw}_{A}(\bar{h}^{-n-k}\bar{\alpha},\bar{h}^{n-k}\bar{\alpha})$.
It remains to check that, given $i,j>0$, $\mathrm{tw}_{A}(\bar{h}^{-i}\bar{\alpha},\bar{h}^{j}\bar{\alpha})\leq p+4$. Here we use the hypothesis that $q\geq 3$. The curve $\tilde{h}^{j}\tilde{\alpha}$ contains the image under $h$ of a good $\beta$-arc of size $1$, and this contains an arc $a^{+}$ originated in $(0,0)$ that crosses the strip $\tilde{A}$ and is included in $0,1]\times[0,3]$. The curve $\tilde{h}^{-i}(\tilde{\alpha}+(0,3))$ contains an arc $a^{-}$ originated from $(0,3)$ that crosses the strip $\tilde{A}$ and included in $p-1,0]\times[0,3]$. The arc $a^{-}$ meet at most $p+1$ translates of the arc $a^{+}$, thus their distance in the arc graph is at most $p+2$. Any arc disjoint from $a^{-}$ is at distance at most $p+4$ from any arc disjoint from $a^{+}$. Thus we have $\mathrm{tw}_{A}(h^{-i}\alpha,h^{j}\alpha)\leq p+4$, as wanted. ∎
Finally, we consider a variation on the family $h_{p,q}$ to provide an example where the twist number are useful to prove non conjugacy. Let $p,q,p^{\prime},q^{\prime}>0$, and consider $h=h_{p,q,p^{\prime},q^{\prime}}=g^{q^{\prime}}f^{p^{\prime}}g^{q}f^{p}$.
Proposition 11.20.
Assume $p,p^{\prime},q,q^{\prime}>B$. Then
$\mathrm{Rot}(h)=[0,p+p^{\prime}]\times[0,q+q^{\prime}],$
$\mathrm{tl}(h)=4,$
$\mathrm{tw}(h)\in[\max(p,p^{\prime},q,q^{\prime}),\max(p,p^{\prime},q,q^{\prime})+4]$
The proof is very similar to the previous proofs, and is left to the reader. As a corollary, we see that this family contains homeomorphisms that have the same rotation set and the same translation length but whose twist numbers differ by more than 4; in particular they are not conjugate, and by Corollary 11.12 they have pairwise distinct fixed points at infinity.
Let us end this subsection with more free groups. The rotation set of $h_{n,n}=g^{n}f^{n}$ is $0,n]^{2}$ and its translation length is $2$. Thus all the $h_{n,n}$’s have distinct scaled rotation sets. By Theorem 3 they have distinct boudary fixed points, though they have homothetic rotation sets (and thus Theorem 1 does not directly apply). The ping-pong argument tells us that up to taking powers, any two of them generate a free group. Likewise, Proposition 11.20 provides plenty of examples of pairs $h,h^{\prime}$ of maps sharing the same scaled rotation sets, but with distinct twist numbers. Choose $h,h^{\prime}$ so that their twists numbers are more than 4 apart. Then by Corollary 11.12 $h$ and $h^{\prime}$ have no boundary fixed points in common. Again, they have powers generating a free group.
11.5. Schottky groups
We now collect some examples which show that some generalizations of our main results which one might hope for are in fact false.
All of these examples rely on a Schottky group acting on the fine curve graph. We begin with two homeomorphisms $f,g\in\mathrm{Homeo}_{0}(T)$ with distinct rotation sets and independent axes in the fine curve graph (note that this is implied by our main result). A standard ping-pong argument (or by using Lemma 3.1) one shows that by replacing $f,g$ by large enough powers we may assume that $f,g$ generate a free group $F_{2}<\mathrm{Homeo}_{0}(T)$ so that the orbit map
$F_{2}\to\mathcal{C}^{\dagger}(T)$
is a quasi-isometric embedding. For our examples to work in this general setting, we would need to know that the scaled rotation set of $fg^{n}$ converges to that of $g$ when $n$ tends to $\infty$ (see question 2.6). We now specify $f$ and $g$ so that this kind of convergence holds. We will assume that
  1. (1)
    There are two disjoint topological closed disks $C(f),D(f)$ of $T^{2}$, such that $f$ is the identity on $D(f)$, and every rational vector in $\mathrm{Rot}(f)$ is the rotation vector of some periodic point in $C(f)$,
  2. (2)
    Likewise for $g$,
  3. (3)
    $C(f)\subset D(g),C(g)\subset D(f)$.
Examples with these properties may be constructed the following way. Let $P$ be a rational polygon containing 0. Kwapisz [Kwa92] constructs an Axiom A diffeomorphisms $f$ whose rotation set is $P$, whose periodic orbits belong to some Cantor set, except for some sinks and sources that have rotation vector 0, and 0 is also the rotation vector of some fixed point in the Cantor set. One may blow-up one of the source or sink to get a disk $D(f)$ of fixed points disjoint from the Cantor set. Now one can choose $C(f)$ to be another disc that contains the Cantor set and is disjoint from $D(f)$. The map $g$ is constructed analogously, and then conjugated to get property 3.
These properties have the following consequence. Let $h$ be some element of the Schottky group generated by $f$ and $g$, and respectively denote $n,m$ the sum of powers of $f$ and $g$ in $h$. Iterating $h$ on a point of $C(f)$ amounts to iterating just $f$, and likewise for $g$. Thus we get the lower bound
$\mathrm{Rot}(h)\supseteq\frac{1}{n+m}\left(n\mathrm{Rot}(f)+m\mathrm{Rot}(g)\right).$
Combining this with the upper bound lemma 2.4, we get that for every $k$,
$\frac{1}{m}\mathrm{Rot}(f^{k}g^{m})\xrightarrow{m\to\infty}\mathrm{Rot}(g).$
We can now construct our examples.
  • Consider the element $\varphi_{k,m}=f^{k}g^{m}$. By choosing $k$ large, a quasi-axis for $\varphi_{k,m}$ can be made to fellow-travel a quasi-axis for $f$ for an arbitrarily long time (since this is true in the Cayley graph of the free group, and the fact that it quasi-isometrically embeds into the fine curve graph). On the other hand, we have that the scaled rotation set of $f^{k}g^{m}$ converges to the scaled rotation set of $g$ if $k$ is fixed and $m\to\infty$.
  • Consider an infinite path in $F_{2}$ of the form
    $P=f^{n_{1}}g^{m_{1}}f^{n_{2}}g^{m_{2}}\cdots$
    Observe that for any homeomorphisms $A,B$ of the plane, by the Misiurewicz-Ziemian theorem we have
    $\frac{1}{m}AB^{m}([0,1]^{2})=\frac{1}{m}(ABA^{-1})^{m}(A[0,1]^{2})\xrightarrow{m\to\infty}\mathrm{Rot}(B).$
    Thus, given $n_{i},m_{i},i\leq N$ we can choose $n_{i+1},m_{i+1}$ so that
    $f^{n_{1}}g^{m_{1}}\cdots f^{n_{i}}g^{m_{i}}f^{n_{i+1}}[0,1]^{2}$
    is projectively close to the rotation set of $f$, and
    $f^{n_{1}}g^{m_{1}}\cdots f^{n_{i}}g^{m_{i}}f^{n_{i+1}}g^{m_{i+1}}[0,1]^{2}$
    is projectively close to the rotation set of $g$.
    As a consequence, the infinite product of $f,g$ given by $P$ does not have a well-defined projective rotation set. Hence, there is also no obvious way to associate a (projective) rotation set to an arbitrary boundary point of the fine curve graph.